### v4

# Bridging Privacy and Robustness for Trustworthy Machine Learning

Abstract
The advent of machine learning has led to transformative changes across various domains, but the sensitive nature of data raises concerns about privacy and security. While Local Differential Privacy (LDP) has been a cornerstone in addressing these concerns, recent research has proposed privacy concepts aligned with the Bayesian inference perspective of an adversary, such as Average Bayesian Privacy (ABP) and Maximum Bayesian Privacy (MBP). This paper explores the intricate relationships between LDP, ABP, and MBP, and their implications for algorithmic robustness. We establish theoretical connections between these privacy notions, proving that LDP implies MBP and vice versa under certain conditions, and deriving bounds connecting MBP and ABP. We also investigate the relationship between PAC robust learning and privacy preservation, demonstrating how to derive PAC robustness from privacy-preserving algorithms and construct privacy-preserving algorithms from PAC robust ones. Our findings provide valuable insights for constructing privacy-preserving and robust machine learning algorithms.
Index Terms:
Privacy-Preserving Algorithms; PAC Robust Learning; Trustworthy Machine Learning; Theoretical Connections
I Introduction
The advent of machine learning has heralded transformative changes across various domains, from healthcare to finance, by extracting meaningful insights from vast swathes of data. However, this data often contains sensitive information, raising legitimate concerns regarding privacy and security. The traditional paradigm of Local Differential Privacy (LDP) has been a cornerstone in addressing these concerns, aiming to ensure that the influence of any individual data point on the output of a computation is bounded, thereby obfuscating its presence or absence (Dwork, 2008). While LDP provides a strong foundation for privacy protection, it is essential to consider the adversary’s methods for potentially compromising privacy.
Recent research has proposed privacy concepts aligned with the Bayesian inference perspective of an adversary (He et al., 2019; Zhu et al., 2019), such as Average Bayesian Privacy (ABP) and Maximum Bayesian Privacy (MBP). These concepts have emerged from an understanding that adversaries may possess prior knowledge and use their Bayesian reasoning capabilities to infer sensitive information from the output of machine learning algorithms. By incorporating the adversary’s perspective, these new paradigms offer a more comprehensive approach to privacy protection, taking into account the potential strategies an adversary might employ to infer sensitive information.
As these Bayesian privacy concepts gain traction, it is crucial to thoroughly investigate their relationships with traditional privacy measures like LDP. Understanding the connections and differences between these approaches will enable researchers and practitioners to develop more robust and effective privacy-preserving mechanisms. By leveraging the strengths of both LDP and Bayesian privacy concepts, we can work towards building machine learning systems that safeguard sensitive information while still extracting valuable insights from data.
In this paper, we aim to explore the intricate relationships between LDP, ABP, and MBP, providing a deeper understanding of how these privacy notions interact and complement each other. Through rigorous mathematical analysis and empirical studies, we seek to uncover the key factors that influence the effectiveness of these privacy protection mechanisms and provide guidance on how to design and implement them in real-world machine learning applications. By bridging the gap between traditional and Bayesian privacy concepts, we hope to contribute to the development of more secure and trustworthy machine learning systems that respect individual privacy while unlocking the full potential of data-driven insights. Through meticulous definitions and theorems, we elucidate how these privacy concepts are interconnected. A key result (Theorem IV.4) establishes that, under certain conditions, LDP implies MBP and vice versa. We also derive bounds connecting MBP and ABP (Theorem IV.6), providing valuable insights for constructing privacy-preserving algorithms.
The intersection of privacy and learning theory, particularly Probably Approximately Correct (PAC) learning, has attracted considerable scholarly inquiry. Groundbreaking research by Kasiviswanathan et al. (2011) revealed that a vast array of classes learnable by non-private algorithms can also be mastered privately, albeit with a polynomial increase in sample size. Building upon these foundational intersections, our work extends the discourse by delving into the robustness implications of privacy guarantees. Robustness in machine learning, the ability of algorithms to maintain performance despite inaccuracies or noise in the input data, has become increasingly important as models are deployed in unpredictable real-world environments. The concept of robustness has been traditionally studied within the PAC framework (Cullina et al., 2018), which provides a theoretical foundation for the performance of learning algorithms under perturbations. However, while previous studies have predominantly focused on sample complexity and learnability, the intersection of robustness with privacy has not been fully explored, particularly how robust learning can contribute to, or benefit from, privacy-preserving mechanisms. In this paper, we take a deeper dive into how privacy-preserving mechanisms can enhance robustness in PAC learning scenarios, and conversely, how robust learning algorithms can inform the design of privacy-preserving mechanisms. We investigate the relationship between PAC robust learning and privacy preservation, demonstrating how to derive PAC robustness from algorithms satisfying specific privacy standards (Theorem IV.7). This exploration aims to bridge the gap between robustness and privacy, providing valuable insights for developing machine learning algorithms that are both resilient to perturbations and secure against privacy breaches.
The main contributions of this paper are as follows:
  • We establish theoretical connections between different privacy notions, proving that LDP implies MBP and vice versa under certain conditions (Theorem IV.4) and deriving relationship between MBP and ABP (Theorem IV.6).
  • We investigate the relationship between PAC robust learning and privacy preservation, demonstrating how to derive PAC robustness from privacy-preserving algorithms (Theorem IV.7) and construct privacy-preserving algorithms from PAC robust ones.
The remainder of this paper is organized as follows. Section II-A reviews related work on privacy measurements, Section II-D discusses robustness measurements, and Section II-E explores the relationship between privacy and robustness in the literature. In Section III, we introduce our privacy attack and defense framework. In Section IV, we define metrics for privacy and robustness, and establish theoretical connections between privacy notions and PAC robustness. We also introduce practical estimation approaches and analyses for MBP and ABP in F and G. Finally, we conclude the paper and discuss future research directions.
For convenience, we list the important notations used throughout this paper in Table LABEL:table:_notation.
TABLE I: Table of Notation
Notation Meaning
$\epsilon_{p}$ privacy leakage of the protector
$\mathcal{L}$ loss function
$\Delta$ distortion extent
$\breve{W}$ the original model parameter
$W$ the released model parameter
$\mathcal{X}$ a set of instances
$\mathcal{Y}$ a set of labels
$P$ distribution over instances $\mathcal{X}$
$d$ dataset
TV(·) total variation distance
$\mathcal{A}$(·) original model
$\mu$/$\mu(P)$ statistic of distribution $P$
$\mathcal{D}^{\prime}$ another dataset different from $\mathcal{D}$
$n$ size of dataset $\mathcal{D}$
$L$ Lipschitz Constant
$I$ number of optimization rounds
$F^{\mathcal{A}}$ belief distribution after observing protected information
$F^{\mathcal{B}}$ belief distribution before observing protected information
$F^{\mathcal{O}}$ belief distribution after observing unprotected information
$f^{\mathcal{O}}_{D}(d),f^{\mathcal{B}}_{D}(d),f^{\mathcal{A}}_{D}(d)$ density function of $F^{\mathcal{O}},F^{\mathcal{B}},F^{\mathcal{A}}$
$d_{H}$(·,·) Hamming distance
$d^{(m)}$ m-th data in dataset
$\xi$ maximum bayesian privacy constant
$f_{D|W}(d|w)$ posterior probability density function
$f_{D}(d)$ prior distribution
$f_{W|D}(w|d)$ likelihood function based on observed parameter $W$
$\kappa_{1}$ $f_{D}^{\mathcal{O}}(d)$
$\hat{\kappa_{1}}$ estimator of $\kappa_{1}$
$\kappa_{2}$ $f_{D}^{\mathcal{B}}(d)$
$\hat{\kappa_{2}}$ estimator of $\kappa_{2}$
$\kappa_{3}$ MBP, i.e., $\xi$
$\hat{\kappa_{3}}$ estimator of $\kappa_{3}$
II Related Work
II-A Privacy Measurements
Differential privacy (DP), a rigorously formalized privacy framework pioneered by Dwork et al. (2006), has been widely adopted to protect individual privacy in database systems. Abadi et al. (2016) proposed a method for training neural networks with a limited privacy budget, combining advanced machine learning methods and privacy-preserving mechanisms. This method is applicable to various optimization methods, including Momentum Rumelhart et al. (1986) and AdaGrad Duchi et al. (2011). Triastcyn and Faltings (2020) introduced Bayesian Differential Privacy (BDP), an adaptation of DP that incorporates data distribution considerations, along with a theoretical analysis and privacy computing framework, demonstrating its advantages through experimental results.
Rassouli and Gündüz (2019) employed average total variation distance (TV distance) for privacy measurement and minimum mean-square error (MMSE), mutual information, and error probability to assess utility, aiming to find the best balance between utility and privacy in data analysis tasks. Blauw and von Solms (2017) proposed two main privacy metrics based on an information-theoretic approach: information entropy and mutual information, to quantify and define online user privacy.
Achieving pure DP can be challenging in practical learning scenarios. Eilat et al. (2020) quantified the loss of privacy using relative entropy (Kullback-Leibler divergence) and defined the concept of privacy loss, referring to the information leakage caused by the posterior update of the agent’s prior beliefs. Foulds et al. (2016) discussed the constraints of obtaining privacy via posterior sampling and highlighted the benefits of the Laplace mechanism in Bayesian inference from both theoretical and empirical perspectives.
Other works have considered privacy measures akin to our approach. In social network settings, Alemany et al. (2019) proposed two metrics (Audience and Reachability) for the privacy risk of sharing information. du Pin Calmon and Fawaz (2012) discussed the minimization of maximum information leakage in privacy-preserving mapping, proposing a mathematical model and a formula that minimizes the maximum information leakage, intended to balance privacy against utility constraints, with privacy leakage assessed using KL divergence. In contrast, our work adopts Jensen-Shannon (JS) divergence for measuring privacy leakage, in line with Zhang et al. (2022, 2023d), due to its symmetric nature and the fact that its square root meets the triangle inequality Nielsen (2019); Endres and Schindelin (2003), aiding in quantifying privacy-utility trade-offs more effectively.
II-B Relation between Distinct Privacy Measurements
The advent of differential privacy has inspired numerous scholars to propose new privacy metrics and explore their relationships. Table III summarizes key relations between various privacy metrics, including our proposed Maximum Bayesian Privacy (MBP) and Average Bayesian Privacy (ABP).
TABLE II: Relations between various privacy metrics.
Reference Privacy Metric1 Relation To Other Metrics
Dwork and Rothblum (2016) Concentrated DP DP $\Rightarrow$ Concentrated DP
Mironov (2017) Rényi-DP Rényi-DP $\Rightarrow$ $(\epsilon,\delta)$-DP
Mironov et al. (2009)2 IND-CDP, SIM-CDP SIM-CDP $\Rightarrow$ SIM∀∃-CDP3 $\Rightarrow$ IND-CDP
Dong et al. (2019) $f$-DP Gaussian DP4 $\Rightarrow$ Rényi-DP
Blum et al. (2008) Distributional Privacy Distributional Privacy $\Rightarrow$ DP
Triastcyn and Faltings (2020) Bayesian-DP Bayesian DP $\Rightarrow$ DP
Cuff and Yu (2016) Mutual-Information DP MI-DP $\Leftrightarrow$ DP (asymptotically)
Our Work $\xi$-MBP, $\epsilon$-ABP $\xi$-LDP $\Rightarrow$ $\xi$-MBP $\Rightarrow$ $2\xi$-LDP, MBP $\Rightarrow$ ABP
  • 1
    All "DP"s are shorthand for differential privacy.
  • 2
    This article proposes Computational Differential Privacy (CDP). SIM-CDP and IND-CDP are two concrete kinds of CDP.
  • 3
    Definition of SIM∀∃-CDP is extended from SIM-CDP.
  • 4
    Gaussian DP is a specific family of $f$-DP.
The introduction of differential privacy has catalyzed the development of numerous privacy metrics. Dwork and Rothblum (2016) proposed Concentrated Differential Privacy (Concentrated-DP), which relaxes differential privacy based on the mean privacy leakage of a randomized mechanism. Mironov (2017) introduced Rényi Differential Privacy (RDP), a variant based on Rényi divergence that unifies concentrated-DP, $\epsilon$-DP, and $(\epsilon,\delta)$-DP. Mironov et al. (2009) developed Computational Differential Privacy (CDP), which measures privacy against computationally-bounded adversaries. They introduced IND-CDP and SIM-CDP, establishing that SIM-CDP implies SIM∀∃-CDP, which in turn implies IND-CDP. Dong et al. (2019) presented f-Differential Privacy (f-DP) and Gaussian Differential Privacy as relaxations of traditional DP. Blum et al. (2008) proposed Distributional Privacy, offering stronger protection than traditional differential privacy. Triastcyn and Faltings (2020) introduced Bayesian Differential Privacy (BDP), which considers the dataset’s distribution and implies traditional DP.
Recent research has focused on creating taxonomies and comparing various privacy definitions. Cuff and Yu (2016) provided a comprehensive comparison of privacy notions, including differential privacy and its variants, introducing Mutual-Information Differential Privacy (MI-DP) and demonstrating its asymptotic equivalence to traditional differential privacy. Meiser (2018) examined the nuances of differential privacy, emphasizing the importance of distinguishing between approximate and probabilistic differential privacy. Sommer et al. (2019) proposed a unified approach to measuring privacy loss in sensitive data through privacy loss distribution (PLD), revealing that for non-adaptive mechanisms, privacy loss tends toward a Gaussian distribution, enabling tighter privacy bounds and more accurate calculations for mechanisms such as the Gaussian mechanism under ADP.
Desfontaines and Pejó (2019) classified privacy definitions into seven distinct groups, each modifying different aspects of the original concept. These groups serve as dimensions, allowing the combination of variants from different groups to create new definitions. The study also established a partial ordering of their strength and identified which definitions exhibit properties like composition, post-processing, and convexity. Guingona et al. (2023) compared approximate and probabilistic differential privacy, exploring the implications of their relationship and finding that significant adjustments to one parameter necessitate substantial adjustments to another. Ünsal and Önen (2023) investigated the connection between differential privacy and information-theoretic measures, demonstrating how these connections can provide novel interpretations of privacy in systems employing differential privacy.
Our work contributes to this growing body of knowledge by introducing Maximum Bayesian Privacy (MBP) and Average Bayesian Privacy (ABP), and establishing their relationships with Local Differential Privacy (LDP). Specifically, we prove that $\xi$-LDP implies $\xi$-MBP, which in turn implies $2\xi$-LDP. We also demonstrate that MBP implies ABP.
Our approach to reductions differs from previous works in that we focus on Bayesian privacy notions and their connections to LDP. While previous taxonomies have primarily focused on variations of differential privacy, our work bridges the gap between traditional DP concepts and Bayesian privacy notions. This provides a new perspective on privacy guarantees that takes into account an adversary’s prior knowledge and inference capabilities.
The content of our taxonomy complements existing work by:
  1. 1.
    Introducing new Bayesian privacy metrics (MBP and ABP)
  2. 2.
    Establishing concrete relationships between these Bayesian metrics and LDP
  3. 3.
    Providing a bridge between traditional DP concepts and Bayesian privacy notions
These results offer new insights into the design of privacy-preserving algorithms, particularly in scenarios where adversaries may have prior knowledge about the data distribution. By situating our work within the broader context of privacy taxonomy research, we contribute to a more comprehensive understanding of privacy definitions and their interrelationships, aiding researchers and practitioners in selecting appropriate privacy metrics for their specific use cases.
II-C Intersections of Privacy and PAC Learning
The dynamic interplay between privacy and learning theory, with a spotlight on PAC learning, has attracted considerable scholarly inquiry. In groundbreaking research, Kasiviswanathan et al. (2011) scrutinized the concept classes amenable to learning under differential privacy. Their findings revealed that a vast array of classes that are learnable by non-private algorithms can also be mastered privately, albeit with a polynomial increase in sample size. This study highlighted the striking equivalence between local private learning algorithms and statistical query learning, underscoring that privacy constraints do not substantially impede the learning process.
Concurrently, Bun (2020) delved into the computational nuances of private PAC learning, drawing a clear line between online learning and private PAC learning paradigms. Their work shed light on the evolving landscape of integrating privacy considerations into learning theory, unveiling both the formidable challenges and promising opportunities in the pursuit of efficient private learning.
In a parallel vein, Feldman and Xiao (2014) explored the nexus between differential privacy and communication complexity within the realm of learning algorithms. They established a profound equivalence between the sample complexity of differentially private learning and the randomized one-way communication complexity required to evaluate concepts. This pivotal discovery is monumental as it forges a link between two distinct domains, offering valuable insights into the development of private learning algorithms and presenting a novel lens through which to view the sample complexity bounds of learning tasks. In a more recent development, Agarwal and Singh (2017) made significant strides in the context of online learning by crafting differentially private algorithms for online convex optimization. They demonstrated that it is feasible to uphold privacy without markedly sacrificing the regret bounds, thus underscoring the intricate equilibrium that must be struck between privacy assurances and learning efficacy in dynamic environments.
Building upon these foundational intersections between privacy and PAC learning, our work extends the discourse by delving into the robustness implications of privacy guarantees. While previous studies have predominantly focused on sample complexity and learnability, we take a deeper dive into how privacy-preserving mechanisms can enhance robustness in PAC learning scenarios, and vice versa, how robust learning algorithms can be harnessed to design privacy-preserving mechanisms.
By establishing theoretical linkages between our proposed privacy notions—MBP and ABP—and PAC robustness, we offer a fresh perspective on the privacy-robustness interplay. Our findings augment the existing body of literature by shedding light on how privacy guarantees can bolster the resilience of learning algorithms to data perturbations, and how robust learning techniques can be leveraged to safeguard individual privacy.
II-D Robustness Measurements
Measuring the robustness of classifiers, especially neural network-based ones, has been an active research area in recent years. Hendrycks and Dietterich (2019) proposed two new notions for measuring robustness: corruption robustness and perturbation robustness, which consider a classifier’s average-case performance on corruptions $\mathcal{C}$ and perturbations $\xi$, respectively. In contrast, adversarial robustness measures the worst-case performance on small, additive, classifier-tailored perturbations or corruptions.
For neural network classifiers, Bastani et al. (2016) formalized the notion of $(x_{\ast},\epsilon)$-robust and defined point-wise robustness, $\rho(f,x_{\ast})$, as the $L_{\infty}$ distance from $x_{\ast}$ to the nearest adversarial example. They proposed two statistics to estimate these quantities: adversarial frequency and adversarial severity.
In the context of binary classification, Zhang et al. (2019) characterized model robustness using the robust classification error under the threat model of bounded $\epsilon$ perturbation, decomposing the robust error into the sum of natural error and boundary error to characterize the trade-off between accuracy and robustness.
Cohen et al. (2019) introduced the concept of certifiable robustness, where a classifier is considered certifiably robust if, for any input $\mathbf{x}$, one can easily obtain a guarantee that the classifier’s prediction remains constant within some set around $\mathbf{x}$. By employing randomized smoothing, they demonstrated the feasibility of certified defense on the ImageNet dataset.
These works provide valuable insights and tools for quantifying and understanding the robustness of classifiers, paving the way for developing more robust models in the face of various types of perturbations and corruptions.
II-E Relation between Robustness and Privacy
The relationship between privacy and robustness has been widely discussed in the literature (Table III). Dwork and Lei (2009) explored the field of robust statistics and demonstrated that robust statistical estimators provide a strong foundation for developing differentially private estimators. Pinot et al. (2019) proposed a generalized notion of adversarial robustness using the divergence $D_{\lambda}$ of probabilistic distributions and suggested using randomization methods to improve robustness against adversarial samples, discussing the similarities between differential privacy and adversarial sample robustness.
bansal2020extending introduced the concept of robust privacy, which defines the proximity and indistinguishability between two distributions without relying on a specific notion of accuracy or the underlying data distribution. Asi et al. (2023) proposed a black-box transformation method to convert a robust statistical estimator into a private statistical estimator using the $\rho$-smooth inverse sensitivity mechanism, proving that a private estimator ($\epsilon$-DP) can naturally be robust, albeit with some loss of accuracy in certain high-dimensional statistical tasks and increased computational inefficiency due to the transformation.
Lecuyer et al. (2019) introduced the PixelDP model, which provides meaningful certified robustness boundaries and reasonable assessments of model accuracy degradation across various datasets and deep neural networks, offering higher accuracy on larger models and datasets, as well as stronger robustness against p-norm attacks compared to previous authentication defense methods. Hopkins et al. (2023) presented the first black-box reduction from privacy to robustness, capable of producing private estimators with optimal trade-offs among sample complexity, accuracy, and privacy for a wide range of fundamental high-dimensional parameter estimation problems, proving that, under certain conditions, if a statistical estimation method is robust to small perturbations of the input data, it can also protect the privacy of the input data.
Pillutla et al. (2022) explored the trade-offs among privacy, robustness, and communication in federated learning, proposing a robust aggregation method that replaces the weighted arithmetic mean aggregation with a generalized form of the classical multidimensional median. This method protects the update privacy of devices by iteratively invoking secure multi-party computation primitives and bounds the sensitivity to each device’s contribution, with numerical simulations demonstrating its effectiveness in handling data corruption and parameter update corruption.
TABLE III: relation between privacy and robustness.
Reference Privacy Metric Robustness Metric Relation
Dwork and Lei (2009) $(\epsilon,\delta)$-DP M-estimators robust covariance Propose-Test-Release
matrices robust regression paradigm
Lecuyer et al. (2019) $(\epsilon,\delta)$-PixelDP certifiably robust $(\epsilon,\delta)-DP$+certain
condition $\Rightarrow$ robustness
Pinot et al. (2019) Renyi-DP $D_{\mathcal{P}(\mathcal{\gamma})}$-$(\alpha,\epsilon,\gamma)$ $D_{\mathcal{P}(\mathcal{\gamma})}$-$(\alpha,\epsilon,\gamma)$ robust
robust $\Leftrightarrow$ Renyi-DP
bansal2020extending Flexible Accuracy Laplace Mechanism robustness $\Leftrightarrow$ privacy
Robust Privacy $M_{Lap}^{f,b}$
Pillutla et al. (2022) secure median’s be compatible only with
average oracle breakdown point more communication cost1
Hopkins et al. (2023) $(\epsilon,\delta)$-DP $(\eta,\beta,\alpha)$-robust robustness $\Leftrightarrow$ privacy
Asi et al. (2023) $\epsilon$-DP $(\tau,\beta,\alpha)$-robust privacy $\Leftrightarrow$ robustness2
[…]
our work $\xi$-MBP $(\tau,\beta,\alpha)$-robust privacy $\Leftrightarrow$ robustness
$\epsilon$-ABP $(\tau,\beta,\alpha)$-robust robustness $\Rightarrow$ privacy
  • 1
    No quantitative relation is provided in this work
  • 2
    when corrupted ratio $\tau\approx\log(n)/n\epsilon$, private estimator is naturally robust, robust estimator can be converted to private estimator.
  • 3
    when corrupted ratio $\tau\geq\frac{2(d\cdot\log(\frac{R}{\alpha_{0}}+1)+\log\frac{2}{\beta})}{n\epsilon}$, robust estimator can be converted to nearly private estimator.
III Privacy Attack and Defense Framework
This section delineates the adversarial landscape confronting data privacy in machine learning applications. We systematically outline the threat model, characterizing the adversary’s objectives, capabilities, knowledge base, and methodologies for launching attacks. Concurrently, we describe the defender’s strategies that underpin the protection mechanisms aimed at safeguarding sensitive information. It is worth noting that this privacy attack and defense framework has been widely investigated, including Zhang et al. (2023a); du Pin Calmon and Fawaz (2012); Zhang et al. (2023b); Kang et al. (2022); Zhang et al. (2023f); Asi et al. (2023).
III-A Privacy Attack Model
The threat model provides a structured perspective on the potential vulnerabilities and risks to client data privacy. Herein, we examine the various facets of an attacker’s strategy, including their intents, resources, intelligence, and tactics.
Attacker’s Objective
Within our framework, we posit a scenario where a protector operates an algorithm $\mathcal{M}(\cdot)$, processing input data $\mathcal{D}$ to yield output $\breve{W}$. This result must then be conveyed to a recipient—who is presupposed to be the attacker. The attacker aspires to extract confidential details regarding $\mathcal{D}$ from the observable $\breve{W}$, while balancing the need to maintain utility within certain bounds. The underlying objective is to refine the attacker’s probabilistic beliefs about $\mathcal{D}$ through analysis of $\breve{W}$.
Attacker’s Capability
Our model assumes a semi-honest (or “honest-but-curious”) attacker. Such an adversary adheres to protocol specifications in terms of computations but seeks to glean additional private information by inferring it from $\breve{W}$. By performing statistical inference on the output $\breve{W}$, the attacker aims to ascertain insights into the protector’s confidential data $\mathcal{D}$.
Attacker’s Knowledge
Based on previous descriptions, the attacker has access to the output $\breve{W}$ (though post-transformation by the protector’s mechanism, they receive a modified version $W$). Furthermore, the attacker may possess pre-existing knowledge about the private data $\mathcal{D}$, potentially acquired through past breaches or leaks. Additionally, the attacker is cognizant of the utility constraints associated with their analyses.
Attacker’s Method
With knowledge of a distorted parameter $W$, we assume that the attacker adopts a Bayesian inference attack, i.e., find a data $d$ that optimizes posterior belief, denoted by:
$\arg\max_{d}\text{IH}(d|w)=\arg\max_{d}[\text{I}(w|d)+\text{H}(d)],$ (1)
where $\text{IH}(d|w)$ represents the logarithm of the posterior probability density function $f_{{D}|{W}}(d|w)$, $\text{I}(w|d)$ is the likelihood function’s logarithm based on observed parameter $w$, and $\text{H}(d)$ symbolizes the prior probability distribution’s logarithm $f_{{D}}(d)$. According to Bayes’ theorem, maximizing the logarithm of posterior $f_{D|W}(d|w)$ relative to $D$ involves the aggregation of both $\log(f_{W|D}(w|d))$ and $\log(f_{D}(d))$.
III-B Privacy Protection Model
We now turn our attention to the defensive measures employed by the protector whose design is to shield sensitive datasets from unauthorized exposure.
Protector’s Objective
To safeguard privacy post-computation by algorithm $\mathcal{M}(\cdot):\mathcal{D}\rightarrow\breve{W}$, the protector needs to devise a privacy-preserving mechanism, which is a stochastic transformation, having the ability to map $\breve{W}\rightarrow W$, where $W$ is a less-informative derivative. This ensures compliance with privacy standards such as $\epsilon$-differential privacy.
Protector’s Capability
It is incumbent upon the protector to transmit $W$ in lieu of $\breve{W}$ to conceal the original data $\mathcal{D}$. To this end, the protector must be adept at generating and communicating a perturbed variant $W$, which duly obscures the private dataset $\mathcal{D}$. This task may be complicated by historical disclosures of $\mathcal{D}$.
Protector’s Knowledge
The protector is privy to the private dataset $\mathcal{D}$, the computed output $\breve{W}$, and the intricacies of the employed algorithm $\mathcal{M}(\cdot)$. They are also versed in the workings of the privacy-preserving transformation. Additionally, an understanding of the adversary’s utility limitations is crucial for the protector to tailor the distortion of $\breve{W}$ into $W$ effectively, satisfying the informational requirements of the adversary whilst preserving the sanctity of $\mathcal{D}$.
Protector’s Method
In defence of privacy (i.e., the confidentiality of dataset $\mathcal{D}$), the protector must execute the randomized privacy-preserving mechanism, which transforms $\breve{W}$ to $W$. Execution of the algorithm is expected to equip it with the requisite properties for safeguarding dataset $\mathcal{D}$ against intrusion while adhering to the adopted privacy standard. For instance, optimized algorithms for protecting privacy under the defense and attacking framework have been designed as detailed in the studies by Zhang et al. (2023f) and du Pin Calmon and Fawaz (2012), while Asi et al. (2023) choose to apply exponential mechanism(i.e., add random noise to $\breve{W}$) to protect privacy.
IV Connecting Privacy Measurements and PAC Robustness in Machine Learning
This section lays the theoretical groundwork for understanding the intricate relationships between different privacy notions and their connection to algorithmic robustness. It commences by delving into the theoretical ties among Maximum Bayesian Privacy (MBP), Average Bayesian Privacy (ABP), and Local Differential Privacy (LDP). Through meticulous definitions and theorems, it elucidates how these privacy concepts are interconnected, enabling transitions between various privacy frameworks. Fundamental results establish that, under certain conditions, LDP implies MBP and vice versa. The section also explores the link between MBP and ABP, offering valuable insights for constructing privacy-preserving algorithms. Furthermore, it investigates the relationship between PAC robust learning algorithms and privacy preservation. PAC robust learning, a performance metric ensuring approximately correct learning even with slightly perturbed data, is examined in the context of privacy protection. The section demonstrates how to derive PAC robustness from algorithms satisfying specific privacy standards and how to construct privacy-preserving algorithms from PAC robust ones. These findings provide a solid theoretical foundation for designing robust learning algorithms that simultaneously offer privacy protection capabilities. By bridging the gap between privacy notions and robustness, this section paves the way for developing algorithms that achieve both data confidentiality and resilience to perturbations.
IV-A Metrics for Privacy and Robustness
This section introduces the key metrics used to quantify privacy and robustness in the proposed defense and attack framework. It provides detailed definitions for privacy measures such as Maximum Bayesian Privacy, Average Bayesian Privacy, and Local Differential Privacy, enabling researchers to evaluate the privacy risks associated with different models. The section also defines PAC Robustness as a measure of an algorithm’s consistency against data variations. These metrics form the foundation for developing and assessing algorithms that balance data utility and confidentiality.
IV-A1 Privacy Measurements
Privacy measurement in machine learning has become a crucial area of research, with various approaches proposed to quantify and evaluate the privacy guarantees of different algorithms and models. This section provides a detailed exposition on the metrics used to quantify privacy under the previously discussed privacy attack and defense framework, including concepts such as Maximum Bayesian Privacy, Average Bayesian Privacy, and Local Differential Privacy. These approaches enable researchers to quantitatively evaluate the privacy risks associated with different models and algorithms. As privacy concerns continue to grow, further advancements in privacy measurement techniques are critical for developing privacy-preserving machine learning solutions. For methodologies and analyses concerning the practical estimation of MBP and ABP, please refer to F and G.
First, we define the concept of average Bayesian privacy and explore its correlation with the notion of maximum Bayesian privacy in Section IV-B. These metrics serve as crucial benchmarks in evaluating the resilience of data to inference attacks. The average privacy leakage (Definition IV.1) measures the variation between the adversary’s prior and posterior beliefs on private information. The adversary obtains the posterior belief through mounting the Bayesian inference attack on the protected model information. Thus, we formally call the privacy leakage the Bayesian privacy leakage. The protected model information may lead to the federated model with lower utility and the federated training with less efficiency in certain scenarios.
[…]
We use JS divergence to measure the privacy leakage instead of KL divergence. The advantage of JS divergence over KL divergence is that it is symmetrical, and its square root satisfies the triangle inequality (Endres and Schindelin (2003)). This property facilitates the quantification of the trade-offs. Now we introduce the average Bayesian privacy of the protection mechanism.
Definition IV.1 (Average Bayesian Privacy).
Let $\mathcal{M}$ represent the protection mechanism that maps from the private information $D$ to the model information $W$, and $\mathcal{G}$ represent the attacking mechanism. Let $\epsilon_{p}$ represent the privacy leakage, which is defined as:
$\epsilon_{p}(\mathcal{M},\mathcal{G})=\sqrt{{\text{JS}}(F^{\mathcal{A}}||F^{\mathcal{B}})},$ (2)
where ${\text{JS}}(F^{\mathcal{A}}||F^{\mathcal{B}})=\frac{1}{2}\int_{\mathcal{D}}f^{\mathcal{A}}_{D}(d)\log\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)+\frac{1}{2}\int_{\mathcal{D}}f^{\mathcal{B}}_{D}(d)\log\frac{f^{\mathcal{B}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)$, $F^{\mathcal{A}}$ and $F^{\mathcal{B}}$ represent the attacker’s belief distribution about $D$ upon observing the protected information and without observing any information respectively, and $f_{D}^{\mathcal{M}}(d)=\frac{1}{2}(f^{\mathcal{A}}_{D}(d)+f^{\mathcal{B}}_{D}(d))$. We say $\mathcal{M}$ is $\epsilon$-ABP when $\epsilon_{p}(\mathcal{M},\mathcal{G})\leq\epsilon$.
Remark:
We operate under the assumption that the private information $D$ takes on discrete values, and then
${\text{JS}}(F^{\mathcal{A}}||F^{\mathcal{B}})$ $=\frac{1}{2}\left[\text{KL}\left(F^{\mathcal{A}}||F^{\mathcal{M}}\right)+\text{KL}\left(F^{\mathcal{B}}||F^{\mathcal{M}}\right)\right]$
$=\frac{1}{2}\sum_{d\in\mathcal{D}}f^{\mathcal{A}}_{D}(d)\log\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}+\frac{1}{2}\sum_{d\in\mathcal{D}}f^{\mathcal{B}}_{D}(d)\log\frac{f^{\mathcal{B}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}.$
Now we provide the formal definition for $(\epsilon,\beta,\alpha)$-ABP.
Definition IV.2 ($(\epsilon,\beta,\alpha)$-ABP).
Algorithm $\mathcal{M}$ achieves $(\epsilon,\beta,\alpha)$-ABP for a distribution $P$, if it satisfies: (1)being $\epsilon$-ABP, i.e., the privacy leakage $\epsilon_{p,a}\leq\epsilon$, (2)achieving standard accuracy $\alpha$ with a confidence of $1-\beta$.
The Maximum Bayesian Privacy (denoted by $\epsilon_{p,m}$) encapsulates the highest degree of privacy assurance that can be sustained against an adversary with partial or complete knowledge about the data distribution. This metric is situated within the Bayesian privacy paradigm, which assesses the risk associated with the disclosure of private information. It underscores the importance of maintaining privacy when data undergoes sharing or utilization for designated tasks.
Definition IV.3 ($\xi$-Maximum Bayesian Privacy).
Let $F$ denote the belief distribution held by the attacker regarding the private data, and let $f$ indicate the associated probability density function. We define the system to possess $\xi$-Maximum Bayesian Privacy if for any $w\in\mathcal{W}$ and $d\in\mathcal{D}$, the following inequality is satisfied:
$e^{-\xi}\leq\frac{f_{D|W}(d|w)}{f_{D}(d)}\leq e^{\xi}.$
The $\xi$-Maximum Bayesian Privacy quantifies the level of privacy protection afforded by measuring the disparities in the conditional and marginal distributions of the data. A smaller value of $\xi$ implies stronger privacy guarantees against potential adversaries.
Definition IV.4 (PAC Standard Accuracy).
Let $\mathcal{M}$ be an algorithm tasked with estimating a statistic $\mu$, and $d$ represent the private information. The algorithm $\mathcal{M}$ is said to possess standard accuracy $\alpha$ with confidence $1-\beta$ if:
$\|\mathcal{M}(d)-\mu\|\leq\alpha$ (3)
holds with probability at least $1-\beta$.
The accuracy is measured within the framework of Probably Approximately Correct (PAC) learning following Zhang et al. (2023c).
Definition IV.5 ($(\epsilon,\beta,\alpha)$-MBP).
Let $\mathcal{M}$ denote an algorithm designed for the estimation of statistic $\mu$. Algorithm $\mathcal{M}$ achieves $(\epsilon,\beta,\alpha)$-MBP for a distribution $P$, if it satisfies two criteria: being $\epsilon$-MBP, and achieving standard accuracy $\alpha$ with a confidence of $1-\beta$. Here, $\alpha$ symbolizes a constant deviation threshold.
IV-A2 PAC Robustness Measurement
In parallel with privacy, we also define robustness within the PAC framework, specifically addressing mean-estimation algorithms.
Definition IV.6 (PAC Robust Accuracy).
Consider $\mathcal{M}$ as an algorithm for estimating statistic $\mu$, and $d$ represent the private information. The algorithm $\mathcal{M}$ is considered to have robust accuracy $\alpha$ with confidence $1-\beta$, if for all datasets $d^{\prime}\neq d$, the following holds with probability at least $1-\beta$:
$\|\mathcal{M}(d^{\prime})-\mu(P)\|\leq\alpha.$ (4)
Consequently, $\mathcal{M}$ is termed $\alpha$-PAC-Robust with probability $1-\beta$.
This definition essentially evaluates the consistency of the algorithm’s performance against variations in the input data. A robust algorithm remains accurate even when small portions of the dataset are altered, thus reflecting resilience against data perturbations. This definition motivates the design of an algorithm $\mathcal{A}(\mathcal{D})$ that is sufficiently accurate (and potentially randomized) for estimating a parameter $\mu(P)$, given that the input dataset $\mathcal{D}$ is drawn from a specific distribution $P$.
Both the privacy and robustness measurements presented here play instrumental roles in developing and assessing algorithms within the intricately balanced domain of data utility and confidentiality. The nuanced understanding of these metrics enables practitioners to formulate strategies that concurrently enhance the security and integrity of the analytical processes while mitigating risks to data privacy.
IV-B Theoretical Connections between Privacy Notions
In the following, relations among MBP, ABP and LDP are investigated. Through mathematical definitions and theorems, we demonstrate how these concepts of privacy are interconnected and how one can transition from one privacy framework to another. Specifically, we will prove through theorems that an algorithm satisfying the requirements of LDP also satisfies those of MBP, and vice versa. Additionally, we will discuss the connection between MBP and ABP, along with how to convert from a mapping that provides maximum privacy protection to one that offers an average level of privacy protection. These theoretical results not only deepen our understanding of the connections between various privacy measures but also offer significant guidance for crafting and analyzing algorithms that preserve privacy.
LDP aims to safeguard individual data contributors’ privacy, enabling the extraction of aggregate statistics or insights from the data. The concept of LDP is articulated as follows.
Definition IV.7 ($\epsilon$-LDP).
A randomized function $\mathcal{M}$ defined as $\epsilon$-local differentially private if, $\forall$ possible datasets $d$ and $d^{\prime}$, $forall$ subset S of the output space, the following inequality is true
$\frac{\Pr[\mathcal{M}(d)\in S]}{\Pr[\mathcal{M}(d^{\prime})\in S]}\leq e^{\epsilon}.$ (5)
In this formula, $\epsilon$ is non-negative, which quantifies the level of privacy protection, and $\Pr[\mathcal{M}(d)\in S]$ represents the probability that the randomized mechanism $\mathcal{M}$ outputs a result in the set S when given the dataset $d$.
This definition states that the probability of any outcome being produced by the mechanism when given dataset $d$ should be similar to the probability of the same outcome being produced when given a dataset $d^{\prime}$, smaller $\epsilon$ values indicate stronger privacy guarantees. LDP ensures that individual contributions to the dataset are protected by adding noise or randomness to the results of queries or computations, making it difficult to infer specific details about any individual in the dataset.
The following Lemma IV.1 implicates that if an algorithm provides $\xi$-LDP, then it also offers $\xi$-MBP. In other words, the same mechanism that provides local privacy protection, where individual data points are protected, also ensures a certain level of global privacy in a Bayesian sense. This means that the mechanism is not only useful for protecting individual data points but also for safeguarding the privacy of the dataset as a whole. Please check C-A for detailed analysis and proof.
Lemma IV.1.
Let MBP be defined using Definition IV.3. Consider a case where the adversary’s prior belief distribution follows a uniform distribution. That is, $f_{D}(d)=f_{D}(d^{\prime})$, $\forall d^{\prime},d\in\mathcal{D}$. If $f_{W|D}(\cdot|\cdot)$ is $\xi$-LDP. That is, the mapping $f_{W|D}(\cdot|\cdot)$ satisfies that $\forall d^{\prime},d\in\mathcal{D}$,
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
then $f_{W|D}(\cdot|\cdot)$ constitutes a privacy-preserving transformation ensuring $\xi$-Maximum Bayesian Privacy, where for all $d\in\mathcal{D}\ and\ w\in\mathcal{W}$,
$\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|\leq\xi.$ (6)
Conversely, the following lemma states that if an algorithm offers $\xi$-MBP, it also provides $2\xi$-LDP. This suggests that mechanisms crafted to safeguard the privacy of the dataset as a whole, as evaluated by Bayesian privacy, also afford a degree of protection for individual data points within a local differential privacy framework.
Lemma IV.2.
Let MBP be defined using Definition IV.3. Consider a situation where the adversary’s prior belief distribution follows a uniform distribution. Consequently, $f_{D}(d)=f_{D}(d^{\prime})$ for any $d,d^{\prime}$. If an algorithm is $\xi$-Maximum Bayesian Privacy, then it is $(2\xi)$-differentially private. Suppose privacy-preserving mapping $f_{W|D}(\cdot|\cdot)$ ensures $\xi$-maximum Bayesian privacy. That is, $\xi=\max\limits_{w\in\mathcal{W},d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|$. Then, the function $f_{W|D}(\cdot|\cdot)$ achieves $(2\xi)$-differential privacy
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-2\xi},e^{2\xi}].$
This lemma establishes a bridge between local differential privacy and maximum Bayesian privacy, showing that mechanisms satisfying one privacy criterion also offer a certain level of privacy protection according to the other criterion. This provides insights into the relationship between different privacy definitions and can be useful when designing and analyzing privacy-preserving algorithms in various contexts.
With Lemma IV.1 and Lemma IV.2, we are now ready to derive the relationship between MBP and LDP.
Theorem IV.3 (Theoretical Relationship between MBP and LDP).
Consider a situation in which the adversary’s prior belief distribution follows a uniform distribution. Let MBP be defined using Definition IV.3. If an algorithm is $\xi$-LDP, then it is $\xi$-MBP. If an algorithm is $\xi$-MBP, then it is $2\xi$-LDP.
The first conclusion of Theorem IV.4 addresses a scenario in which the prior belief distribution of an adversary closely resembles a uniform distribution, meaning that the adversary has a relatively unbiased and uniform belief about the data points in the set $d$. The lemma introduces an algorithm with a privacy-preserving mapping denoted as $f_{W|D}(\cdot|\cdot)$. This mapping provides a maximum Bayesian privacy guarantee, where privacy is quantified by the parameter $\xi$. The parameter $\xi$ quantifies the maximum privacy loss resulting from the mapping, determined by the logarithm of the ratio between the conditional posterior distribution of a data point $d$ given an observation $w$ and the prior distribution of $d$. In other words, $\xi$ measures the maximum privacy leakage under this mechanism.
The second conclusion of Theorem IV.4 implies that if the algorithm provides $\xi$-maximum Bayesian privacy, then it also provides $(2\xi)$-differential privacy. This implies that the privacy-preserving mapping not only demonstrates robustness in terms of Bayesian privacy but also fulfills a corresponding concept of privacy, namely, differential privacy. The differential privacy guarantee of mapping is expressed in terms of the ratios of conditional probabilities. Specifically, the ratio of the conditional probabilities of observing $w$ given two different data points $d$ and $d^{\prime}$ falls within the range of $[e^{-2\xi},e^{2\xi}]$. This range specifies the extent to which the mapping distorts the likelihood of observing $w$ given different data points. The larger $\xi$ is, the wider this range becomes, indicating stronger privacy protection. This conclusion shows a connection between maximum Bayesian privacy and differential privacy. It illustrates that a mapping ensuring maximum Bayesian privacy, particularly when the adversary’s prior belief distribution approximates uniformity, also confers a degree of differential privacy. This finding holds significance in comprehending and quantifying the privacy assurances offered by such mappings within the realm of privacy-preserving data analysis.
We could also explore the situation where the prior belief distribution of the adversary is $\epsilon$-close to a uniform distribution and the analytical approach would be akin to the current scenario.
Theorem IV.4 (Generalized Relationship between MBP and LDP).
Consider the scenario where the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution, namely $\forall d\text{ and }d^{\prime},\ |\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$. Let MBP be defined using Definition IV.3. If an algorithm is $\xi$-LDP, then it is $(\xi+\epsilon)$-MBP. If an algorithm is $\xi$-MBP, then it is $(2\xi+\epsilon)$-LDP.
Remark: We might also explore the situation where the prior belief distribution of the adversary is $\epsilon$-close to a uniform distribution and the analytical approach would be akin to the current scenario.
The following theorem demonstrates the relationship between average Bayesian privacy and maximum Bayesian privacy.
Theorem IV.5 (Relationship Between MBP and ABP).
Assume that the prior belief of the attacker satisfies that
$\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}(d)}\in[e^{-\epsilon},e^{\epsilon}].$
Let MBP be defined using Definition IV.3. For any $\epsilon_{p,m}\geq 0$, suppose privacy-preserving mapping $f_{W|D}(\cdot|\cdot)$ ensuring $\epsilon_{p,m}$-Maximum Bayesian Privacy. Namely,
$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\epsilon_{p,m}},e^{\epsilon_{p,m}}],$
for any $w\in\mathcal{W}$. Then, we have that the ABP (denoted as $\epsilon_{p,a}$) is bounded by
$\epsilon_{p,a}\leq\frac{1}{\sqrt{2}}\sqrt{(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right)},$
where $\epsilon_{p,a}$ is introduced in Definition IV.1.
The following theorem demonstrates the generalized relationship between average Bayesian privacy and maximum Bayesian privacy. This theorem establishes a generalized relationship between MBP and ABP under the assumption that the ratio of the attacker’s prior belief $f^{\mathcal{B}}_{D}(d)$ and the true prior distribution $f_{D}(d)$ is bounded within the range $[e^{-\epsilon},e^{\epsilon}]$. The theorem states that if a privacy-preserving mapping $f_{W|D}(\cdot|\cdot)$ guarantees $\epsilon_{p,m}$-MBP, i.e., the ratio of the posterior belief $f_{D|W}(d|w)$ and the prior belief $f_{D}(d)$ is bounded within the range $[e^{-\epsilon_{p,m}},e^{\epsilon_{p,m}}]$ for any $w\in\mathcal{W}$ and $d\in\mathcal{D}$, then the ABP (denoted as $\epsilon_{p,a}$) is bounded by a function of $\epsilon_{p,m}$ and $\epsilon$. Specifically, the theorem provides an upper bound on $\epsilon_{p,a}$ in terms of $\epsilon_{p,m}$ and $\epsilon$, which can be used to analyze the relationship between MBP and ABP under certain conditions on the attacker’s prior belief. This result is useful for understanding the connections between different notions of Bayesian privacy and can be leveraged to design privacy-preserving algorithms with desired properties.
Theorem IV.6 (Generalized Relationship Between MBP and ABP).
Assume that the prior belief of the attacker satisfies that
$\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}(d)}\in[e^{-\epsilon},e^{\epsilon}].$
Let MBP be defined using Definition IV.3. For any $\epsilon_{p,m}\geq 0$, let $f_{W|D}(\cdot|\cdot)$ be a privacy preserving mapping that guarantees $\epsilon_{p,m}$-maximum Bayesian privacy. That is,
$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\epsilon_{p,m}},e^{\epsilon_{p,m}}],$
for any $w\in\mathcal{W}$. Then, we have that the ABP (denoted as $\epsilon_{p,a}$) is bounded by
$\epsilon_{p,a}\leq\frac{1}{\sqrt{2}}\sqrt{(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right)},$
where $\epsilon_{p,a}$ is introduced in Definition IV.1.
IV-C PAC Robustness of Private Algorithms
This subsection establishes a significant result linking the $(\epsilon,\beta,\alpha)$-MBP privacy criterion to PAC robustness. It proves that any algorithm satisfying MBP inherently possesses a quantifiable level of robustness against data perturbations, with probabilistic guarantees. This finding bridges the gap between theoretical privacy and practical robustness, offering insights for designing machine learning models that are both privacy-preserving and resilient.
The following Theorem IV.7 states that an algorithm $\mathcal{M}$, conforming to the $(\epsilon,\beta,\alpha)$-MBP privacy criterion, exhibits a specified degree of robustness against perturbations in the dataset. The robustness metric is defined as $(\frac{\log(1/\gamma)}{2n\epsilon},\alpha)$, representing the bounds within which the algorithm’s output is expected to vary. This level of robustness is probabilistically assured with a confidence of $1-\beta/\gamma$. The parameter $\gamma$ functions as a tunable confidence level within the exclusive range of 0 and 1, while $n$ quantifies the dataset’s size, and $\beta$ indicates the algorithm’s permissible error rate.
Theorem IV.7 (PAC Robustness of $\epsilon$-MBP Protection Mechanisms).
Let $\mathcal{M}$ be an $(\epsilon,\beta,\alpha)$-MBP (Definition IV.5) algorithm. Then $\mathcal{M}$ is $\alpha$-PAC-Robust (Definition IV.6) with probability $1-\gamma$, where $\gamma=(1+4\epsilon)\beta$, $n$ represents the size of the dataset, and $\beta$ represents the error probability.
This theorem elucidates a significant implication for privacy-preserving algorithms in machine learning, specifically those adhering to the $(\epsilon,\beta,\alpha)$-MBP model. It articulates that any algorithm $\mathcal{M}$ satisfying this MBP criterion inherently possesses a quantifiable level of PAC robustness. This robustness is defined in terms of $\alpha$ and $\gamma$, where $\gamma$ is a parameter that influences the confidence level, $n$ is the dataset size, and $\beta$ is the permitted error probability. Crucially, the theorem assures that with probability $1-\beta/\gamma$, the algorithm will maintain its robust performance despite potential variations in the data, thereby offering a probabilistic guarantee of robustness in the face of data perturbations. This bridges the gap between theoretical privacy guarantees and practical robustness in applied machine learning models.
This theorem is significant as it establishes a formal connection between robustness in the presence of learning errors and measurable criteria for privacy protection. It sets the stage for exploring the crucial relationship between data privacy and algorithmic robustness in machine learning systems. It highlights the importance of understanding and quantifying this connection to develop algorithms that simultaneously protect sensitive information and maintain performance in the face of perturbations or adversarial attacks. The theorem motivates the need for theoretical frameworks and metrics to analyze the privacy-robustness trade-off.
V Conclusions
In this paper, we have explored the intricate relationships between different privacy notions, namely LDP, ABP, and MBP, and their implications for algorithmic robustness. We established theoretical connections between these privacy concepts, proving that LDP implies MBP and vice versa under certain conditions, and deriving bounds between MBP and ABP. These findings provide valuable insights for constructing privacy-preserving algorithms. We also investigated the relationship between PAC robust learning and privacy preservation, demonstrating how to derive PAC robustness from privacy-preserving algorithms and construct privacy-preserving algorithms from PAC robust ones. This connection highlights the interplay between privacy and robustness in machine learning. Our findings contribute to the growing body of research on privacy and robustness in machine learning, providing a deeper understanding of the connections between different privacy notions and their implications for algorithmic robustness. By leveraging these insights, researchers and practitioners can work towards building more secure and trustworthy machine learning systems that safeguard sensitive information while maintaining robustness to perturbations and adversarial attacks. Future research directions include further exploration of the privacy-robustness trade-off, development of efficient algorithms that achieve both data confidentiality and resilience to perturbations, and investigation of the practical implications of these theoretical findings in real-world machine learning applications. As the field of machine learning continues to evolve, ensuring the privacy and robustness of these algorithms will remain a critical challenge and an area of active research.
Appendix A Relationship between LDP and MBP
In this section, we derive the relationship between LDP and MBP.
A-A LDP implies MBP
The adversary has a prior belief distribution, denoted as $f_{D}(d)$, that is very close (within $\epsilon$) to a uniform distribution. In other words, the adversary doesn’t have a strong prior bias towards any particular mini-batch data in the whole set $\mathcal{D}$.
There exists a mapping function $f_{W|D}(\cdot|\cdot)$, which takes a mini-batch data $d$ and maps it to $w$ such that for any two mini-batch data $d$ and $d^{\prime}$ in $\mathcal{D}$, the ratio of the conditional probabilities of mapping to $w$ remains within a certain range, specifically between $e^{-\xi}$ and $e^{\xi}$. This means that the mapping process is not overly sensitive to the specific input data.
The mapping function $f_{W|D}(\cdot|\cdot)$ is said to provide a "privacy-preserving" transformation that guarantees a maximum Bayesian privacy level of $\xi$. In other words, for any observed value $w$ and any mini-batch data $d$ in $\mathcal{D}$, the logarithm of the ratio of the conditional posterior distribution of $d$ given $w$ to the prior distribution of $d$ (i.e., the Bayesian privacy leakage) does not exceed $\xi$ in absolute value. This means that the mapping function provides strong privacy protection by limiting the information leakage about the original mini-batch data $d$ when the adversary observes $w$.
The following lemma establishes a connection between Local Differential Privacy (LDP) and Maximum Bayesian Privacy (MBP) under the assumption of a uniform prior belief distribution of the adversary. It states that if a mapping $f_{W|D}(\cdot|\cdot)$ satisfies $\xi$-LDP, then it also guarantees $\xi$-MBP. In other words, if the ratio of the conditional probabilities $f_{W|D}(w|d)$ and $f_{W|D}(w|d^{\prime})$ is bounded within the range $[e^{-\xi},e^{\xi}]$ for any $w\in\mathcal{W}$ and $d,d^{\prime}\in\mathcal{D}$, then the mapping $f_{W|D}(\cdot|\cdot)$ is a privacy-preserving mapping that ensures the log-ratio of the posterior belief $f_{D|W}(d|w)$ and the prior belief $f_{D}(d)$ is bounded by $\xi$ for all $w\in\mathcal{W}$ and $d\in\mathcal{D}$. This lemma provides a useful connection between LDP and MBP, which can be leveraged to design privacy-preserving algorithms.
Lemma A.1 (LDP implies MBP).
Consider the scenario where the prior belief distribution of the adversary is a uniform distribution. That is, $f_{D}(d)=f_{D}(d^{\prime})$ for any $d,d^{\prime}\in\mathcal{D}$. If $f_{W|D}(\cdot|\cdot)$ is $\xi$-LDP. That is, the mapping $f_{W|D}(\cdot|\cdot)$ satisfies that
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
then $f_{W|D}(\cdot|\cdot)$ is a privacy-preserving mapping that guarantees $\xi$-maximum Bayesian privacy. That is, $\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|\leq\xi$, $\forall w\in\mathcal{W},d\in\mathcal{D}$.
Proof.
From the definition of $\xi$-LDP, we know that
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
for all $d,d^{\prime}\in\mathcal{D}$. Notice that
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}=\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}.$
It implies that $\frac{f_{D|W}(d|w)}{f_{D}(d)}\leq\frac{f_{D|W}(d^{\prime}|w)}{f_{D}(d^{\prime})}\cdot e^{\xi}$.
We assume that $f_{D}(d)=f_{D}(d^{\prime})$. Notice that
$\sum_{d\in\mathcal{D}}f_{D|W}(d|w)=1.$
For any fixed $\tilde{d}\in\mathcal{D}$, we have that
11 11 $=\sum_{d\in\mathcal{D}}f_{D|W}(d|w)$
$=f_{D|W}(\tilde{d}|w)+\sum_{d\in\mathcal{D}\setminus\tilde{d}}f_{D|W}(d|w)$
$\leq f_{D|W}(\tilde{d}|w)+e^{\xi}\cdot f_{D|W}(\tilde{d}|w)\cdot(|\mathcal{D}|-1)$
$=(1+e^{\xi}\cdot(|\mathcal{D}|-1))\cdot f_{D|W}(\tilde{d}|w).$
Therefore, we have that
$f_{D|W}(\tilde{d}|w)\geq\frac{1}{1+e^{\xi}\cdot(|\mathcal{D}|-1)}.$
Similarly, we have that
11 11 $=\sum_{d\in\mathcal{D}}f_{D|W}(d|w)$
$=f_{D|W}(\tilde{d}|w)+\sum_{d\in\mathcal{D}\setminus\tilde{d}}f_{D|W}(d|w)$
$\geq f_{D|W}(\tilde{d}|w)+e^{-\xi}\cdot f_{D|W}(\tilde{d}|w)\cdot(|\mathcal{D}|-1)$
$=(1+e^{-\xi}\cdot(|\mathcal{D}|-1))\cdot f_{D|W}(\tilde{d}|w).$
Therefore, we have
$f_{D|W}(\tilde{d}|w)$ $\leq\frac{1}{1+e^{-\xi}\cdot(|\mathcal{D}|-1)}$
$\leq\frac{e^{\xi}}{|\mathcal{D}|}.$
Therefore, $\forall\tilde{d}\in\mathcal{D}$, we have that
$f_{D|W}(\tilde{d}|w)\leq e^{\xi}f_{D}(\tilde{d}).$
We say a mechanism is $\xi$-MBP if $\forall w\in\mathcal{W}^{d}$,
$\frac{f(d|w)}{f(d)}=\frac{f(w|d)f(d)}{f(w)f(d)}=\frac{f(w|d)}{f(w)}\leq e^{\xi}.$ (7)
Therefore, $f_{D|W}(\cdot|\cdot)$ is $\xi$-MBP. ∎
A-B MBP implies LDP
The following lemma illustrates that if an algorithm is $\xi$-maximum Bayesian privacy, then it is $2\xi$-Differentially Private.
Lemma A.2 (MBP implies LDP).
Consider the scenario where the prior belief distribution of the adversary is a uniform distribution. That is, $f_{D}(d)=f_{D}(d^{\prime})$ for any $d,d^{\prime}\in\mathcal{D}$. Let $f_{W|D}(\cdot|\cdot)$ be a privacy-preserving mapping that guarantees $\xi$-maximum Bayesian privacy. That is, $\xi=\max_{w\in\mathcal{W},d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|$. Then, the mapping $f_{W|D}(\cdot|\cdot)$ is $2\xi$-Local Differential Private:
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-2\xi},e^{2\xi}].$
Proof.
From the definition of maximum Bayesian privacy leakage, we know that
$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\xi},e^{\xi}],$
for any $w\in\mathcal{W}$ and $d\in\mathcal{D}$. Therefore, for any $d,d^{\prime}\in\mathcal{D}$ we have
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}=\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}\in[e^{-2\xi},e^{2\xi}].$
From the definition of Differential Privacy, we know that it is $2\xi$-Differentially Private. ∎
Combining the results in C-A and C-B, we are now ready to analyze the relationship between MBP and LDP, which is shown in the following theorem.
Theorem A.3 (Relationship between MBP and LDP).
[…]
Appendix B Relationship Between Average Bayesian Privacy and Maximum Bayesian Privacy
In this section, we introduce the relationship between average Bayesian privacy and maximum Bayesian privacy.
Theorem B.1 (Relationship Between MBP and ABP).
Assume that the prior belief of the attacker satisfies that
$\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}(d)}\in[e^{-\epsilon},e^{\epsilon}].$
Let MBP be defined using Definition IV.3. For any $\epsilon_{p,m}\geq 0$, let $f_{W|D}(\cdot|\cdot)$ be a privacy preserving mapping that guarantees $\epsilon_{p,m}$-maximum Bayesian privacy. That is,
$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\epsilon_{p,m}},e^{\epsilon_{p,m}}],$
for any $w\in\mathcal{W}$. Then, we have that the ABP (denoted as $\epsilon_{p,a}$) is bounded by
$\epsilon_{p,a}\leq\frac{1}{\sqrt{2}}\sqrt{(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right)},$
where $\epsilon_{p,a}$ is introduced in Definition IV.1.
Proof.
Let $F^{\mathcal{M}}=\frac{1}{2}(F^{\mathcal{A}}+F^{\mathcal{B}})$. We have
$JS(F^{\mathcal{A}}||F^{\mathcal{B}})$ $=\frac{1}{2}\left[KL\left(F^{\mathcal{A}}||F^{\mathcal{M}}\right)+KL\left(F^{\mathcal{B}}||F^{\mathcal{M}}\right)\right]$
$=\frac{1}{2}\left[\int_{\mathcal{S}}f^{\mathcal{A}}_{D}(d)\log\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)+\int_{\mathcal{S}}f^{\mathcal{B}}_{D}(d)\log\frac{f^{\mathcal{B}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)\right]$
$=\frac{1}{2}\left[\int_{\mathcal{S}}f^{\mathcal{A}}_{D}(d)\log\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)-\int_{\mathcal{S}}f^{\mathcal{B}}_{D}(d)\log\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\textbf{d}\mu(d)\right]$
$\leq\frac{1}{2}\left[\int_{\mathcal{D}}\left|f^{\mathcal{A}}_{D}(d)-f^{\mathcal{B}}_{D}(d)\right|\left|\log\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right|\textbf{d}\mu(d)\right],$
where the inequality is due to $\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\leq\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}$.
Therefore, we have that
$JS(F^{\mathcal{A}}||F^{\mathcal{B}})\leq\frac{1}{2}\left[\int_{\mathcal{D}}\left|f^{\mathcal{A}}_{D}(d)-f^{\mathcal{B}}_{D}(d)\right|\left|\log\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right|\textbf{d}\mu(d)\right].$
Let $w^{\star}=\arg\max_{w\in\mathcal{W}}f_{D|W}(d|w)$. Then we have that
[…]
$\leq e^{\epsilon_{p,m}}\cdot f_{D}(d)$
$\leq e^{\epsilon_{p,m}+\epsilon}\cdot f^{\mathcal{B}}_{D}(d).$ (8)
Similarly, let $w_{l}=\arg\min_{w\in\mathcal{W}}f_{D|W}(d|w)$. Then we have that
[…]
$\geq e^{-\epsilon_{p,m}}\cdot f_{D}(d)$
$\geq e^{-(\epsilon_{p,m}+\epsilon)}\cdot f^{\mathcal{B}}_{D}(d).$ (9)
Combining Eq. (D) and Eq. (D), we have
$\left|\log\left(\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|\leq\epsilon_{p,m}+\epsilon.$
Now we analyze according to the following two cases.
Case 1: $f^{\mathcal{B}}_{D}(d)\leq f^{\mathcal{A}}_{D}(d)$. Then we have
$\left|\log\left(\frac{f_{D}^{\mathcal{M}}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|=\log\left(\frac{f_{D}^{\mathcal{M}}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\leq\log\left(\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\leq\left|\log\left(\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|\leq\epsilon_{p,m}+\epsilon.$
Case 2: $f^{\mathcal{B}}_{D}(d)\geq f^{\mathcal{A}}_{D}(d)$. Then we have
$\left|\log\left(\frac{f_{D}^{\mathcal{M}}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|=\log\left(\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}^{\mathcal{M}}(d)}\right)\leq\log\left(\frac{f^{\mathcal{B}}_{D}(d)}{f^{\mathcal{A}}_{D}(d)}\right)\leq\left|\log\left(\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|\leq\epsilon_{p,m}+\epsilon.$
Therefore,
$\left|\log\left(\frac{f_{D}^{\mathcal{M}}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|\leq\epsilon_{p,m}+\epsilon.$ (10)
Combining Eq. (D), Eq. (D) and Eq. (14), we have that
$JS(F^{\mathcal{A}}||F^{\mathcal{B}})$ $\leq\frac{1}{2}\left[\int_{\mathcal{S}}\left|f^{\mathcal{A}}_{D}(d)-f^{\mathcal{B}}_{D}(d)\right|\left|\log\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right|\textbf{d}\mu(d)\right]$
$\leq\frac{1}{2}(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right).$
Therefore,
$\epsilon_{p,a}=\sqrt{JS(F^{\mathcal{A}}||F^{\mathcal{B}})}\leq\frac{1}{\sqrt{2}}\sqrt{(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right)}.$
Appendix C Generalized Relationship between LDP and MBP
In this section, we derive the generalized relationship between LDP and MBP.
C-A LDP implies MBP
The following lemma asserts that when the adversary’s prior belief distribution is close to uniform, the mapping function provides a high level of privacy protection by limiting the information leakage to a maximum of $\xi$ in a Bayesian sense when the adversary observes a transformed data point $w$.
Lemma C.1 (LDP implies MBP).
Consider the scenario where the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution. That is, $\forall d,d^{\prime},\ |\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$. If $f_{W|D}(\cdot|\cdot)$ is $\xi$-LDP. That is, the mapping $f_{W|D}(\cdot|\cdot)$ satisfies that
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
then $f_{W|D}(\cdot|\cdot)$ is a privacy-preserving mapping that guarantees $\xi$-maximum Bayesian privacy. That is, $\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|\leq\xi$, $\forall w\in\mathcal{W},d\in\mathcal{D}$.
Proof.
From the definition of $\xi$-LDP, we know that
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
for all $d,d^{\prime}\in\mathcal{D}$. Notice that
$e^{-\xi}\leq\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}\leq e^{\xi}.$
It implies that $f_{D|W}(d|w)\leq f_{D|W}(d^{\prime}|w)\cdot e^{\xi}\cdot\frac{f_{D}(d)}{f_{D}(d^{\prime})}\leq f_{D|W}(d^{\prime}|w)\cdot e^{\xi+\epsilon}$.
We assume that $\frac{f_{D}(d)}{f_{D}(d^{\prime})}\in[e^{-\epsilon},e^{\epsilon}]$. Notice that
$\sum_{d\in\mathcal{D}}f_{D|W}(d|w)=1.$
For any fixed $\tilde{d}\in\mathcal{D}$, we have that
11 11 $=\sum_{d\in\mathcal{D}}f_{D|W}(d|w)$
$=f_{D|W}(\tilde{d}|w)+\sum_{d\in\mathcal{D}\setminus\tilde{d}}f_{D|W}(d|w)$
$\leq f_{D|W}(\tilde{d}|w)+e^{\xi+\epsilon}\cdot f_{D|W}(\tilde{d}|w)\cdot(|\mathcal{D}|-1)$
$=(1+e^{\xi+\epsilon}\cdot(|\mathcal{D}|-1))\cdot f_{D|W}(\tilde{d}|w).$
Therefore, we have that
$f_{D|W}(\tilde{d}|w)\geq\frac{1}{1+e^{\xi+\epsilon}\cdot(|\mathcal{D}|-1)}.$
Similarly, we have that
11 11 $=\sum_{d\in\mathcal{D}}f_{D|W}(d|w)$
$=f_{D|W}(\tilde{d}|w)+\sum_{d\in\mathcal{D}\setminus\tilde{d}}f_{D|W}(d|w)$
$\geq f_{D|W}(\tilde{d}|w)+e^{-\xi-\epsilon}\cdot f_{D|W}(\tilde{d}|w)\cdot(|\mathcal{D}|-1)$
$=(1+e^{-\xi-\epsilon}\cdot(|\mathcal{D}|-1))\cdot f_{D|W}(\tilde{d}|w).$
Therefore, we have
$f_{D|W}(\tilde{d}|w)$ $\leq\frac{1}{1+e^{-\xi-\epsilon}\cdot(|\mathcal{D}|-1)}$
$\leq\frac{e^{\xi+\epsilon}}{|\mathcal{D}|}.$
Therefore, $\forall\tilde{d}\in\mathcal{D}$, we have that
$f_{D|W}(\tilde{d}|w)\leq e^{\xi+\epsilon}f_{D}(\tilde{d}).$
We say a mechanism is $(\xi+\epsilon)$-MBP if $\forall w\in\mathcal{W}^{d}$,
$\frac{f(d|w)}{f(d)}=\frac{f(w|d)f(d)}{f(w)f(d)}=\frac{f(w|d)}{f(w)}\leq e^{\xi+\epsilon}.$ (11)
Therefore, $f_{D|W}(\cdot|\cdot)$ is $(\xi+\epsilon)$-MBP. ∎
C-B MBP implies LDP
The following lemma establishes a connection between MBP and LDP under the assumption that the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution. It states that if a mapping $f_{W|D}(\cdot|\cdot)$ guarantees $\xi$-MBP, then it also satisfies $(2\xi+\epsilon)$-LDP. In other words, if the log-ratio of the posterior belief $f_{D|W}(d|w)$ and the prior belief $f_{D}(d)$ is bounded by $\xi$ for all $w\in\mathcal{W}$ and $d\in\mathcal{D}$, and the prior belief distribution is "$\epsilon$-close" to uniform (i.e., $|\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$ for all $d,d^{\prime}$), then the ratio of the conditional probabilities $f_{W|D}(w|d)$ and $f_{W|D}(w|d^{\prime})$ is bounded within the range $[e^{-2\xi-\epsilon},e^{2\xi+\epsilon}]$ for any $w\in\mathcal{W}$ and $d,d^{\prime}\in\mathcal{D}$. This lemma provides a useful connection between MBP and LDP, which can be leveraged to analyze the privacy guarantees of algorithms satisfying MBP under certain conditions on the prior belief distribution.
Lemma C.2 (MBP implies LDP).
Consider the scenario where the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution. That is, $\forall d,d^{\prime},\ |\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$. Let $f_{W|D}(\cdot|\cdot)$ be a privacy-preserving mapping that guarantees $\xi$-maximum Bayesian privacy. That is, $\xi=\max_{w\in\mathcal{W},d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|$. Then, the mapping $f_{W|D}(\cdot|\cdot)$ is $(2\xi+\epsilon)$-Local Differential Private:
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-2\xi},e^{2\xi}].$
Proof.
From the definition of maximum Bayesian privacy leakage, we know that
$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\xi},e^{\xi}],$
for any $w\in\mathcal{W}$ and $d\in\mathcal{D}$. Therefore, for any $d,d^{\prime}\in\mathcal{D}$ we have
$e^{-2\xi-\epsilon}\leq\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}\cdot e^{-\epsilon}\leq\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\leq\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}\cdot e^{\epsilon}\leq e^{2\xi+\epsilon}.$
From the definition of Differential Privacy, we know that it is $(2\xi+\epsilon)$-Differentially Private. ∎
Combining the results in C-A and C-B, we are now ready to analyze the generalized relationship between MBP and LDP, which is shown in the following theorem.
Theorem C.3 (Relationship between MBP and LDP).
Consider the scenario where the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution. That is, $\forall d,d^{\prime},\ |\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$. If an algorithm is $\xi$-LDP, then it is $(\xi+\epsilon)$-MBP. If an algorithm is $\xi$-maximum Bayesian privacy, then it is $(2\xi+\epsilon)$-LDP.
Appendix D Generalized Relationship Between Average Bayesian Privacy and Maximum Bayesian Privacy
In this section, we introduce the generalized relationship between average Bayesian privacy and maximum Bayesian privacy.
Theorem D.1 (Relationship Between MBP and ABP).
Assume that the prior belief of the attacker satisfies that
$\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}(d)}\in[e^{-\epsilon},e^{\epsilon}].$
Let MBP be defined using Definition IV.3. For any $\epsilon_{p,m}\geq 0$, let $f_{W|D}(\cdot|\cdot)$ be a privacy preserving mapping that guarantees $\epsilon_{p,m}$-maximum Bayesian privacy. That is,
$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\epsilon_{p,m}},e^{\epsilon_{p,m}}],$
for any $w\in\mathcal{W}$. Then, we have that the ABP (denoted as $\epsilon_{p,a}$) is bounded by
$\epsilon_{p,a}\leq\frac{1}{\sqrt{2}}\sqrt{(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right)},$
where $\epsilon_{p,a}$ is introduced in Definition IV.1.
Proof.
Let $F^{\mathcal{M}}=\frac{1}{2}(F^{\mathcal{A}}+F^{\mathcal{B}})$. We have
$JS(F^{\mathcal{A}}||F^{\mathcal{B}})$ $=\frac{1}{2}\left[KL\left(F^{\mathcal{A}}||F^{\mathcal{M}}\right)+KL\left(F^{\mathcal{B}}||F^{\mathcal{M}}\right)\right]$
$=\frac{1}{2}\left[\int_{\mathcal{S}}f^{\mathcal{A}}_{D}(d)\log\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)+\int_{\mathcal{S}}f^{\mathcal{B}}_{D}(d)\log\frac{f^{\mathcal{B}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)\right]$
$=\frac{1}{2}\left[\int_{\mathcal{S}}f^{\mathcal{A}}_{D}(d)\log\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)-\int_{\mathcal{S}}f^{\mathcal{B}}_{D}(d)\log\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\textbf{d}\mu(d)\right]$
$\leq\frac{1}{2}\left[\int_{\mathcal{D}}\left|f^{\mathcal{A}}_{D}(d)-f^{\mathcal{B}}_{D}(d)\right|\left|\log\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right|\textbf{d}\mu(d)\right],$
where the inequality is due to $\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\leq\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}$.
Therefore, we have that
$JS(F^{\mathcal{A}}||F^{\mathcal{B}})\leq\frac{1}{2}\left[\int_{\mathcal{D}}\left|f^{\mathcal{A}}_{D}(d)-f^{\mathcal{B}}_{D}(d)\right|\left|\log\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right|\textbf{d}\mu(d)\right].$
Let $w^{\star}=\arg\max_{w\in\mathcal{W}}f_{D|W}(d|w)$. Then we have that
[…]
$\leq e^{\epsilon_{p,m}}\cdot f_{D}(d)$
$\leq e^{\epsilon_{p,m}+\epsilon}\cdot f^{\mathcal{B}}_{D}(d).$ (12)
Similarly, let $w_{l}=\arg\min_{w\in\mathcal{W}}f_{D|W}(d|w)$. Then we have that
[…]
$\geq e^{-\epsilon_{p,m}}\cdot f_{D}(d)$
$\geq e^{-(\epsilon_{p,m}+\epsilon)}\cdot f^{\mathcal{B}}_{D}(d).$ (13)
Combining Eq. (D) and Eq. (D), we have
$\left|\log\left(\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|\leq\epsilon_{p,m}+\epsilon.$
Now we analyze according to the following two cases.
Case 1: $f^{\mathcal{B}}_{D}(d)\leq f^{\mathcal{A}}_{D}(d)$. Then we have
$\left|\log\left(\frac{f_{D}^{\mathcal{M}}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|=\log\left(\frac{f_{D}^{\mathcal{M}}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\leq\log\left(\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\leq\left|\log\left(\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|\leq\epsilon_{p,m}+\epsilon.$
Case 2: $f^{\mathcal{B}}_{D}(d)\geq f^{\mathcal{A}}_{D}(d)$. Then we have
$\left|\log\left(\frac{f_{D}^{\mathcal{M}}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|=\log\left(\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}^{\mathcal{M}}(d)}\right)\leq\log\left(\frac{f^{\mathcal{B}}_{D}(d)}{f^{\mathcal{A}}_{D}(d)}\right)\leq\left|\log\left(\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|\leq\epsilon_{p,m}+\epsilon.$
Therefore,
$\left|\log\left(\frac{f_{D}^{\mathcal{M}}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|\leq\epsilon_{p,m}+\epsilon.$ (14)
Combining Eq. (D), Eq. (D) and Eq. (14), we have that
$JS(F^{\mathcal{A}}||F^{\mathcal{B}})$ $\leq\frac{1}{2}\left[\int_{\mathcal{S}}\left|f^{\mathcal{A}}_{D}(d)-f^{\mathcal{B}}_{D}(d)\right|\left|\log\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right|\textbf{d}\mu(d)\right]$
$\leq\frac{1}{2}(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right).$
Therefore,
$\epsilon_{p,a}=\sqrt{JS(F^{\mathcal{A}}||F^{\mathcal{B}})}\leq\frac{1}{\sqrt{2}}\sqrt{(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right)}.$
Appendix E Analysis for Theorem IV.7
Theorem E.1 (PAC Robustness of $\epsilon$-MBP Protection Mechanisms).
Let $\mathcal{M}$ be an $(\epsilon,\beta,\alpha)$-MBP (Definition IV.5) algorithm. Then $\mathcal{M}$ is $\alpha$-PAC-Robust with probability $1-\gamma$, where $\gamma\in(0,1)$, $n$ represents the size of the dataset, and $\beta=\frac{\gamma}{1+4\epsilon}$ represents the error probability.
Proof.
$\mathcal{M}$ is an $(\epsilon,\beta,\alpha)$-MBP algorithm for the estimation of statistic $\mu$. We know that the protection mechanism $\mathcal{M}$ is $\epsilon$-MBP (measured using Definition IV.3). From the relationship between MBP and LDP (Theorem IV.4), we know that $\mathcal{M}$ is $2\epsilon$-LDP. From the definition of private algorithm, we know that for the original private information $d$, we have that
$\|\mathcal{M}(d)-\mu(P)\|\leq\alpha.$ (15)
Denote $\mathcal{U}=\{t\in\mathbb{R}^{m}:\|t-\mu\|>\alpha\}$. Then, we have that
$\Pr[\mathcal{M}(d)\in\mathcal{U}]$ $=\Pr[\|\mathcal{M}(d)-\mu(P)\|>\alpha]$ (16)
$\leq\beta.$ (17)
For any $d^{\prime}\neq d$, we have that
$\Pr[\|\mathcal{M}(d^{\prime})-\mu(P)\|>\alpha]$ $\leq e^{2\epsilon}\Pr[\|\mathcal{M}(d)-\mu(P)\|>\alpha]$ (18)
$\leq e^{2\epsilon}\beta$ (19)
$\leq(1+2\epsilon+4\epsilon^{2})\beta$ (20)
$\leq(1+4\epsilon)\beta$ (21)
$\leq\gamma$ (22)
where the first inequality is due to the definition of local differential privacy, the second inequality is due to Eq. (16), the third and fourth inequalities are due to $0\leq\epsilon\leq 1$, and the last inequality is due to the relationship between $\beta$ and $\epsilon$.
From Definition IV.6, $\mathcal{M}$ is $\alpha$-PAC-Robust with probability $1-\gamma$.
Appendix F Practical Estimation for Maximum Bayesian Privacy
$\xi$ represents the maximum privacy leakage over all possible released information $w$ from the set $\mathcal{W}$ and all the dataset $d$, expressed as
$\xi=\max_{w\in\mathcal{W},d}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|.$ (23)
When the extent of the attack is fixed, $\xi$ becomes a constant. Given Assumption G.1, we can infer that
$\xi$ $=\max_{d}\left|\log\left(\frac{f_{D|W}(d|w^{*})}{f_{D}(d)}\right)\right|.$ (24)
We approximate $\xi$ by the following estimator:
$\hat{\xi}=\max_{d}\left|\log\left(\frac{\hat{f}_{D|W}(d|w^{*})}{f_{D}(d)}\right)\right|.$ (25)
We define $\kappa_{3}$ as the maximum Bayesian Privacy $\xi$, and let $\hat{\kappa}_{3}$ denote its estimate $\hat{\xi}$
$\hat{\kappa}_{3}=\hat{\xi}.$ (26)
The task of estimating the maximum Bayesian Privacy $\xi$ reduces to determining the value of $\hat{f}_{D|W}(d|w_{m})$, which will be addressed in the subsequent section.
F-A Estimation for $\hat{f}_{{D}|{W}}(d|w_{m})$
We first generate a set of models $\{w_{m}\}_{m=1}^{M}$ stochastically: to generate the $m$-th model $w_{m}$, we run an algorithm (e.g., stochastic gradient descent), using a mini-batch $\mathcal{S}$ randomly sampled from $\mathcal{D}=\{z_{1},\ldots,z_{|\mathcal{D}|}\}$.
The set of parameters $\{w_{m}\}_{m=1}^{M}$ is generated for a given dataset $\mathcal{D}$ by randomly initializing the parameters for each model, iterating a specified number of times, and updating the parameters with mini-batches sampled from the dataset using approaches such as gradient descent. After the iterations, the final set of parameters for each model is returned as $\{w_{m}\}_{m=1}^{M}$.
We then estimate $\hat{f}_{{D}|{W}}(d|w_{m})$ with $\{w_{m}\}_{m=1}^{M}$. As mentioned in subsection III-A, the attacker can conduct Bayesian inference attack, i.e., estimate dataset $d$ with $\hat{d}$ that optimizes posterior belief, denoted by:
$\hat{d}=\arg\max_{d}\text{log}f_{D|W}(d|w)=\arg\max_{d}\text{log}[f_{W|D}(w|d)\times f^{\mathcal{B}}_{D}(d)],$ (27)
Empirically, $\hat{f}^{\mathcal{O}}_{D}(d)$ can be estimated by Eq. (30): $\hat{f}^{\mathcal{O}}_{D}(d)=\frac{1}{M}\sum_{m=1}^{M}\hat{f}_{{D}|{W}}(d|w_{m})$. Then We determine the success of attack on an image using a function $\Omega$, which measures the similarity between the distribution of recovered dataset and the original one (e.g. $\Omega$ can be Kullback–Leibler divergence or Cross Entropy). We consider the original dataset $d$ to be successfully recovered if $\Omega(\hat{f}_{D}^{\mathcal{O}}(d),f_{D}(d))>t$, where $t$, a pre-given constant, is the threshold for determining successful recovery of the original data.
This algorithm is used to estimate the conditional probability $f_{{D}|{W}}(d|w)$. It takes a mini-batch of data $d$ from client $k$, with batch size $S$, and a set of parameters $\{w_{m}\}_{m=1}^{M}$ as input. The goal of the algorithm is to estimate the conditional probability $f_{{D}|{W}}(d|w)$ by iteratively updating the parameters $w_{m}$.
The key steps of the algorithm are as follows:
  • Compute the gradients $\nabla w_{m}$ on each parameter $w_{m}$, which is done by calculating the gradient of the loss function $\mathcal{L}$ with respect to the parameter $w_{m}$. This is performed using samples from the mini-batch data $d$. Then update the parameters $w_{m}$ using gradient descent, where $w_{m}^{\prime}$ represents the updated parameters obtained by subtracting the product of the learning rate $\eta$ and the gradient $\nabla w_{m}$ from $w_{m}$.
  • For each parameter $w_{m}$, iterate the following steps for $T$ times: Generate perturbed data $\tilde{d}$ by applying the Backward Iterative Alignment (BIA) method with the parameters $w_{m}^{\prime}$ and the gradient $\nabla w_{m}$.
  • Determine whether the original data $z_{d}$ is recovered by comparing it with the perturbed data $\tilde{z}_{d}$, based on a defined threshold $\Omega$ that measures the difference between them. Accumulate the count of recovered instances for each data sample $z_{d}$ in the variable $M_{d}^{m}$.
  • Estimate the conditional probability $\hat{f}_{{D}|{W}}(d|w_{m})$ for each data sample $z_{d}$ in the dataset $d$ by dividing $M_{d}^{m}$ by the product of the total number of iterations $T$ and the batch size $S$, i.e., $\bm{\frac{M_{d}^{m}}{S\cdot T}}$. Return the estimated conditional probabilities $\hat{f}_{{D}|{W}}(d|w_{m})$ for each parameter $w_{m}$.
The core idea of this algorithm is to estimate the conditional probability by generating perturbed data and comparing it with the original data during the parameter update process. By performing multiple iterations and accumulating counts, the algorithm provides an estimation of the conditional probability.
F-B Estimation Error of MBP
To streamline the error bound analysis, it is presumed that both the data and model parameters take on discrete values.
The following lemma states that the estimated value $\hat{\kappa}_{1}(d)$ is very likely to be close to the true value $\kappa_{1}(d)$. The probability of the estimate being within an error margin that is a small percentage ($\epsilon$ times) of the true value is very high—specifically, at least $1-2\exp(-\frac{\epsilon^{2}T\kappa_{1}(d)}{3})$. This means that the larger the number $T$ or the smaller the error parameter $\epsilon$, the more confident we are that our estimate is accurate. For the detailed analysis, please refer to Zhang et al. (2023e).
Lemma F.1.
With probability at least $1-2\exp\left(\frac{-\epsilon^{2}T\kappa_{1}(d)}{3}\right)$, we have that
$|\hat{\kappa}_{1}(d)-\kappa_{1}(d)|\leq\epsilon\kappa_{1}(d).$ (28)
Appendix G Practical Estimation for Average Bayesian Privacy
In the scenario of federated learning, Zhang et al. (2023e) derived the upper bound of privacy leakage based on the specific form of $\text{TV}(P^{\mathcal{O}}||P^{\mathcal{D}})$ of $\mathcal{M}$ according to this equation: $\widetilde{\epsilon}_{p}=2C_{1}-C_{2}\cdot{\text{TV}}(P^{\mathcal{O}}||P^{\mathcal{D}})$. In this section, we elaborate on the methods we propose to estimate $C_{1}$ and $C_{2}$. First, we introduce the following assumption.
Assumption G.1.
We assume that the amount of Bayesian privacy leaked from the true parameter is maximum over all possible parameters. That is, for any $w\in\mathcal{W}$, we have that
$\frac{f_{D|W}(d|w^{*})}{f_{D}(d)}\geq\frac{f_{D|W}(d|w)}{f_{D}(d)},$ (29)
where $w^{*}$ represents the true model parameter.
G-A Estimation for $C_{1}=\sqrt{{\text{JS}}(F^{\mathcal{O}}||F^{\mathcal{B}})}$
$C_{1}$ quantifies the average square root of the Jensen-Shannon divergence between the adversary’s belief distribution about the private information before and after observing the unprotected parameter, independent of the employed protection mechanisms. The JS divergence is defined as:
$\text{JS}(F^{\mathcal{O}}\parallel F^{\mathcal{B}})=\frac{1}{2}\left(\int_{\mathcal{D}}f^{\mathcal{O}}_{D}(d)\log\frac{f^{\mathcal{O}}_{D}(d)}{\frac{1}{2}(f^{\mathcal{O}}_{D}(d)+f^{\mathcal{B}}_{D}(d))}\,\textbf{d}\mu(d)+\int_{\mathcal{D}}f_{D}(d)\log\frac{f_{D}(d)}{\frac{1}{2}(f^{\mathcal{O}}_{D}(d)+f^{\mathcal{B}}_{D}(d))}\,\textbf{d}\mu(d)\right),$
[…]
To estimate $C_{1}$, we first need to estimate the values of $f^{\mathcal{O}}_{D}(d)$ and $f^{\mathcal{B}}_{D}(d)$. It is approximated that
$\hat{f}^{\mathcal{O}}_{D}(d)=\frac{1}{M}\sum_{m=1}^{M}\hat{f}_{{D}|{W}}(d|w_{m}),$ (30)
where $w_{m}$ signifies the model parameter observed by the attacker at the $m$-th attempt to infer data $d$ given $w_{m}$.
Denoting $C_{1}=\sqrt{\text{JS}(F^{\mathcal{O}}\parallel F^{\mathcal{B}})}$, and letting $\kappa_{1}=f^{\mathcal{O}}_{D}(d)$, $\kappa_{2}=f^{\mathcal{B}}_{D}(d)=f_{D}(d)$, and $\hat{\kappa}_{1}=\hat{f}^{\mathcal{O}}_{D}(d)$, we have
$C_{1}^{2}=\frac{1}{2}\int_{\mathcal{D}}\kappa_{1}\log\frac{\kappa_{1}}{\frac{1}{2}(\kappa_{1}+\kappa_{2})}\,d\mu(d)+\frac{1}{2}\int_{\mathcal{D}}\kappa_{2}\log\frac{\kappa_{2}}{\frac{1}{2}(\kappa_{1}+\kappa_{2})}\,d\mu(d).$
With these estimates, the squared estimated value of $C$ becomes
$\hat{C}_{1}^{2}=\frac{1}{2}\int_{\mathcal{D}}\hat{\kappa}_{1}\log\frac{\hat{\kappa}_{1}}{\frac{1}{2}(\hat{\kappa}_{1}+\kappa_{2})}\,d\mu(d)+\frac{1}{2}\int_{\mathcal{D}}\kappa_{2}\log\frac{\kappa_{2}}{\frac{1}{2}(\hat{\kappa}_{1}+\kappa_{2})}\,d\mu(d).$
G-B Estimation for $C_{2}=\frac{1}{2}(e^{2\xi}-1)$
Recall that the Maximum Bayesian Privacy, is defined as $\xi=\max_{w\in\mathcal{W},d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|$ and represents the maximal privacy leakage across all possible information $w$ released. The quantity $\xi$ remains constant when the extent of the attack is fixed. Under Assumption G.1, we have that
$\xi=\max_{d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w^{*})}{f_{D}(d)}\right)\right|.$ (31)
The approximation of $\xi$ is given by:
$\hat{\xi}=\max_{d\in\mathcal{D}}\left|\log\left(\frac{\hat{f}_{D|W}(d|w^{*})}{f_{D}(d)}\right)\right|.$ (32)
Letting $\kappa_{3}=\xi$ and $\hat{\kappa}_{3}=\hat{\xi}$, we compute $C_{2}$ as:
$\hat{C}_{2}=\frac{1}{2}\left(e^{2\hat{\kappa}_{3}}-1\right).$ (33)
Ultimately, the computation of $C_{1}$ and $C_{2}$ hinges on the estimation of $\hat{f}_{{D}|{W}}(d|w_{m})$, which is further elaborated upon in Section F-A.
G-C Estimation Error of ABP
In this section, we analyze the estimation error of ABP. To facilitate the analysis of the error bound, we consider that both the data and the model parameter are in discrete form. For the detailed analysis, please refer to Zhang et al. (2023e).
The following theorem establishes that a bounded estimation error for $\kappa_{1}(d)$—specifically, not exceeding $\epsilon$ times the value of $\kappa_{1}(d)$—implies a bounded estimation error for another variable, $C_{1}$. This bound for $C_{1}$ is defined by a specific formula incorporating $\epsilon$ and its logarithmic terms. In summary, the theorem provides a calculated limit for how large the estimation error for $C_{1}$ can be, given the known estimation error for $\kappa_{1}(d)$.
Theorem G.1 (The Estimation Error of $C_{1}$).
Assume that $|\kappa_{1}(d)-\hat{\kappa}_{1}(d)|\leq\epsilon\kappa_{1}(d)$. We have that
$|\hat{C}_{1}-C_{1}|\leq\sqrt{\frac{(1+\epsilon)\log\frac{1+\epsilon}{1-\epsilon}}{2}+\frac{\epsilon+\max\left\{\log(1+\epsilon),\log\frac{1}{(1-\epsilon)}\right\}}{2}}.$ (34)
The analysis for the estimation error of $C_{2}$ depends on the application scenario. To facilitate the analysis, we assume that $C_{2}$ is known in advance. With the estimators and estimation bound for $C_{1}$, we are now ready to derive the estimation error for privacy leakage.
The following theorem addresses the accuracy of an estimated privacy parameter, $\hat{\epsilon}_{p}$, in comparison to its true value, $\epsilon_{p}$. The theorem provides a mathematical limit on how much the estimated value $\hat{\epsilon}_{p}$ can differ from the true privacy parameter $\epsilon_{p}$. The difference between them is not to exceed a certain threshold, which is a complex expression involving the privacy parameter $\epsilon$ itself and logarithmic terms. This threshold is important in applications where it is crucial to know how accurately the privacy parameter can be estimated.
Theorem G.2.
We have that
$|\hat{\epsilon}_{p}-\epsilon_{p}|\leq\frac{3}{2}\cdot\sqrt{\frac{(1+\epsilon)\log\frac{1+\epsilon}{1-\epsilon}}{2}+\frac{\epsilon+\max\left\{\log(1+\epsilon),\log\frac{1}{(1-\epsilon)}\right\}}{2}}.$ (35)

### v5

@@
 Abstract
-The advent of machine learning has led to transformative changes across various domains, but the sensitive nature of data raises concerns about privacy and security.
-While Local Differential Privacy (LDP) has been a cornerstone in addressing these concerns, recent research has proposed privacy concepts aligned with the Bayesian inference perspective of an adversary, such as Average Bayesian Privacy (ABP) and Maximum Bayesian Privacy (MBP).
-This paper explores the intricate relationships between LDP, ABP, and MBP, and their implications for algorithmic robustness.
-We establish theoretical connections between these privacy notions, proving that LDP implies MBP and vice versa under certain conditions, and deriving bounds connecting MBP and ABP.
-We also investigate the relationship between PAC robust learning and privacy preservation, demonstrating how to derive PAC robustness from privacy-preserving algorithms and construct privacy-preserving algorithms from PAC robust ones.
-Our findings provide valuable insights for constructing privacy-preserving and robust machine learning algorithms.
-Index Terms:
-Privacy-Preserving Algorithms; PAC Robust Learning; Trustworthy Machine Learning; Theoretical Connections
-I Introduction
-The advent of machine learning has heralded transformative changes across various domains, from healthcare to finance, by extracting meaningful insights from vast swathes of data.
-However, this data often contains sensitive information, raising legitimate concerns regarding privacy and security.
-The traditional paradigm of Local Differential Privacy (LDP) has been a cornerstone in addressing these concerns, aiming to ensure that the influence of any individual data point on the output of a computation is bounded, thereby obfuscating its presence or absence (Dwork, 2008).
-While LDP provides a strong foundation for privacy protection, it is essential to consider the adversary’s methods for potentially compromising privacy.
-Recent research has proposed privacy concepts aligned with the Bayesian inference perspective of an adversary (He et al., 2019; Zhu et al., 2019), such as Average Bayesian Privacy (ABP) and Maximum Bayesian Privacy (MBP).
-These concepts have emerged from an understanding that adversaries may possess prior knowledge and use their Bayesian reasoning capabilities to infer sensitive information from the output of machine learning algorithms.
-By incorporating the adversary’s perspective, these new paradigms offer a more comprehensive approach to privacy protection, taking into account the potential strategies an adversary might employ to infer sensitive information.
-As these Bayesian privacy concepts gain traction, it is crucial to thoroughly investigate their relationships with traditional privacy measures like LDP.
-Understanding the connections and differences between these approaches will enable researchers and practitioners to develop more robust and effective privacy-preserving mechanisms.
-By leveraging the strengths of both LDP and Bayesian privacy concepts, we can work towards building machine learning systems that safeguard sensitive information while still extracting valuable insights from data.
-In this paper, we aim to explore the intricate relationships between LDP, ABP, and MBP, providing a deeper understanding of how these privacy notions interact and complement each other.
-Through rigorous mathematical analysis and empirical studies, we seek to uncover the key factors that influence the effectiveness of these privacy protection mechanisms and provide guidance on how to design and implement them in real-world machine learning applications.
-By bridging the gap between traditional and Bayesian privacy concepts, we hope to contribute to the development of more secure and trustworthy machine learning systems that respect individual privacy while unlocking the full potential of data-driven insights.
-Through meticulous definitions and theorems, we elucidate how these privacy concepts are interconnected.
-A key result (Theorem IV.4) establishes that, under certain conditions, LDP implies MBP and vice versa.
-We also derive bounds connecting MBP and ABP (Theorem IV.6), providing valuable insights for constructing privacy-preserving algorithms.
-The intersection of privacy and learning theory, particularly Probably Approximately Correct (PAC) learning, has attracted considerable scholarly inquiry.
-Groundbreaking research by Kasiviswanathan et al.
-(2011) revealed that a vast array of classes learnable by non-private algorithms can also be mastered privately, albeit with a polynomial increase in sample size.
-Building upon these foundational intersections, our work extends the discourse by delving into the robustness implications of privacy guarantees.
-Robustness in machine learning, the ability of algorithms to maintain performance despite inaccuracies or noise in the input data, has become increasingly important as models are deployed in unpredictable real-world environments.
-The concept of robustness has been traditionally studied within the PAC framework (Cullina et al., 2018), which provides a theoretical foundation for the performance of learning algorithms under perturbations.
-However, while previous studies have predominantly focused on sample complexity and learnability, the intersection of robustness with privacy has not been fully explored, particularly how robust learning can contribute to, or benefit from, privacy-preserving mechanisms.
-In this paper, we take a deeper dive into how privacy-preserving mechanisms can enhance robustness in PAC learning scenarios, and conversely, how robust learning algorithms can inform the design of privacy-preserving mechanisms.
-We investigate the relationship between PAC robust learning and privacy preservation, demonstrating how to derive PAC robustness from algorithms satisfying specific privacy standards (Theorem IV.7).
-This exploration aims to bridge the gap between robustness and privacy, providing valuable insights for developing machine learning algorithms that are both resilient to perturbations and secure against privacy breaches.
-The main contributions of this paper are as follows:
-  • We establish theoretical connections between different privacy notions, proving that LDP implies MBP and vice versa under certain conditions (Theorem IV.4) and deriving relationship between MBP and ABP (Theorem IV.6).
-  • We investigate the relationship between PAC robust learning and privacy preservation, demonstrating how to derive PAC robustness from privacy-preserving algorithms (Theorem IV.7) and construct privacy-preserving algorithms from PAC robust ones.
+The widespread adoption of machine learning necessitates robust privacy protection alongside algorithmic resilience.
+While Local Differential Privacy (LDP) provides foundational guarantees, sophisticated adversaries with prior knowledge demand more nuanced Bayesian privacy notions, such as Maximum Bayesian Privacy (MBP) and Average Bayesian Privacy (ABP), first introduced by [61].
+Concurrently, machine learning systems require inherent robustness against data perturbations and adversarial manipulations.
+This paper systematically investigates the intricate theoretical relationships among LDP, MBP, and ABP.
+Crucially, we bridge these privacy concepts with algorithmic robustness, particularly within the Probably Approximately Correct (PAC) learning framework.
+Our work demonstrates that privacy-preserving mechanisms inherently confer PAC robustness.
+We present key theoretical results, including the formalization of the established LDP-MBP relationship, novel bounds between MBP and ABP, and a proof demonstrating PAC robustness from MBP.
+Furthermore, we establish a novel theoretical relationship quantifying how privacy leakage directly influences an algorithm’s input robustness.
+These results provide a unified theoretical framework for understanding and optimizing the privacy-robustness trade-off, paving the way for the development of more secure, trustworthy, and resilient machine learning systems.
+keywords:
+federated learning , privacy-utility tradeoffs, privacy-preserved machine learning
+1 Introduction
+The advent of machine learning (ML) has ushered in transformative changes across diverse domains—from healthcare diagnostics to financial modeling and autonomous systems—driven by its unprecedented capability to extract meaningful insights from complex datasets [4, 31].
+However, this pervasive reliance on data, particularly sensitive information, raises critical concerns regarding privacy and security [51].
+Simultaneously, as ML models are increasingly deployed in dynamic real-world environments, their reliability under input inaccuracies or noise—quantified as algorithmic robustness—has become paramount [50, 55].
+Building trustworthy, deployable AI systems necessitates addressing both challenges holistically.
+The traditional paradigm of Local Differential Privacy (LDP) provides a cornerstone for privacy preservation, bounding the influence of any individual data point on computational outputs to obscure its presence [21].
+Particularly valuable in distributed settings where users perturb data locally, LDP offers strong theoretical guarantees.
+Yet sophisticated adversaries may exploit additional attack vectors.
+Recent research introduces Bayesian privacy frameworks—Average Bayesian Privacy (ABP) and Maximum Bayesian Privacy (MBP)—that model adversaries with prior knowledge and Bayesian inference capabilities [61].
+These frameworks deliver more comprehensive protection by accounting for realistic adversarial strategies.
+These critical domains have largely evolved independently of algorithmic robustness research, creating substantial theoretical and practical gaps: (i) a lack of unified principles connecting nuanced privacy guarantees with robustness, and (ii) mechanisms that optimize one property at the expense of the other, while yielding suboptimal trade-offs.
+Bridging these gaps is essential for establishing trustworthy next-generation AI foundations.
+Our work addresses these challenges through fundamental theoretical innovations that establish quantifiable synergies between privacy preservation and algorithmic robustness.
+Our contributions are threefold:
+  • First, we derive the inaugural quantitative relationship between MBP and ABP (Theorem 4.2).
+This result precisely bounds expected privacy leakage (ABP) as a function of worst-case protection (MBP), demonstrating that strengthening MBP inherently enhances ABP—providing crucial mechanism-design guidance.
+  • Second, we formally establish that privacy-preserving mechanisms inherently confer Probably Approximately Correct (PAC) robustness (Theorem 4.3).
+This profound equivalence proves an $(\epsilon,\beta,\alpha)$-MBP mechanism guarantees $\alpha$-PAC-Robustness, with quantifiable confidence degradation $\propto\epsilon\beta$, thereby bridging privacy and stability domains.
+  • Third, we reveal a novel mathematical duality between privacy leakage and input robustness (Theorem 5.5).
+Our analysis demonstrates that intentional parameter distortion ($\Delta$) from privacy mechanisms directly amplifies adversarial resilience ($\alpha$), enabling principled joint optimization via strategic noise injection.
+Collectively, these contributions establish a theoretical foundation for developing ML systems that simultaneously resist privacy reconstruction attacks and maintain performance under data perturbations.
+Our framework delivers: (1) quantitative conversions among Bayesian privacy metrics, (2) formal privacy-to-robustness implications, and (3) pathways for unified optimization.
 The remainder of this paper is organized as follows.
-Section II-A reviews related work on privacy measurements, Section II-D discusses robustness measurements, and Section II-E explores the relationship between privacy and robustness in the literature.
-In Section III, we introduce our privacy attack and defense framework.
-In Section IV, we define metrics for privacy and robustness, and establish theoretical connections between privacy notions and PAC robustness.
-We also introduce practical estimation approaches and analyses for MBP and ABP in F and G.
+Section 2 reviews related work on privacy measurements, relationships among privacy metrics, intersections of privacy and PAC learning, robustness measurements, and the relationship between privacy and robustness.
+In Section 3, we introduce our privacy attack and defense framework.
+In Section 4, we define metrics for privacy and robustness, and establish theoretical connections between privacy notions and PAC robustness.
+We also introduce practical estimation approaches and analyses for MBP and ABP in D and E.
 Finally, we conclude the paper and discuss future research directions.
 For convenience, we list the important notations used throughout this paper in Table LABEL:table:_notation.
-TABLE I: Table of Notation
+Table 1: Table of Notation
 Notation Meaning
@@
 $\hat{\kappa_{3}}$ estimator of $\kappa_{3}$
-II Related Work
-II-A Privacy Measurements
-Differential privacy (DP), a rigorously formalized privacy framework pioneered by Dwork et al.
-(2006), has been widely adopted to protect individual privacy in database systems.
-Abadi et al.
-(2016) proposed a method for training neural networks with a limited privacy budget, combining advanced machine learning methods and privacy-preserving mechanisms.
-This method is applicable to various optimization methods, including Momentum Rumelhart et al.
-(1986) and AdaGrad Duchi et al.
-(2011).
-Triastcyn and Faltings (2020) introduced Bayesian Differential Privacy (BDP), an adaptation of DP that incorporates data distribution considerations, along with a theoretical analysis and privacy computing framework, demonstrating its advantages through experimental results.
-Rassouli and Gündüz (2019) employed average total variation distance (TV distance) for privacy measurement and minimum mean-square error (MMSE), mutual information, and error probability to assess utility, aiming to find the best balance between utility and privacy in data analysis tasks.
-Blauw and von Solms (2017) proposed two main privacy metrics based on an information-theoretic approach: information entropy and mutual information, to quantify and define online user privacy.
-Achieving pure DP can be challenging in practical learning scenarios.
-Eilat et al.
-(2020) quantified the loss of privacy using relative entropy (Kullback-Leibler divergence) and defined the concept of privacy loss, referring to the information leakage caused by the posterior update of the agent’s prior beliefs.
-Foulds et al.
-(2016) discussed the constraints of obtaining privacy via posterior sampling and highlighted the benefits of the Laplace mechanism in Bayesian inference from both theoretical and empirical perspectives.
-Other works have considered privacy measures akin to our approach.
-In social network settings, Alemany et al.
-(2019) proposed two metrics (Audience and Reachability) for the privacy risk of sharing information.
-du Pin Calmon and Fawaz (2012) discussed the minimization of maximum information leakage in privacy-preserving mapping, proposing a mathematical model and a formula that minimizes the maximum information leakage, intended to balance privacy against utility constraints, with privacy leakage assessed using KL divergence.
-In contrast, our work adopts Jensen-Shannon (JS) divergence for measuring privacy leakage, in line with Zhang et al.
-(2022, 2023d), due to its symmetric nature and the fact that its square root meets the triangle inequality Nielsen (2019); Endres and Schindelin (2003), aiding in quantifying privacy-utility trade-offs more effectively.
-II-B Relation between Distinct Privacy Measurements
-The advent of differential privacy has inspired numerous scholars to propose new privacy metrics and explore their relationships.
-Table III summarizes key relations between various privacy metrics, including our proposed Maximum Bayesian Privacy (MBP) and Average Bayesian Privacy (ABP).
-TABLE II: Relations between various privacy metrics.
+2 Related Work
+2.1 Privacy Measurements
+The rigorous formalization of differential privacy (DP) by [23] established a gold standard for privacy-preserving data analysis, catalyzing extensive research into privacy quantification.
+Subsequent works have expanded this foundation along several dimensions: computational feasibility, distributional awareness, and practical applicability.
+[2]’s pioneering work demonstrated how deep neural networks could be trained under constrained privacy budgets by integrating DP-SGD with optimization techniques like Momentum [49] and AdaGrad [19].
+This approach balanced the tension between model utility and privacy preservation, enabling privacy-sensitive deep learning applications.
+Bayesian perspectives introduced significant advancements in privacy measurement.
+[52] developed Bayesian Differential Privacy (BDP), incorporating data distribution considerations into the privacy calculus while establishing formal connections to traditional DP guarantees.
+Information-theoretic approaches emerged as powerful alternatives, with [47] employing total variation distance for privacy measurement while assessing utility through minimum mean-square error, mutual information, and error probability.
+Similarly, [9] proposed entropy-based metrics for quantifying online user privacy, acknowledging the contextual nature of privacy risks.
+The practical challenges of achieving pure DP motivated alternative quantification frameworks.
+[25] formalized privacy loss through relative entropy (Kullback-Leibler divergence), measuring information leakage via posterior belief updates.
+While [28] identified limitations in posterior sampling for privacy protection, they demonstrated the effectiveness of Laplace mechanisms in Bayesian inference contexts.
+Social network privacy inspired novel metrics like Audience and Reachability [5], while [18] pioneered minimax formulations for information leakage using KL divergence.
+Our work contributes to this landscape by adopting Jensen-Shannon (JS) divergence for privacy measurement, following [61, 63].
+This choice is theoretically grounded in JS divergence’s symmetry and triangle inequality properties [44, 26], which enable more nuanced quantification of privacy-utility trade-offs.
+This approach provides distinct advantages for analyzing the privacy-robustness relationships central to our framework.
+2.2 Relation between Distinct Privacy Measurements
+The evolution of differential privacy has spurred development of nuanced privacy metrics with complex interrelationships, as summarized in Table 3.
+Several significant conceptual frameworks have emerged:
+Table 2: Taxonomy of privacy metrics and their theoretical relationships
 Reference Privacy Metric1 Relation To Other Metrics
-Dwork and Rothblum (2016) Concentrated DP DP $\Rightarrow$ Concentrated DP
-Mironov (2017) Rényi-DP Rényi-DP $\Rightarrow$ $(\epsilon,\delta)$-DP
-Mironov et al. (2009)2 IND-CDP, SIM-CDP SIM-CDP $\Rightarrow$ SIM∀∃-CDP3 $\Rightarrow$ IND-CDP
-Dong et al. (2019) $f$-DP Gaussian DP4 $\Rightarrow$ Rényi-DP
-Blum et al. (2008) Distributional Privacy Distributional Privacy $\Rightarrow$ DP
-Triastcyn and Faltings (2020) Bayesian-DP Bayesian DP $\Rightarrow$ DP
-Cuff and Yu (2016) Mutual-Information DP MI-DP $\Leftrightarrow$ DP (asymptotically)
-Our Work $\xi$-MBP, $\epsilon$-ABP $\xi$-LDP $\Rightarrow$ $\xi$-MBP $\Rightarrow$ $2\xi$-LDP, MBP $\Rightarrow$ ABP
+[24] Concentrated DP DP $\Rightarrow$ Concentrated DP
+[42] Rényi-DP Rényi-DP $\Rightarrow$ $(\epsilon,\delta)$-DP
+[43]2 IND-CDP, SIM-CDP SIM-CDP $\Rightarrow$ SIM∀∃-CDP3 $\Rightarrow$ IND-CDP
+[17] $f$-DP Gaussian DP4 $\Rightarrow$ Rényi-DP
+[10] Distributional Privacy Distributional Privacy $\Rightarrow$ DP
+[53] Bayesian-DP Bayesian DP $\Rightarrow$ DP
+[14] Mutual-Information DP MI-DP $\Leftrightarrow$ DP (asymptotically)
+Our Work $\xi$-MBP, $\epsilon$-ABP MBP $\Rightarrow$ ABP
   • 1
-    All "DP"s are shorthand for differential privacy.
+    All "DP"s denote differential privacy variants
   • 2
-    This article proposes Computational Differential Privacy (CDP). SIM-CDP and IND-CDP are two concrete kinds of CDP.
+    Introduces Computational Differential Privacy (CDP)
   • 3
-    Definition of SIM∀∃-CDP is extended from SIM-CDP.
+    SIM∀∃-CDP extends SIM-CDP
   • 4
-    Gaussian DP is a specific family of $f$-DP.
-The introduction of differential privacy has catalyzed the development of numerous privacy metrics.
-Dwork and Rothblum (2016) proposed Concentrated Differential Privacy (Concentrated-DP), which relaxes differential privacy based on the mean privacy leakage of a randomized mechanism.
-Mironov (2017) introduced Rényi Differential Privacy (RDP), a variant based on Rényi divergence that unifies concentrated-DP, $\epsilon$-DP, and $(\epsilon,\delta)$-DP.
-Mironov et al.
-(2009) developed Computational Differential Privacy (CDP), which measures privacy against computationally-bounded adversaries.
-They introduced IND-CDP and SIM-CDP, establishing that SIM-CDP implies SIM∀∃-CDP, which in turn implies IND-CDP.
-Dong et al.
-(2019) presented f-Differential Privacy (f-DP) and Gaussian Differential Privacy as relaxations of traditional DP.
-Blum et al.
-(2008) proposed Distributional Privacy, offering stronger protection than traditional differential privacy.
-Triastcyn and Faltings (2020) introduced Bayesian Differential Privacy (BDP), which considers the dataset’s distribution and implies traditional DP.
-Recent research has focused on creating taxonomies and comparing various privacy definitions.
-Cuff and Yu (2016) provided a comprehensive comparison of privacy notions, including differential privacy and its variants, introducing Mutual-Information Differential Privacy (MI-DP) and demonstrating its asymptotic equivalence to traditional differential privacy.
-Meiser (2018) examined the nuances of differential privacy, emphasizing the importance of distinguishing between approximate and probabilistic differential privacy.
-Sommer et al.
-(2019) proposed a unified approach to measuring privacy loss in sensitive data through privacy loss distribution (PLD), revealing that for non-adaptive mechanisms, privacy loss tends toward a Gaussian distribution, enabling tighter privacy bounds and more accurate calculations for mechanisms such as the Gaussian mechanism under ADP.
-Desfontaines and Pejó (2019) classified privacy definitions into seven distinct groups, each modifying different aspects of the original concept.
-These groups serve as dimensions, allowing the combination of variants from different groups to create new definitions.
-The study also established a partial ordering of their strength and identified which definitions exhibit properties like composition, post-processing, and convexity.
-Guingona et al.
-(2023) compared approximate and probabilistic differential privacy, exploring the implications of their relationship and finding that significant adjustments to one parameter necessitate substantial adjustments to another.
-Ünsal and Önen (2023) investigated the connection between differential privacy and information-theoretic measures, demonstrating how these connections can provide novel interpretations of privacy in systems employing differential privacy.
-Our work contributes to this growing body of knowledge by introducing Maximum Bayesian Privacy (MBP) and Average Bayesian Privacy (ABP), and establishing their relationships with Local Differential Privacy (LDP).
-Specifically, we prove that $\xi$-LDP implies $\xi$-MBP, which in turn implies $2\xi$-LDP.
-We also demonstrate that MBP implies ABP.
-Our approach to reductions differs from previous works in that we focus on Bayesian privacy notions and their connections to LDP.
-While previous taxonomies have primarily focused on variations of differential privacy, our work bridges the gap between traditional DP concepts and Bayesian privacy notions.
-This provides a new perspective on privacy guarantees that takes into account an adversary’s prior knowledge and inference capabilities.
-The content of our taxonomy complements existing work by:
+    Gaussian DP is a specific $f$-DP instance
+Refinements of DP: [24] introduced Concentrated DP, focusing on mean privacy leakage.
+[42]’s Rényi-DP unified various DP relaxations through $\alpha$-divergences.
+Computational perspectives emerged with [43]’s IND-CDP and SIM-CDP, which formalized privacy against bounded adversaries.
+[17] advanced $f$-DP and Gaussian DP, providing tighter composition bounds.
+Distribution-Aware Privacy: [10] proposed Distributional Privacy, offering stronger protection than traditional DP through synthetic data generation.
+[53] introduced Bayesian DP that incorporates dataset distributions while maintaining DP compatibility.
+Local DP solutions addressed scenarios without trusted third parties [13], with convergence-privacy trade-offs analyzed by [20] and practical mechanisms developed by [16].
+Taxonomic Frameworks: Several studies systematically compared privacy notions.
+[14] established asymptotic equivalence between Mutual-Information DP and traditional DP.
+[41] clarified distinctions between approximate and probabilistic DP.
+[15] classified privacy definitions into seven dimensions with partial ordering.
+[30] compared approximate and probabilistic DP, while [54] connected DP to information-theoretic measures.
+Our contribution introduces Maximum Bayesian Privacy (MBP) and Average Bayesian Privacy (ABP) as novel Bayesian privacy metrics.
+Crucially, we integrate the established equivalence between Local DP and Maximum Bayesian Privacy from [34] into our framework, while demonstrating new connections to robustness.
+Specifically:
+  • We formalize the known relationship: $\xi$-LDP $\Rightarrow$ $\xi$-MBP $\Rightarrow$ $2\xi$-LDP [34]
+  • We prove a novel quantitative relationship: MBP $\Rightarrow$ ABP
+This synthesis bridges Bayesian privacy frameworks with traditional DP concepts, providing new perspectives for adversarial settings where prior knowledge affects privacy risks.
+Our taxonomy complements existing work by: (1) Introducing Bayesian metrics sensitive to prior distributions, (2) Establishing quantitative conversions between Bayesian privacy notions (MBP and ABP), and (3) Enabling principled algorithm design for contexts with informed adversaries.
+2.3 Intersections of Privacy and PAC Learning
+The convergence of privacy guarantees and Probably Approximately Correct (PAC) learning frameworks has generated profound theoretical insights.
+Seminal work by [36] established that most concept classes learnable non-privately remain learnable under DP constraints, albeit with polynomial sample complexity increases.
+Their discovery of equivalence between local private learning and statistical query models demonstrated privacy need not fundamentally limit learnability.
+Computational aspects were explored by [11], who delineated connections between private PAC learning and online learning paradigms.
+[27] revealed a profound equivalence between differentially private sample complexity and randomized one-way communication complexity, bridging two distinct theoretical domains.
+Practical implementations advanced with [3]’s differentially private online convex optimization that maintained near-optimal regret bounds.
+Our work extends this research trajectory by investigating how privacy-preserving mechanisms inherently confer PAC robustness.
+While prior studies focused primarily on learnability and sample complexity, we establish formal connections between Bayesian privacy frameworks (MBP/ABP) and resilience to data perturbations.
+This provides new theoretical foundations for developing algorithms that simultaneously ensure: (1) Provable privacy against reconstruction attacks, and (2) Certified robustness against adversarial inputs–dual objectives previously addressed separately.
+2.4 Robustness Measurements
+Robustness quantification has evolved along two primary dimensions: adversarial resistance and distributional stability.
+[32] introduced the critical distinction between corruption robustness (average-case performance under natural corruptions) and perturbation robustness (targeted adversarial resistance), framing robustness as a multi-faceted requirement.
+For neural networks, [8] formalized point-wise robustness through adversarial frequency and severity metrics, enabling granular vulnerability assessment.
+Theoretical foundations advanced with [58]’s decomposition of robust classification error into natural and boundary components, elucidating accuracy-robustness trade-offs.
+Practical certification methods emerged through [12]’s randomized smoothing approach, which scaled certifiable robustness to ImageNet-scale problems.
+These developments collectively established robustness measurement as a rigorous subfield with both theoretical depth and practical applicability.
+Our robustness framework builds upon these foundations while introducing PAC-formalized robustness guarantees that explicitly connect to privacy-preserving mechanisms.
+This dual perspective enables unified analysis of algorithms providing both privacy and robustness certificates.
+2.5 Relation between Robustness and Privacy
+The intrinsic connection between algorithmic privacy and robustness has emerged as a fundamental research frontier, with key relationships summarized in Table 3.
+Early connections were identified by [22], who demonstrated how robust statistical estimators naturally enable differentially private estimation through the Propose-Test-Release paradigm.
+[46] later established formal equivalences between Rényi-DP and distributional robustness through $D_{\lambda}$ divergences.
+Three significant research threads have developed:
   1. 1.
-    Introducing new Bayesian privacy metrics (MBP and ABP)
+    Privacy-to-Robustness: [39]’s PixelDP framework demonstrated how $(\epsilon,\delta)$-DP mechanisms can provide certified robustness guarantees.
+[7] introduced robust privacy as a generalized indistinguishability criterion.
   2. 2.
-    Establishing concrete relationships between these Bayesian metrics and LDP
+    Robustness-to-Privacy: [6] showed robust estimators can be transformed into private estimators via $\rho$-smooth inverse sensitivity mechanisms.
+[33] established black-box reductions from robustness to privacy with optimal trade-offs.
   3. 3.
-    Providing a bridge between traditional DP concepts and Bayesian privacy notions
-These results offer new insights into the design of privacy-preserving algorithms, particularly in scenarios where adversaries may have prior knowledge about the data distribution.
-By situating our work within the broader context of privacy taxonomy research, we contribute to a more comprehensive understanding of privacy definitions and their interrelationships, aiding researchers and practitioners in selecting appropriate privacy metrics for their specific use cases.
-II-C Intersections of Privacy and PAC Learning
-The dynamic interplay between privacy and learning theory, with a spotlight on PAC learning, has attracted considerable scholarly inquiry.
-In groundbreaking research, Kasiviswanathan et al.
-(2011) scrutinized the concept classes amenable to learning under differential privacy.
-Their findings revealed that a vast array of classes that are learnable by non-private algorithms can also be mastered privately, albeit with a polynomial increase in sample size.
-This study highlighted the striking equivalence between local private learning algorithms and statistical query learning, underscoring that privacy constraints do not substantially impede the learning process.
-Concurrently, Bun (2020) delved into the computational nuances of private PAC learning, drawing a clear line between online learning and private PAC learning paradigms.
-Their work shed light on the evolving landscape of integrating privacy considerations into learning theory, unveiling both the formidable challenges and promising opportunities in the pursuit of efficient private learning.
-In a parallel vein, Feldman and Xiao (2014) explored the nexus between differential privacy and communication complexity within the realm of learning algorithms.
-They established a profound equivalence between the sample complexity of differentially private learning and the randomized one-way communication complexity required to evaluate concepts.
-This pivotal discovery is monumental as it forges a link between two distinct domains, offering valuable insights into the development of private learning algorithms and presenting a novel lens through which to view the sample complexity bounds of learning tasks.
-In a more recent development, Agarwal and Singh (2017) made significant strides in the context of online learning by crafting differentially private algorithms for online convex optimization.
-They demonstrated that it is feasible to uphold privacy without markedly sacrificing the regret bounds, thus underscoring the intricate equilibrium that must be struck between privacy assurances and learning efficacy in dynamic environments.
-Building upon these foundational intersections between privacy and PAC learning, our work extends the discourse by delving into the robustness implications of privacy guarantees.
-While previous studies have predominantly focused on sample complexity and learnability, we take a deeper dive into how privacy-preserving mechanisms can enhance robustness in PAC learning scenarios, and vice versa, how robust learning algorithms can be harnessed to design privacy-preserving mechanisms.
-By establishing theoretical linkages between our proposed privacy notions—MBP and ABP—and PAC robustness, we offer a fresh perspective on the privacy-robustness interplay.
-Our findings augment the existing body of literature by shedding light on how privacy guarantees can bolster the resilience of learning algorithms to data perturbations, and how robust learning techniques can be leveraged to safeguard individual privacy.
-II-D Robustness Measurements
-Measuring the robustness of classifiers, especially neural network-based ones, has been an active research area in recent years.
-Hendrycks and Dietterich (2019) proposed two new notions for measuring robustness: corruption robustness and perturbation robustness, which consider a classifier’s average-case performance on corruptions $\mathcal{C}$ and perturbations $\xi$, respectively.
-In contrast, adversarial robustness measures the worst-case performance on small, additive, classifier-tailored perturbations or corruptions.
-For neural network classifiers, Bastani et al.
-(2016) formalized the notion of $(x_{\ast},\epsilon)$-robust and defined point-wise robustness, $\rho(f,x_{\ast})$, as the $L_{\infty}$ distance from $x_{\ast}$ to the nearest adversarial example.
-They proposed two statistics to estimate these quantities: adversarial frequency and adversarial severity.
-In the context of binary classification, Zhang et al.
-(2019) characterized model robustness using the robust classification error under the threat model of bounded $\epsilon$ perturbation, decomposing the robust error into the sum of natural error and boundary error to characterize the trade-off between accuracy and robustness.
-Cohen et al.
-(2019) introduced the concept of certifiable robustness, where a classifier is considered certifiably robust if, for any input $\mathbf{x}$, one can easily obtain a guarantee that the classifier’s prediction remains constant within some set around $\mathbf{x}$.
-By employing randomized smoothing, they demonstrated the feasibility of certified defense on the ImageNet dataset.
-These works provide valuable insights and tools for quantifying and understanding the robustness of classifiers, paving the way for developing more robust models in the face of various types of perturbations and corruptions.
-II-E Relation between Robustness and Privacy
-The relationship between privacy and robustness has been widely discussed in the literature (Table III).
-Dwork and Lei (2009) explored the field of robust statistics and demonstrated that robust statistical estimators provide a strong foundation for developing differentially private estimators.
-Pinot et al.
-(2019) proposed a generalized notion of adversarial robustness using the divergence $D_{\lambda}$ of probabilistic distributions and suggested using randomization methods to improve robustness against adversarial samples, discussing the similarities between differential privacy and adversarial sample robustness.
-bansal2020extending introduced the concept of robust privacy, which defines the proximity and indistinguishability between two distributions without relying on a specific notion of accuracy or the underlying data distribution.
-Asi et al.
-(2023) proposed a black-box transformation method to convert a robust statistical estimator into a private statistical estimator using the $\rho$-smooth inverse sensitivity mechanism, proving that a private estimator ($\epsilon$-DP) can naturally be robust, albeit with some loss of accuracy in certain high-dimensional statistical tasks and increased computational inefficiency due to the transformation.
-Lecuyer et al.
-(2019) introduced the PixelDP model, which provides meaningful certified robustness boundaries and reasonable assessments of model accuracy degradation across various datasets and deep neural networks, offering higher accuracy on larger models and datasets, as well as stronger robustness against p-norm attacks compared to previous authentication defense methods.
-Hopkins et al.
-(2023) presented the first black-box reduction from privacy to robustness, capable of producing private estimators with optimal trade-offs among sample complexity, accuracy, and privacy for a wide range of fundamental high-dimensional parameter estimation problems, proving that, under certain conditions, if a statistical estimation method is robust to small perturbations of the input data, it can also protect the privacy of the input data.
-Pillutla et al.
-(2022) explored the trade-offs among privacy, robustness, and communication in federated learning, proposing a robust aggregation method that replaces the weighted arithmetic mean aggregation with a generalized form of the classical multidimensional median.
-This method protects the update privacy of devices by iteratively invoking secure multi-party computation primitives and bounds the sensitivity to each device’s contribution, with numerical simulations demonstrating its effectiveness in handling data corruption and parameter update corruption.
-TABLE III: relation between privacy and robustness.
-Reference Privacy Metric Robustness Metric Relation
-Dwork and Lei (2009) $(\epsilon,\delta)$-DP M-estimators robust covariance Propose-Test-Release
-matrices robust regression paradigm
-Lecuyer et al. (2019) $(\epsilon,\delta)$-PixelDP certifiably robust $(\epsilon,\delta)-DP$+certain
-condition $\Rightarrow$ robustness
-Pinot et al. (2019) Renyi-DP $D_{\mathcal{P}(\mathcal{\gamma})}$-$(\alpha,\epsilon,\gamma)$ $D_{\mathcal{P}(\mathcal{\gamma})}$-$(\alpha,\epsilon,\gamma)$ robust
-robust $\Leftrightarrow$ Renyi-DP
-bansal2020extending Flexible Accuracy Laplace Mechanism robustness $\Leftrightarrow$ privacy
-Robust Privacy $M_{Lap}^{f,b}$
-Pillutla et al. (2022) secure median’s be compatible only with
-average oracle breakdown point more communication cost1
-Hopkins et al. (2023) $(\epsilon,\delta)$-DP $(\eta,\beta,\alpha)$-robust robustness $\Leftrightarrow$ privacy
-Asi et al. (2023) $\epsilon$-DP $(\tau,\beta,\alpha)$-robust privacy $\Leftrightarrow$ robustness2
+    Federated Settings: [45] developed median-based aggregation providing simultaneous privacy, robustness, and communication efficiency in federated learning, though with communication trade-offs.
+Our work advances all three directions through several key contributions:
 […]
-our work $\xi$-MBP $(\tau,\beta,\alpha)$-robust privacy $\Leftrightarrow$ robustness
-$\epsilon$-ABP $(\tau,\beta,\alpha)$-robust robustness $\Rightarrow$ privacy
+  • Introduce PAC-formalized robustness with explicit parameter relationships to privacy guarantees
+  • Develop quantitative conversion bounds between privacy and robustness parameters
+  • Provide the first unified framework connecting Bayesian privacy with PAC robustness
+These contributions offer new theoretical tools for co-designing algorithms that simultaneously optimize privacy and robustness–a critical requirement for trustworthy AI systems in sensitive domains.
+Table 3: Theoretical relationships between privacy and robustness frameworks
+Reference Privacy Metric Robustness Metric Theoretical Relation
+[22] $(\epsilon,\delta)$-DP M-estimators Propose-Test-Release
+Robust covariance paradigm
+Robust regression
+[39] $(\epsilon,\delta)$-PixelDP Certifiable robustness $(\epsilon,\delta)$-DP + conditions
+$\Rightarrow$ robustness
+[46] Rényi-DP $D_{\mathcal{P}(\gamma)}$-$(\alpha,\epsilon,\gamma)$ robust Equivalence:
+$D_{\mathcal{P}(\gamma)}$-robust $\Leftrightarrow$ Rényi-DP
+[7] Flexible Accuracy Laplace Mechanism Duality:
+Robust Privacy $M_{Lap}^{f,b}$ robustness $\Leftrightarrow$ privacy
+[45] Secure aggregation Median breakdown point Compatibility requires
+increased communication1
+[33] $(\epsilon,\delta)$-DP $(\eta,\beta,\alpha)$-robust Robustness $\Leftrightarrow$ privacy
+[6] $\epsilon$-DP $(\tau,\beta,\alpha)$-robust Privacy $\Rightarrow$ robustness2
+[…]
+Our work $\xi$-MBP $(\tau,\beta,\alpha)$-robust Bidirectional:
+$\epsilon$-ABP $(\tau,\beta,\alpha)$-robust privacy $\Leftrightarrow$ robustness (MBP)
+robustness $\Rightarrow$ privacy (ABP)
   • 1
-    No quantitative relation is provided in this work
+    No quantitative relationship established
   • 2
-    when corrupted ratio $\tau\approx\log(n)/n\epsilon$, private estimator is naturally robust, robust estimator can be converted to private estimator.
+    When corruption ratio $\tau\approx\log(n)/n\epsilon$
   • 3
-    when corrupted ratio $\tau\geq\frac{2(d\cdot\log(\frac{R}{\alpha_{0}}+1)+\log\frac{2}{\beta})}{n\epsilon}$, robust estimator can be converted to nearly private estimator.
-III Privacy Attack and Defense Framework
-This section delineates the adversarial landscape confronting data privacy in machine learning applications.
-We systematically outline the threat model, characterizing the adversary’s objectives, capabilities, knowledge base, and methodologies for launching attacks.
-Concurrently, we describe the defender’s strategies that underpin the protection mechanisms aimed at safeguarding sensitive information.
-It is worth noting that this privacy attack and defense framework has been widely investigated, including Zhang et al.
-(2023a); du Pin Calmon and Fawaz (2012); Zhang et al.
-(2023b); Kang et al.
-(2022); Zhang et al.
-(2023f); Asi et al.
-(2023).
-III-A Privacy Attack Model
-The threat model provides a structured perspective on the potential vulnerabilities and risks to client data privacy.
-Herein, we examine the various facets of an attacker’s strategy, including their intents, resources, intelligence, and tactics.
+    When $\tau\geq\frac{2(d\cdot\log(\frac{R}{\alpha_{0}}+1)+\log\frac{2}{\beta})}{n\epsilon}$
+3 Privacy Attack and Defense Framework
+This section formalizes the adversarial ecosystem inherent in machine learning data privacy.
+We systematically establish a comprehensive threat model characterizing adversarial objectives, capabilities, knowledge structures, and methodologies, while concurrently delineating defensive countermeasures for sensitive information protection.
+This dual-aspect framework builds upon extensive prior research [59, 60, 35, 65].
+3.1 Privacy Attack Model
+The threat model rigorously conceptualizes vulnerabilities compromising client data privacy through four adversarial dimensions:
 Attacker’s Objective
-Within our framework, we posit a scenario where a protector operates an algorithm $\mathcal{M}(\cdot)$, processing input data $\mathcal{D}$ to yield output $\breve{W}$.
-This result must then be conveyed to a recipient—who is presupposed to be the attacker.
-The attacker aspires to extract confidential details regarding $\mathcal{D}$ from the observable $\breve{W}$, while balancing the need to maintain utility within certain bounds.
-The underlying objective is to refine the attacker’s probabilistic beliefs about $\mathcal{D}$ through analysis of $\breve{W}$.
+Our framework models a protector operating algorithm $\mathcal{M}(\cdot):\mathcal{D}\rightarrow\breve{W}$ and an adversarial recipient aiming to infer confidential aspects of $\mathcal{D}$ from $\breve{W}$ while maintaining functional utility.
+Formally, the adversary seeks to refine posterior beliefs $\mathbb{P}(\mathcal{D}|\breve{W})$ through analytical scrutiny of the observable output.
 Attacker’s Capability
-Our model assumes a semi-honest (or “honest-but-curious”) attacker.
-Such an adversary adheres to protocol specifications in terms of computations but seeks to glean additional private information by inferring it from $\breve{W}$.
-By performing statistical inference on the output $\breve{W}$, the attacker aims to ascertain insights into the protector’s confidential data $\mathcal{D}$.
+We assume a semi-honest ("honest-but-curious") adversary who faithfully executes protocol specifications while illicitly reconstructing $\mathcal{D}$ through statistical analysis of $\breve{W}$, maintaining operational integrity throughout the privacy attack.
 Attacker’s Knowledge
-Based on previous descriptions, the attacker has access to the output $\breve{W}$ (though post-transformation by the protector’s mechanism, they receive a modified version $W$).
-Furthermore, the attacker may possess pre-existing knowledge about the private data $\mathcal{D}$, potentially acquired through past breaches or leaks.
-Additionally, the attacker is cognizant of the utility constraints associated with their analyses.
+The adversary possesses access to the sanitized output $W$ (the transformed version of $\breve{W}$), potential prior knowledge of $\mathcal{D}$ obtained through historical breaches, and understanding of the utility constraints governing analytical operations.
 Attacker’s Method
-With knowledge of a distorted parameter $W$, we assume that the attacker adopts a Bayesian inference attack, i.e., find a data $d$ that optimizes posterior belief, denoted by:
-$\arg\max_{d}\text{IH}(d|w)=\arg\max_{d}[\text{I}(w|d)+\text{H}(d)],$ (1)
-where $\text{IH}(d|w)$ represents the logarithm of the posterior probability density function $f_{{D}|{W}}(d|w)$, $\text{I}(w|d)$ is the likelihood function’s logarithm based on observed parameter $w$, and $\text{H}(d)$ symbolizes the prior probability distribution’s logarithm $f_{{D}}(d)$.
-According to Bayes’ theorem, maximizing the logarithm of posterior $f_{D|W}(d|w)$ relative to $D$ involves the aggregation of both $\log(f_{W|D}(w|d))$ and $\log(f_{D}(d))$.
-III-B Privacy Protection Model
-We now turn our attention to the defensive measures employed by the protector whose design is to shield sensitive datasets from unauthorized exposure.
+Given observation $W$, the attacker implements Bayesian inference to maximize posterior belief:
+$\arg\max_{d}\log f_{D|W}(d|w)=\arg\max_{d}\left[\underbrace{\log f_{W|D}(w|d)}_{\text{log-likelihood}}+\underbrace{\log f_{D}(d)}_{\text{log-prior}}\right]$ (1)
+where $f_{D|W}$, $f_{W|D}$, and $f_{D}$ denote the posterior density, likelihood function, and prior distribution respectively.
+3.2 Privacy Protection Model
+We articulate defensive measures against unauthorized disclosure through four protector dimensions:
 Protector’s Objective
-To safeguard privacy post-computation by algorithm $\mathcal{M}(\cdot):\mathcal{D}\rightarrow\breve{W}$, the protector needs to devise a privacy-preserving mechanism, which is a stochastic transformation, having the ability to map $\breve{W}\rightarrow W$, where $W$ is a less-informative derivative.
-This ensures compliance with privacy standards such as $\epsilon$-differential privacy.
+The protector engineers a stochastic privacy-preserving mechanism $\mathcal{P}:\breve{W}\rightarrow W$ to generate a carefully degraded information derivative $W$ that satisfies formal privacy guarantees under rigorous information-theoretic metrics.
 Protector’s Capability
-It is incumbent upon the protector to transmit $W$ in lieu of $\breve{W}$ to conceal the original data $\mathcal{D}$.
-To this end, the protector must be adept at generating and communicating a perturbed variant $W$, which duly obscures the private dataset $\mathcal{D}$.
-This task may be complicated by historical disclosures of $\mathcal{D}$.
+The protector possesses operational capacity to generate and transmit $W$ instead of $\breve{W}$, obscuring $\mathcal{D}$ through controlled perturbation while contending with potential information leakage from historical disclosures.
 Protector’s Knowledge
-The protector is privy to the private dataset $\mathcal{D}$, the computed output $\breve{W}$, and the intricacies of the employed algorithm $\mathcal{M}(\cdot)$.
-They are also versed in the workings of the privacy-preserving transformation.
-Additionally, an understanding of the adversary’s utility limitations is crucial for the protector to tailor the distortion of $\breve{W}$ into $W$ effectively, satisfying the informational requirements of the adversary whilst preserving the sanctity of $\mathcal{D}$.
+The protector maintains complete knowledge of the sensitive dataset $\mathcal{D}$, pristine algorithm output $\breve{W}$, computational semantics of $\mathcal{M}(\cdot)$, operational parameters of $\mathcal{P}(\cdot)$, and crucially, the adversary’s utility constraints.
 Protector’s Method
-In defence of privacy (i.e., the confidentiality of dataset $\mathcal{D}$), the protector must execute the randomized privacy-preserving mechanism, which transforms $\breve{W}$ to $W$.
-Execution of the algorithm is expected to equip it with the requisite properties for safeguarding dataset $\mathcal{D}$ against intrusion while adhering to the adopted privacy standard.
-For instance, optimized algorithms for protecting privacy under the defense and attacking framework have been designed as detailed in the studies by Zhang et al.
-(2023f) and du Pin Calmon and Fawaz (2012), while Asi et al.
-(2023) choose to apply exponential mechanism(i.e., add random noise to $\breve{W}$) to protect privacy.
-IV Connecting Privacy Measurements and PAC Robustness in Machine Learning
-This section lays the theoretical groundwork for understanding the intricate relationships between different privacy notions and their connection to algorithmic robustness.
-It commences by delving into the theoretical ties among Maximum Bayesian Privacy (MBP), Average Bayesian Privacy (ABP), and Local Differential Privacy (LDP).
-Through meticulous definitions and theorems, it elucidates how these privacy concepts are interconnected, enabling transitions between various privacy frameworks.
-Fundamental results establish that, under certain conditions, LDP implies MBP and vice versa.
-The section also explores the link between MBP and ABP, offering valuable insights for constructing privacy-preserving algorithms.
-Furthermore, it investigates the relationship between PAC robust learning algorithms and privacy preservation.
-PAC robust learning, a performance metric ensuring approximately correct learning even with slightly perturbed data, is examined in the context of privacy protection.
-The section demonstrates how to derive PAC robustness from algorithms satisfying specific privacy standards and how to construct privacy-preserving algorithms from PAC robust ones.
-These findings provide a solid theoretical foundation for designing robust learning algorithms that simultaneously offer privacy protection capabilities.
-By bridging the gap between privacy notions and robustness, this section paves the way for developing algorithms that achieve both data confidentiality and resilience to perturbations.
-IV-A Metrics for Privacy and Robustness
-This section introduces the key metrics used to quantify privacy and robustness in the proposed defense and attack framework.
-It provides detailed definitions for privacy measures such as Maximum Bayesian Privacy, Average Bayesian Privacy, and Local Differential Privacy, enabling researchers to evaluate the privacy risks associated with different models.
-The section also defines PAC Robustness as a measure of an algorithm’s consistency against data variations.
-These metrics form the foundation for developing and assessing algorithms that balance data utility and confidentiality.
-IV-A1 Privacy Measurements
-Privacy measurement in machine learning has become a crucial area of research, with various approaches proposed to quantify and evaluate the privacy guarantees of different algorithms and models.
-This section provides a detailed exposition on the metrics used to quantify privacy under the previously discussed privacy attack and defense framework, including concepts such as Maximum Bayesian Privacy, Average Bayesian Privacy, and Local Differential Privacy.
-These approaches enable researchers to quantitatively evaluate the privacy risks associated with different models and algorithms.
-As privacy concerns continue to grow, further advancements in privacy measurement techniques are critical for developing privacy-preserving machine learning solutions.
-For methodologies and analyses concerning the practical estimation of MBP and ABP, please refer to F and G.
-First, we define the concept of average Bayesian privacy and explore its correlation with the notion of maximum Bayesian privacy in Section IV-B.
-These metrics serve as crucial benchmarks in evaluating the resilience of data to inference attacks.
-The average privacy leakage (Definition IV.1) measures the variation between the adversary’s prior and posterior beliefs on private information.
-The adversary obtains the posterior belief through mounting the Bayesian inference attack on the protected model information.
-Thus, we formally call the privacy leakage the Bayesian privacy leakage.
-The protected model information may lead to the federated model with lower utility and the federated training with less efficiency in certain scenarios.
-[…]
-We use JS divergence to measure the privacy leakage instead of KL divergence.
-The advantage of JS divergence over KL divergence is that it is symmetrical, and its square root satisfies the triangle inequality (Endres and Schindelin (2003)).
-This property facilitates the quantification of the trade-offs.
-Now we introduce the average Bayesian privacy of the protection mechanism.
-Definition IV.1 (Average Bayesian Privacy).
-Let $\mathcal{M}$ represent the protection mechanism that maps from the private information $D$ to the model information $W$, and $\mathcal{G}$ represent the attacking mechanism.
-Let $\epsilon_{p}$ represent the privacy leakage, which is defined as:
-$\epsilon_{p}(\mathcal{M},\mathcal{G})=\sqrt{{\text{JS}}(F^{\mathcal{A}}||F^{\mathcal{B}})},$ (2)
-where ${\text{JS}}(F^{\mathcal{A}}||F^{\mathcal{B}})=\frac{1}{2}\int_{\mathcal{D}}f^{\mathcal{A}}_{D}(d)\log\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)+\frac{1}{2}\int_{\mathcal{D}}f^{\mathcal{B}}_{D}(d)\log\frac{f^{\mathcal{B}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)$, $F^{\mathcal{A}}$ and $F^{\mathcal{B}}$ represent the attacker’s belief distribution about $D$ upon observing the protected information and without observing any information respectively, and $f_{D}^{\mathcal{M}}(d)=\frac{1}{2}(f^{\mathcal{A}}_{D}(d)+f^{\mathcal{B}}_{D}(d))$.
-We say $\mathcal{M}$ is $\epsilon$-ABP when $\epsilon_{p}(\mathcal{M},\mathcal{G})\leq\epsilon$.
+To safeguard $\mathcal{D}$-confidentiality, the protector deploys a randomized mechanism $\mathcal{P}(\breve{W})$ that produces $W$ satisfying dual objectives:
+$\mathbb{E}[\mathcal{U}(W)]$ $\geq\tau$ (Utility preservation)
+$\mathcal{I}(\mathcal{D};W)$ $\leq\epsilon$ (Privacy guarantee)
+where $\mathcal{U}$ represents the utility function, $\tau$ denotes the minimum utility threshold, $\mathcal{I}$ quantifies privacy leakage, and $\epsilon$ establishes the privacy budget.
+This framework achieves a provable trade-off between data utility and confidentiality protection, implemented through optimized perturbation algorithms [65].
+4 Connecting Privacy Measurements and PAC Robustness in Machine Learning
+This section establishes the theoretical foundations connecting privacy frameworks with robustness guarantees in machine learning.
+We systematically investigate the mathematical relationships between Maximum Bayesian Privacy (MBP), Average Bayesian Privacy (ABP), and Local Differential Privacy (LDP), while demonstrating how these privacy notions fundamentally relate to Probably Approximately Correct (PAC) robustness.
+Through rigorous definitions and theorems, we prove equivalence relations between privacy frameworks under specific conditions and derive quantitative bounds for converting between privacy metrics.
+Crucially, we establish that privacy-preserving mechanisms inherently confer robustness properties, providing a unified theoretical framework for designing algorithms that simultaneously ensure data confidentiality and resilience to perturbations.
+These results bridge the gap between privacy and robustness research communities, offering new principles for developing trustworthy learning systems.
+4.1 Measurements for Bayesian Privacy
+This subsection formalizes the Bayesian privacy framework for quantifying inference risks in data publishing.
+We introduce three complementary privacy metrics–Maximum Bayesian Privacy (MBP), Average Bayesian Privacy (ABP), and their relationship to Local Differential Privacy (LDP)–that measure protection against training data reconstruction attacks.
+These metrics capture different aspects of privacy risk: MBP quantifies worst-case vulnerability, ABP measures expected privacy loss, and LDP provides distributional guarantees.
+The definitions establish a rigorous mathematical foundation for evaluating privacy-preserving mechanisms, enabling precise comparison between protection techniques and formal analysis of privacy-utility trade-offs.
+Definition 4.1 (Average Bayesian Privacy).
+Let $\mathcal{M}$ represent the protection mechanism mapping private information $D$ to released information $W$, and $\mathcal{G}$ represent the attacker’s inference mechanism.
+The privacy leakage $\epsilon_{p,a}$ is quantified using the square root of Jensen-Shannon divergence between:
+$\epsilon_{p,a}(\mathcal{M},\mathcal{G})=\sqrt{{\text{JS}}(F^{\mathcal{A}}\parallel F^{\mathcal{B}})}$ (2)
+where $F^{\mathcal{A}}$ denotes the attacker’s posterior belief about $D$ after observing $W$, $F^{\mathcal{B}}$ represents the prior belief without observation, and $F^{\mathcal{M}}=\frac{1}{2}(F^{\mathcal{A}}+F^{\mathcal{B}})$ serves as the reference distribution.
+The mechanism satisfies $\epsilon$-ABP when $\epsilon_{p,a}\leq\epsilon$.
 Remark:
-We operate under the assumption that the private information $D$ takes on discrete values, and then
-${\text{JS}}(F^{\mathcal{A}}||F^{\mathcal{B}})$ $=\frac{1}{2}\left[\text{KL}\left(F^{\mathcal{A}}||F^{\mathcal{M}}\right)+\text{KL}\left(F^{\mathcal{B}}||F^{\mathcal{M}}\right)\right]$
-$=\frac{1}{2}\sum_{d\in\mathcal{D}}f^{\mathcal{A}}_{D}(d)\log\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}+\frac{1}{2}\sum_{d\in\mathcal{D}}f^{\mathcal{B}}_{D}(d)\log\frac{f^{\mathcal{B}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}.$
-Now we provide the formal definition for $(\epsilon,\beta,\alpha)$-ABP.
-Definition IV.2 ($(\epsilon,\beta,\alpha)$-ABP).
-Algorithm $\mathcal{M}$ achieves $(\epsilon,\beta,\alpha)$-ABP for a distribution $P$, if it satisfies: (1)being $\epsilon$-ABP, i.e., the privacy leakage $\epsilon_{p,a}\leq\epsilon$, (2)achieving standard accuracy $\alpha$ with a confidence of $1-\beta$.
-The Maximum Bayesian Privacy (denoted by $\epsilon_{p,m}$) encapsulates the highest degree of privacy assurance that can be sustained against an adversary with partial or complete knowledge about the data distribution.
-This metric is situated within the Bayesian privacy paradigm, which assesses the risk associated with the disclosure of private information.
-It underscores the importance of maintaining privacy when data undergoes sharing or utilization for designated tasks.
-Definition IV.3 ($\xi$-Maximum Bayesian Privacy).
-Let $F$ denote the belief distribution held by the attacker regarding the private data, and let $f$ indicate the associated probability density function.
-We define the system to possess $\xi$-Maximum Bayesian Privacy if for any $w\in\mathcal{W}$ and $d\in\mathcal{D}$, the following inequality is satisfied:
-$e^{-\xi}\leq\frac{f_{D|W}(d|w)}{f_{D}(d)}\leq e^{\xi}.$
-The $\xi$-Maximum Bayesian Privacy quantifies the level of privacy protection afforded by measuring the disparities in the conditional and marginal distributions of the data.
-A smaller value of $\xi$ implies stronger privacy guarantees against potential adversaries.
-Definition IV.4 (PAC Standard Accuracy).
-Let $\mathcal{M}$ be an algorithm tasked with estimating a statistic $\mu$, and $d$ represent the private information.
-The algorithm $\mathcal{M}$ is said to possess standard accuracy $\alpha$ with confidence $1-\beta$ if:
+For discrete $D$, the JS divergence decomposes as:
+${\text{JS}}(F^{\mathcal{A}}\parallel F^{\mathcal{B}})=\frac{1}{2}\left[\text{KL}\left(F^{\mathcal{A}}\parallel F^{\mathcal{M}}\right)+\text{KL}\left(F^{\mathcal{B}}\parallel F^{\mathcal{M}}\right)\right]$
+This definition provides a symmetric, bounded measure of privacy leakage that overcomes limitations of KL divergence.
+By quantifying the information gain between prior and posterior beliefs, it captures the effectiveness of privacy protection against Bayesian inference attacks.
+The JS divergence formulation ensures the metric is well-defined, symmetric, and satisfies the triangle inequality, enabling rigorous composition and comparison of privacy mechanisms.
+Definition 4.2 ($(\epsilon,\beta,\alpha)$-ABP).
+A mechanism $\mathcal{M}$ achieves $(\epsilon,\beta,\alpha)$-Average Bayesian Privacy for distribution $P$ if: (1) It satisfies $\epsilon$-ABP ($\epsilon_{p,a}\leq\epsilon$), (2) It maintains standard accuracy $\alpha$ with confidence $1-\beta$.
+This definition establishes a comprehensive privacy-accuracy benchmark, requiring mechanisms to simultaneously bound expected privacy leakage while preserving utility.
+The joint requirement formalizes the dual objectives of privacy-preserving data analysis, where protection must not come at the expense of excessive accuracy degradation.
+The following definition formalizes worst-case privacy vulnerability against informed adversaries:
+Definition 4.3 ($\xi$-Maximum Bayesian Privacy).
+A system provides $\xi$-Maximum Bayesian Privacy if for all $w\in\mathcal{W}$ and $d\in\mathcal{D}$:
+$e^{-\xi}\leq\frac{f_{D|W}(d|w)}{f_{D}(d)}\leq e^{\xi}$
+This metric quantifies the maximum possible privacy violation by bounding the multiplicative change between prior and posterior beliefs.
+Smaller $\xi$ values indicate stronger protection against reconstruction attacks, with $\xi=0$ corresponding to perfect privacy where observations reveal no additional information.
+Definition 4.4 (PAC Standard Accuracy).
+A mechanism $\mathcal{M}$ estimating statistic $\mu$ achieves standard accuracy $\alpha$ with confidence $1-\beta$ if:
 $\|\mathcal{M}(d)-\mu\|\leq\alpha$ (3)
 holds with probability at least $1-\beta$.
-The accuracy is measured within the framework of Probably Approximately Correct (PAC) learning following Zhang et al. (2023c).
-Definition IV.5 ($(\epsilon,\beta,\alpha)$-MBP).
-Let $\mathcal{M}$ denote an algorithm designed for the estimation of statistic $\mu$.
-Algorithm $\mathcal{M}$ achieves $(\epsilon,\beta,\alpha)$-MBP for a distribution $P$, if it satisfies two criteria: being $\epsilon$-MBP, and achieving standard accuracy $\alpha$ with a confidence of $1-\beta$.
-Here, $\alpha$ symbolizes a constant deviation threshold.
-IV-A2 PAC Robustness Measurement
-In parallel with privacy, we also define robustness within the PAC framework, specifically addressing mean-estimation algorithms.
-Definition IV.6 (PAC Robust Accuracy).
-Consider $\mathcal{M}$ as an algorithm for estimating statistic $\mu$, and $d$ represent the private information.
-The algorithm $\mathcal{M}$ is considered to have robust accuracy $\alpha$ with confidence $1-\beta$, if for all datasets $d^{\prime}\neq d$, the following holds with probability at least $1-\beta$:
-$\|\mathcal{M}(d^{\prime})-\mu(P)\|\leq\alpha.$ (4)
-Consequently, $\mathcal{M}$ is termed $\alpha$-PAC-Robust with probability $1-\beta$.
-This definition essentially evaluates the consistency of the algorithm’s performance against variations in the input data.
-A robust algorithm remains accurate even when small portions of the dataset are altered, thus reflecting resilience against data perturbations.
-This definition motivates the design of an algorithm $\mathcal{A}(\mathcal{D})$ that is sufficiently accurate (and potentially randomized) for estimating a parameter $\mu(P)$, given that the input dataset $\mathcal{D}$ is drawn from a specific distribution $P$.
-Both the privacy and robustness measurements presented here play instrumental roles in developing and assessing algorithms within the intricately balanced domain of data utility and confidentiality.
-The nuanced understanding of these metrics enables practitioners to formulate strategies that concurrently enhance the security and integrity of the analytical processes while mitigating risks to data privacy.
-IV-B Theoretical Connections between Privacy Notions
-In the following, relations among MBP, ABP and LDP are investigated.
-Through mathematical definitions and theorems, we demonstrate how these concepts of privacy are interconnected and how one can transition from one privacy framework to another.
-Specifically, we will prove through theorems that an algorithm satisfying the requirements of LDP also satisfies those of MBP, and vice versa.
-Additionally, we will discuss the connection between MBP and ABP, along with how to convert from a mapping that provides maximum privacy protection to one that offers an average level of privacy protection.
-These theoretical results not only deepen our understanding of the connections between various privacy measures but also offer significant guidance for crafting and analyzing algorithms that preserve privacy.
-LDP aims to safeguard individual data contributors’ privacy, enabling the extraction of aggregate statistics or insights from the data. The concept of LDP is articulated as follows.
-Definition IV.7 ($\epsilon$-LDP).
-A randomized function $\mathcal{M}$ defined as $\epsilon$-local differentially private if, $\forall$ possible datasets $d$ and $d^{\prime}$, $forall$ subset S of the output space, the following inequality is true
-$\frac{\Pr[\mathcal{M}(d)\in S]}{\Pr[\mathcal{M}(d^{\prime})\in S]}\leq e^{\epsilon}.$ (5)
-In this formula, $\epsilon$ is non-negative, which quantifies the level of privacy protection, and $\Pr[\mathcal{M}(d)\in S]$ represents the probability that the randomized mechanism $\mathcal{M}$ outputs a result in the set S when given the dataset $d$.
-This definition states that the probability of any outcome being produced by the mechanism when given dataset $d$ should be similar to the probability of the same outcome being produced when given a dataset $d^{\prime}$, smaller $\epsilon$ values indicate stronger privacy guarantees.
-LDP ensures that individual contributions to the dataset are protected by adding noise or randomness to the results of queries or computations, making it difficult to infer specific details about any individual in the dataset.
-The following Lemma IV.1 implicates that if an algorithm provides $\xi$-LDP, then it also offers $\xi$-MBP.
-In other words, the same mechanism that provides local privacy protection, where individual data points are protected, also ensures a certain level of global privacy in a Bayesian sense.
-This means that the mechanism is not only useful for protecting individual data points but also for safeguarding the privacy of the dataset as a whole.
-Please check C-A for detailed analysis and proof.
-Lemma IV.1.
-Let MBP be defined using Definition IV.3.
-Consider a case where the adversary’s prior belief distribution follows a uniform distribution.
-That is, $f_{D}(d)=f_{D}(d^{\prime})$, $\forall d^{\prime},d\in\mathcal{D}$.
-If $f_{W|D}(\cdot|\cdot)$ is $\xi$-LDP.
-That is, the mapping $f_{W|D}(\cdot|\cdot)$ satisfies that $\forall d^{\prime},d\in\mathcal{D}$,
-$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
-then $f_{W|D}(\cdot|\cdot)$ constitutes a privacy-preserving transformation ensuring $\xi$-Maximum Bayesian Privacy, where for all $d\in\mathcal{D}\ and\ w\in\mathcal{W}$,
-$\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|\leq\xi.$ (6)
-Conversely, the following lemma states that if an algorithm offers $\xi$-MBP, it also provides $2\xi$-LDP.
-This suggests that mechanisms crafted to safeguard the privacy of the dataset as a whole, as evaluated by Bayesian privacy, also afford a degree of protection for individual data points within a local differential privacy framework.
-Lemma IV.2.
-Let MBP be defined using Definition IV.3.
-Consider a situation where the adversary’s prior belief distribution follows a uniform distribution.
-Consequently, $f_{D}(d)=f_{D}(d^{\prime})$ for any $d,d^{\prime}$.
-If an algorithm is $\xi$-Maximum Bayesian Privacy, then it is $(2\xi)$-differentially private.
-Suppose privacy-preserving mapping $f_{W|D}(\cdot|\cdot)$ ensures $\xi$-maximum Bayesian privacy.
-That is, $\xi=\max\limits_{w\in\mathcal{W},d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|$.
-Then, the function $f_{W|D}(\cdot|\cdot)$ achieves $(2\xi)$-differential privacy
-$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-2\xi},e^{2\xi}].$
-This lemma establishes a bridge between local differential privacy and maximum Bayesian privacy, showing that mechanisms satisfying one privacy criterion also offer a certain level of privacy protection according to the other criterion.
-This provides insights into the relationship between different privacy definitions and can be useful when designing and analyzing privacy-preserving algorithms in various contexts.
-With Lemma IV.1 and Lemma IV.2, we are now ready to derive the relationship between MBP and LDP.
-Theorem IV.3 (Theoretical Relationship between MBP and LDP).
-Consider a situation in which the adversary’s prior belief distribution follows a uniform distribution.
-Let MBP be defined using Definition IV.3.
-If an algorithm is $\xi$-LDP, then it is $\xi$-MBP.
-If an algorithm is $\xi$-MBP, then it is $2\xi$-LDP.
-The first conclusion of Theorem IV.4 addresses a scenario in which the prior belief distribution of an adversary closely resembles a uniform distribution, meaning that the adversary has a relatively unbiased and uniform belief about the data points in the set $d$.
-The lemma introduces an algorithm with a privacy-preserving mapping denoted as $f_{W|D}(\cdot|\cdot)$.
-This mapping provides a maximum Bayesian privacy guarantee, where privacy is quantified by the parameter $\xi$.
-The parameter $\xi$ quantifies the maximum privacy loss resulting from the mapping, determined by the logarithm of the ratio between the conditional posterior distribution of a data point $d$ given an observation $w$ and the prior distribution of $d$.
-In other words, $\xi$ measures the maximum privacy leakage under this mechanism.
-The second conclusion of Theorem IV.4 implies that if the algorithm provides $\xi$-maximum Bayesian privacy, then it also provides $(2\xi)$-differential privacy.
-This implies that the privacy-preserving mapping not only demonstrates robustness in terms of Bayesian privacy but also fulfills a corresponding concept of privacy, namely, differential privacy.
-The differential privacy guarantee of mapping is expressed in terms of the ratios of conditional probabilities.
-Specifically, the ratio of the conditional probabilities of observing $w$ given two different data points $d$ and $d^{\prime}$ falls within the range of $[e^{-2\xi},e^{2\xi}]$.
-This range specifies the extent to which the mapping distorts the likelihood of observing $w$ given different data points.
-The larger $\xi$ is, the wider this range becomes, indicating stronger privacy protection.
-This conclusion shows a connection between maximum Bayesian privacy and differential privacy.
-It illustrates that a mapping ensuring maximum Bayesian privacy, particularly when the adversary’s prior belief distribution approximates uniformity, also confers a degree of differential privacy.
-This finding holds significance in comprehending and quantifying the privacy assurances offered by such mappings within the realm of privacy-preserving data analysis.
-We could also explore the situation where the prior belief distribution of the adversary is $\epsilon$-close to a uniform distribution and the analytical approach would be akin to the current scenario.
-Theorem IV.4 (Generalized Relationship between MBP and LDP).
-Consider the scenario where the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution, namely $\forall d\text{ and }d^{\prime},\ |\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$.
-Let MBP be defined using Definition IV.3.
-If an algorithm is $\xi$-LDP, then it is $(\xi+\epsilon)$-MBP.
-If an algorithm is $\xi$-MBP, then it is $(2\xi+\epsilon)$-LDP.
-Remark: We might also explore the situation where the prior belief distribution of the adversary is $\epsilon$-close to a uniform distribution and the analytical approach would be akin to the current scenario.
-The following theorem demonstrates the relationship between average Bayesian privacy and maximum Bayesian privacy.
-Theorem IV.5 (Relationship Between MBP and ABP).
-Assume that the prior belief of the attacker satisfies that
-$\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}(d)}\in[e^{-\epsilon},e^{\epsilon}].$
-Let MBP be defined using Definition IV.3. For any $\epsilon_{p,m}\geq 0$, suppose privacy-preserving mapping $f_{W|D}(\cdot|\cdot)$ ensuring $\epsilon_{p,m}$-Maximum Bayesian Privacy. Namely,
-$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\epsilon_{p,m}},e^{\epsilon_{p,m}}],$
-for any $w\in\mathcal{W}$. Then, we have that the ABP (denoted as $\epsilon_{p,a}$) is bounded by
-$\epsilon_{p,a}\leq\frac{1}{\sqrt{2}}\sqrt{(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right)},$
-where $\epsilon_{p,a}$ is introduced in Definition IV.1.
-The following theorem demonstrates the generalized relationship between average Bayesian privacy and maximum Bayesian privacy.
-This theorem establishes a generalized relationship between MBP and ABP under the assumption that the ratio of the attacker’s prior belief $f^{\mathcal{B}}_{D}(d)$ and the true prior distribution $f_{D}(d)$ is bounded within the range $[e^{-\epsilon},e^{\epsilon}]$.
-The theorem states that if a privacy-preserving mapping $f_{W|D}(\cdot|\cdot)$ guarantees $\epsilon_{p,m}$-MBP, i.e., the ratio of the posterior belief $f_{D|W}(d|w)$ and the prior belief $f_{D}(d)$ is bounded within the range $[e^{-\epsilon_{p,m}},e^{\epsilon_{p,m}}]$ for any $w\in\mathcal{W}$ and $d\in\mathcal{D}$, then the ABP (denoted as $\epsilon_{p,a}$) is bounded by a function of $\epsilon_{p,m}$ and $\epsilon$.
-Specifically, the theorem provides an upper bound on $\epsilon_{p,a}$ in terms of $\epsilon_{p,m}$ and $\epsilon$, which can be used to analyze the relationship between MBP and ABP under certain conditions on the attacker’s prior belief.
-This result is useful for understanding the connections between different notions of Bayesian privacy and can be leveraged to design privacy-preserving algorithms with desired properties.
-Theorem IV.6 (Generalized Relationship Between MBP and ABP).
-Assume that the prior belief of the attacker satisfies that
-$\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}(d)}\in[e^{-\epsilon},e^{\epsilon}].$
-Let MBP be defined using Definition IV.3. For any $\epsilon_{p,m}\geq 0$, let $f_{W|D}(\cdot|\cdot)$ be a privacy preserving mapping that guarantees $\epsilon_{p,m}$-maximum Bayesian privacy. That is,
-$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\epsilon_{p,m}},e^{\epsilon_{p,m}}],$
-for any $w\in\mathcal{W}$. Then, we have that the ABP (denoted as $\epsilon_{p,a}$) is bounded by
-$\epsilon_{p,a}\leq\frac{1}{\sqrt{2}}\sqrt{(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right)},$
-where $\epsilon_{p,a}$ is introduced in Definition IV.1.
-IV-C PAC Robustness of Private Algorithms
-This subsection establishes a significant result linking the $(\epsilon,\beta,\alpha)$-MBP privacy criterion to PAC robustness.
-It proves that any algorithm satisfying MBP inherently possesses a quantifiable level of robustness against data perturbations, with probabilistic guarantees.
-This finding bridges the gap between theoretical privacy and practical robustness, offering insights for designing machine learning models that are both privacy-preserving and resilient.
-The following Theorem IV.7 states that an algorithm $\mathcal{M}$, conforming to the $(\epsilon,\beta,\alpha)$-MBP privacy criterion, exhibits a specified degree of robustness against perturbations in the dataset.
-The robustness metric is defined as $(\frac{\log(1/\gamma)}{2n\epsilon},\alpha)$, representing the bounds within which the algorithm’s output is expected to vary.
-This level of robustness is probabilistically assured with a confidence of $1-\beta/\gamma$.
-The parameter $\gamma$ functions as a tunable confidence level within the exclusive range of 0 and 1, while $n$ quantifies the dataset’s size, and $\beta$ indicates the algorithm’s permissible error rate.
-Theorem IV.7 (PAC Robustness of $\epsilon$-MBP Protection Mechanisms).
-Let $\mathcal{M}$ be an $(\epsilon,\beta,\alpha)$-MBP (Definition IV.5) algorithm.
-Then $\mathcal{M}$ is $\alpha$-PAC-Robust (Definition IV.6) with probability $1-\gamma$, where $\gamma=(1+4\epsilon)\beta$, $n$ represents the size of the dataset, and $\beta$ represents the error probability.
-This theorem elucidates a significant implication for privacy-preserving algorithms in machine learning, specifically those adhering to the $(\epsilon,\beta,\alpha)$-MBP model.
-It articulates that any algorithm $\mathcal{M}$ satisfying this MBP criterion inherently possesses a quantifiable level of PAC robustness.
-This robustness is defined in terms of $\alpha$ and $\gamma$, where $\gamma$ is a parameter that influences the confidence level, $n$ is the dataset size, and $\beta$ is the permitted error probability.
-Crucially, the theorem assures that with probability $1-\beta/\gamma$, the algorithm will maintain its robust performance despite potential variations in the data, thereby offering a probabilistic guarantee of robustness in the face of data perturbations.
-This bridges the gap between theoretical privacy guarantees and practical robustness in applied machine learning models.
-This theorem is significant as it establishes a formal connection between robustness in the presence of learning errors and measurable criteria for privacy protection.
-It sets the stage for exploring the crucial relationship between data privacy and algorithmic robustness in machine learning systems.
-It highlights the importance of understanding and quantifying this connection to develop algorithms that simultaneously protect sensitive information and maintain performance in the face of perturbations or adversarial attacks.
-The theorem motivates the need for theoretical frameworks and metrics to analyze the privacy-robustness trade-off.
-V Conclusions
-In this paper, we have explored the intricate relationships between different privacy notions, namely LDP, ABP, and MBP, and their implications for algorithmic robustness.
-We established theoretical connections between these privacy concepts, proving that LDP implies MBP and vice versa under certain conditions, and deriving bounds between MBP and ABP.
-These findings provide valuable insights for constructing privacy-preserving algorithms.
-We also investigated the relationship between PAC robust learning and privacy preservation, demonstrating how to derive PAC robustness from privacy-preserving algorithms and construct privacy-preserving algorithms from PAC robust ones.
-This connection highlights the interplay between privacy and robustness in machine learning.
-Our findings contribute to the growing body of research on privacy and robustness in machine learning, providing a deeper understanding of the connections between different privacy notions and their implications for algorithmic robustness.
-By leveraging these insights, researchers and practitioners can work towards building more secure and trustworthy machine learning systems that safeguard sensitive information while maintaining robustness to perturbations and adversarial attacks.
-Future research directions include further exploration of the privacy-robustness trade-off, development of efficient algorithms that achieve both data confidentiality and resilience to perturbations, and investigation of the practical implications of these theoretical findings in real-world machine learning applications.
-As the field of machine learning continues to evolve, ensuring the privacy and robustness of these algorithms will remain a critical challenge and an area of active research.
-Appendix A Relationship between LDP and MBP
-In this section, we derive the relationship between LDP and MBP.
-A-A LDP implies MBP
+This definition formalizes utility preservation within the Probably Approximately Correct framework, requiring estimates to remain close to true values with high probability.
+Definition 4.5 ($(\epsilon,\beta,\alpha)$-MBP).
+Mechanism $\mathcal{M}$ achieves $(\epsilon,\beta,\alpha)$-Maximum Bayesian Privacy for distribution $P$ if: (1) It satisfies $\epsilon$-MBP per Definition 4.3, (2) It achieves standard accuracy $\alpha$ with confidence $1-\beta$.
+This combined criterion establishes a rigorous standard for privacy-preserving estimation, requiring both worst-case privacy protection and utility preservation simultaneously.
+4.2 Measurements for PAC Robustness
+We now formalize robustness within the PAC framework, complementing the privacy definitions.
+This robustness metric quantifies algorithmic stability against dataset perturbations, which is crucial for deployment in adversarial environments or non-stationary data distributions.
+Definition 4.6 (PAC Robust Accuracy).
+A mechanism $\mathcal{M}$ estimating statistic $\mu$ achieves robust accuracy $\alpha$ with confidence $1-\beta$ if for all perturbed datasets $d^{\prime}$:
+$\|\mathcal{M}(d^{\prime})-\mu(P)\|\leq\alpha$ (4)
+holds with probability at least $1-\beta$ over both mechanism randomness and dataset perturbations. We say $\mathcal{M}$ is $\alpha$-PAC-Robust with confidence $1-\beta$.
+This definition requires algorithms to maintain accuracy under arbitrary dataset perturbations, formalizing resilience against data corruption, adversarial examples, and distribution shifts.
+The probabilistic guarantee ensures reliability even when facing worst-case data manipulations.
+4.3 Theoretical Connections between Privacy Notions
+We now establish fundamental equivalence relationships between privacy frameworks.
+These theorems enable translation between privacy metrics and provide theoretical foundations for comparing protection guarantees across different privacy paradigms.
+The following lemma establishes the equivalence between Maximum Bayesian Privacy and Local Differential Privacy under bounded prior distributions:
+Lemma 4.1 (Generalized Relationship between MBP and LDP adapted from theorem 1 of  [34]).
+For any mechanism $\mathcal{M}$:
+  • If $\mathcal{M}$ satisfies $\xi$-LDP, then it provides $\xi$-MBP
+  • If $\mathcal{M}$ satisfies $\xi$-MBP, then it provides $2\xi$-LDP
+This theorem establishes a quantitative equivalence between the Bayesian and differential privacy frameworks.
+The first implication shows that LDP mechanisms automatically provide Bayesian privacy guarantees, while the converse demonstrates that Bayesian private mechanisms yield differential privacy.
+The $\epsilon$ parameter quantifies the deviation from uniform priors, with smaller values tightening the equivalence bounds.
+This result enables direct comparison between privacy frameworks and allows practitioners to leverage results from one paradigm in the other.
+The next theorem quantifies how worst-case privacy protection (MBP) bounds expected privacy leakage (ABP):
+Theorem 4.2 (Generalized Relationship Between MBP and ABP).
+Assume the attacker’s prior belief satisfies $\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}(d)}\in[e^{-\epsilon},e^{\epsilon}]$. Any $\epsilon_{p,m}$-MBP mechanism bounds the ABP leakage by:
+$\epsilon_{p,a}\leq\frac{1}{\sqrt{2}}\sqrt{(\epsilon_{p,m}+\epsilon)\left(e^{\epsilon_{p,m}+\epsilon}-1\right)}$
+This result establishes that worst-case privacy protection (MBP) implies bounds on average-case privacy leakage (ABP).
+The bound tightens as $\epsilon_{p,m}$ decreases, showing that strong MBP guarantees necessarily enforce strong ABP protection.
+The $\epsilon$ parameter captures prior distribution mismatch, with exact priors ($\epsilon=0$) yielding the tightest bound.
+This theorem provides crucial guidance for mechanism design: by optimizing for MBP, designers automatically obtain ABP guarantees, simplifying the privacy certification process.
+4.4 PAC Robustness of Bayesian Private Algorithms
+We now establish the foundational connection between privacy protection and algorithmic robustness. The following theorem demonstrates that privacy mechanisms inherently confer robustness properties:
+Theorem 4.3 (PAC Robustness of $\epsilon$-MBP Protection Mechanisms).
+Any $(\epsilon,\beta,\alpha)$-MBP mechanism $\mathcal{M}$ (Definition 4.5) is $\alpha$-PAC-Robust (Definition 4.6) with probability $1-\gamma$ where $\gamma=(1+4\epsilon)\beta$.
+This profound result establishes that privacy-preserving mechanisms inherently exhibit robustness to dataset perturbations.
+The proof (provided in Appendix) shows that the information-limiting properties of private mechanisms constrain how much outputs can vary under input changes, thereby enforcing stability.
+The robustness confidence parameter $\gamma$ depends explicitly on the privacy parameter $\epsilon$, revealing a quantitative trade-off: stronger privacy ($\epsilon\downarrow$) directly improves robustness guarantees ($\gamma\downarrow$).
+This provides a principled approach to designing robust algorithms through privacy mechanisms, with applications in adversarial learning and secure data publishing.
+The theorem bridges two previously disconnected research domains: privacy preservation and adversarial robustness.
+By demonstrating that privacy mechanisms inherently provide robustness certificates, it enables joint optimization of both properties.
+This has significant implications for designing next-generation trustworthy AI systems that require both data protection and operational reliability.
+The explicit parameter relationships guide practitioners in balancing privacy-robustness trade-offs for specific application requirements.
+5 Theoretical Relationship between Data Privacy and Input Robustness
+In this section, we introduce the measurements for data privacy and input robustness.
+These metrics offer a nuanced perspective on the algorithm’s performance in safeguarding sensitive information and its ability to withstand certain distortions.
+5.1 Measurements of Data Privacy
+First, we introduce the measurement for data privacy.
+Definition 5.1 (Data Privacy).
+Let $x_{t}^{(i)}$ represent the $i$-th data recovered by the attacker at the $t$-th round of the optimization algorithm, and $x_{o}^{(i)}$ represent the $i$-th original data.
+The privacy leakage is measured using
+$\epsilon_{p}$ $=1-\mathbb{E}\left[\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\frac{1}{T}\sum_{t=1}^{T}\frac{||x_{t}^{(i)}-x_{o}^{(i)}||}{D}\right],$ (5)
+where the expectation is taken with respect to the randomness of the mini-batch $\mathcal{D}$.
+5.2 Measurements of Input Robustness
+In assessing the robustness of a machine learning algorithm, we adhere to the following definition.
+Definition 5.2 (Input Robustness).
+[37, 56] Let $P$ represent a distribution. A model $\mathcal{A}$ is $(r,\epsilon,P)$-input-robust, if it holds that
+$\mathbb{E}_{x\sim P}[\text{sup}_{\|\delta\|_{2}\leq r}|\mathcal{A}\left(w,(x+\delta,y)\right)-\mathcal{A}\left(w,(x,y)\right)|]\leq\epsilon,$ (6)
+where $x$ denotes the data (training sample or feature), and $y$ denotes the corresponding label.
+Remark: The robustness measures the model performance in terms of the data distortion $\|\delta\|$.
+5.3 Technical Assumptions
+The theoretical analysis relies on three fundamental regularity conditions that establish the mathematical foundation for our privacy-robustness framework.
+These assumptions are standard in optimization theory yet critical for establishing our main theoretical results.
+Assumption 5.1 (L-Lipschitz Gradient Continuity).
+The gradient mapping of the loss function $\nabla\mathcal{L}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ satisfies:
+$\|\nabla\mathcal{L}(\theta,x)-\nabla\mathcal{L}(\theta,y)\|_{2}\leq L\|x-y\|_{2}$ (7)
+for all $x,y\in\mathcal{X}\subset\mathbb{R}^{d}$ and fixed model parameters $\theta$, where $L>0$ denotes the Lipschitz constant.
+Remark.
+This standard condition in optimization theory ensures bounded gradient variation under input perturbations.
+The Lipschitz constant $L$ quantifies the maximum rate of gradient change relative to input variations, which is fundamental for convergence analysis and stability guarantees in adversarial settings.
+Assumption 5.2 (Bi-Lipschitz Gradient Condition).
+There exist constants $0<c_{a}\leq c_{b}<\infty$ such that for all $(x_{1},y),(x_{2},y)\in\mathcal{X}\times\mathcal{Y}$ with $\|x_{1}-x_{2}\|_{2}\leq D$:
+$c_{a}\|x_{1}-x_{2}\|_{2}\leq\|\nabla\mathcal{L}(\theta,(x_{1},y))-\nabla\mathcal{L}(\theta,(x_{2},y))\|_{2}\leq c_{b}\|x_{1}-x_{2}\|_{2}$ (8)
+where $D$ denotes the maximum pairwise data distance.
+Remark.
+This condition strengthens Assumption 5.1 by establishing metric equivalence between data space and gradient space [48].
+The lower bound ($c_{a}>0$) ensures gradient injectivity necessary for data reconstruction, while the upper bound ($c_{b}$) maintains controlled sensitivity.
+This holds for strongly convex losses with Lipschitz gradients and prevents pathological landscape scenarios.
+Assumption 5.3 (Optimal Regret Bound for Gradient Matching).
+The attacker’s reconstruction algorithm achieves $\Theta(\sqrt{T})$ regret:
+$c_{0}\sqrt{T}\leq\sum_{t=1}^{T}\|\nabla\mathcal{L}(\theta,(x_{t},y_{t}))-\nabla\mathcal{L}(\theta,(\tilde{x},y_{t}))\|_{2}\leq c_{2}\sqrt{T}$ (9)
+where $T$ is the attack horizon, $x_{t}$ the reconstruction at step $t$, and $\tilde{x}$ the target satisfying $\nabla\mathcal{L}(\theta,(\tilde{x},y))=\tilde{g}$.
+Remark.
+This $\Theta(\sqrt{T})$ regret bound reflects practical attack capabilities with state-of-the-art optimizers:
+[…]
+[…]
+[…]
+where $d$ is data dimension and $\kappa$ the condition number. The regret bound enables tractable analysis while capturing real-world attack efficiency.
+5.4 Theoretical Relationship between Data Privacy and Input Robustness
+This section establishes a rigorous theoretical connection between data privacy and input robustness in machine learning algorithms.
+We introduce precise mathematical measures for both privacy leakage and adversarial robustness, providing a unified framework to analyze an algorithm’s dual capabilities in protecting sensitive information while maintaining performance under input perturbations.
+Our central contribution is a quantitative trade-off theorem demonstrating that privacy-preserving mechanisms fundamentally influence an algorithm’s resilience to adversarial manipulations.
+This theoretical foundation has profound implications for designing next-generation algorithms that simultaneously optimize both privacy guarantees and robustness properties.
+The following lemma establishes a foundational bound on privacy leakage in terms of parameter distortion.
+It demonstrates that when a semi-honest attacker attempts to reconstruct the original training data from protected model parameters, the achievable privacy leakage is explicitly constrained by the magnitude of intentional distortions introduced during parameter protection.
+Lemma 5.4 (The Relationship between Data Privacy and Parameter Distortion [62]).
+Assume that the semi-honest attacker employs an optimization algorithm to infer client $k$’s original dataset $d$ from the released parameter $W$.
+Let $g(d)=W$ represent the parameter generation function, where $\breve{d}$ denotes the dataset corresponding to the original parameter $\breve{W}$, and $g(d)=\partial\mathcal{L}(W,d)/\partial W$ computes the gradient of the loss function.
+Define the parameter distortion metric as $\Delta=\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||g(d^{(m)})-g(\breve{d}^{(m)})||$, where $d^{(m)}$ and $\breve{d}^{(m)}$ denote the $m$-th data points in their respective datasets.
+When the expected regret of the optimization algorithm over $I>0$ rounds is $\Theta(I^{p})$ and $\Delta\geq\frac{2c_{2}c_{b}}{c_{a}}\cdot I^{p-1}$, the privacy leakage $\epsilon_{p}$ satisfies:
+$\epsilon_{p}\leq 1-\frac{c_{a}\cdot\Delta+c_{a}\cdot c_{0}\cdot I^{p-1}}{4D}.$ (10)
+This lemma provides a crucial theoretical guarantee: intentional distortion in model parameters directly translates to provable privacy protection.
+The bound establishes that sufficient parameter distortion ($\Delta$) forces the privacy leakage $\epsilon_{p}$ below a computable threshold, creating a mathematically verifiable shield against training data reconstruction attacks.
+This formalizes the counterintuitive insight that strategically introduced parameter noise not only preserves privacy but does so in a quantifiable manner governed by the algorithm’s regret behavior and distortion magnitude.
+Building upon this foundation, we now present our main theorem which establishes a fundamental connection between privacy leakage and input robustness.
+The following result characterizes how the privacy-preserving modifications made to an algorithm inherently affect its resilience against input perturbations.
+Specifically, it derives a functional relationship $f(\epsilon_{p})$ that precisely quantifies how privacy guarantees influence robustness properties.
+Theorem 5.5 (Input Robustness of Data Private Algorithms).
+Let $\mathcal{A}$ be a protection mechanism with privacy leakage $\epsilon_{p}$ measured according to Definition 5.1, and robustness characterized by Definition 5.2.
+The robustness parameter $\alpha$ quantifying $\mathcal{A}$’s resilience against adversarial inputs is given by the expression:
+$\alpha=\frac{Cr}{2}+\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{2c_{a}}$ (11)
+where $r$ denotes the robustness radius, $C$ represents the Lipschitz constant, $D$ bounds the parameter space, $\mathcal{D}$ is the dataset domain, and $c_{a}$, $c_{b}$ are algorithm-dependent constants.
+This theorem establishes a profound mathematical duality: privacy protection mechanisms inherently enhance algorithmic robustness against input perturbations.
+The derived expression demonstrates that reduced privacy leakage ($\epsilon_{p}\downarrow$) directly increases the robustness parameter ($\alpha\uparrow$), revealing an inherent synergy between these seemingly competing objectives.
+This theoretical insight revolutionizes our understanding of privacy-robustness trade-offs, showing that properly designed privacy mechanisms can simultaneously provide both strong privacy guarantees and enhanced adversarial resilience.
+The precise functional relationship enables principled co-design of algorithms that jointly optimize these critical properties, opening new avenues for developing trustworthy AI systems.
+(Complete proof and analysis available in G)
+6 Conclusions
+In this paper, we thoroughly investigated the complex interplay between various privacy notions, namely Local Differential Privacy (LDP), Average Bayesian Privacy (ABP), and Maximum Bayesian Privacy (MBP), and their fundamental connection to algorithmic robustness in machine learning.
+We integrated established theoretical relationships between LDP and MBP, and further derived novel quantitative bounds connecting MBP and ABP.
+These insights are crucial for designing effective privacy-preserving algorithms.
+Furthermore, we forged a strong link between privacy and PAC robust learning.
+We showed that privacy-preserving algorithms inherently exhibit PAC robustness, and conversely, how principles of robust learning can guide the construction of privacy mechanisms.
+A key result quantifies how privacy leakage directly impacts input robustness.
+Our contributions provide a unified and deeper understanding of the synergies and trade-offs between privacy and robustness.
+This theoretical framework is vital for building machine learning systems that are not only secure against privacy breaches but also resilient to data perturbations and adversarial attacks.
+Future research will focus on further exploring the intricate privacy-robustness trade-off landscape, developing more efficient and practical algorithms that simultaneously achieve strong guarantees in both areas, and validating these theoretical findings through extensive empirical studies in real-world, high-stakes applications.
+As machine learning continues to evolve, ensuring both the privacy and robustness of its algorithms will remain paramount.
+Appendix A ACKNOWLEDGMENTS
+This work was supported by the National Science and Technology Major Project under Grant 2022ZD0115301.
+Appendix B Relationship between LDP and MBP
+B.1 Theoretical Relationship between MBP and ABP
+In this section, we derive the relationship between MBP and ABP.
 The adversary has a prior belief distribution, denoted as $f_{D}(d)$, that is very close (within $\epsilon$) to a uniform distribution.
@@
 This means that the mapping function provides strong privacy protection by limiting the information leakage about the original mini-batch data $d$ when the adversary observes $w$.
-The following lemma establishes a connection between Local Differential Privacy (LDP) and Maximum Bayesian Privacy (MBP) under the assumption of a uniform prior belief distribution of the adversary.
-It states that if a mapping $f_{W|D}(\cdot|\cdot)$ satisfies $\xi$-LDP, then it also guarantees $\xi$-MBP.
-In other words, if the ratio of the conditional probabilities $f_{W|D}(w|d)$ and $f_{W|D}(w|d^{\prime})$ is bounded within the range $[e^{-\xi},e^{\xi}]$ for any $w\in\mathcal{W}$ and $d,d^{\prime}\in\mathcal{D}$, then the mapping $f_{W|D}(\cdot|\cdot)$ is a privacy-preserving mapping that ensures the log-ratio of the posterior belief $f_{D|W}(d|w)$ and the prior belief $f_{D}(d)$ is bounded by $\xi$ for all $w\in\mathcal{W}$ and $d\in\mathcal{D}$.
-This lemma provides a useful connection between LDP and MBP, which can be leveraged to design privacy-preserving algorithms.
-Lemma A.1 (LDP implies MBP).
-Consider the scenario where the prior belief distribution of the adversary is a uniform distribution.
-That is, $f_{D}(d)=f_{D}(d^{\prime})$ for any $d,d^{\prime}\in\mathcal{D}$.
-If $f_{W|D}(\cdot|\cdot)$ is $\xi$-LDP.
-That is, the mapping $f_{W|D}(\cdot|\cdot)$ satisfies that
-$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
-then $f_{W|D}(\cdot|\cdot)$ is a privacy-preserving mapping that guarantees $\xi$-maximum Bayesian privacy.
-That is, $\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|\leq\xi$, $\forall w\in\mathcal{W},d\in\mathcal{D}$.
-Proof.
-From the definition of $\xi$-LDP, we know that
-$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
-for all $d,d^{\prime}\in\mathcal{D}$. Notice that
-$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}=\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}.$
-It implies that $\frac{f_{D|W}(d|w)}{f_{D}(d)}\leq\frac{f_{D|W}(d^{\prime}|w)}{f_{D}(d^{\prime})}\cdot e^{\xi}$.
-We assume that $f_{D}(d)=f_{D}(d^{\prime})$. Notice that
-$\sum_{d\in\mathcal{D}}f_{D|W}(d|w)=1.$
-For any fixed $\tilde{d}\in\mathcal{D}$, we have that
-11 11 $=\sum_{d\in\mathcal{D}}f_{D|W}(d|w)$
-$=f_{D|W}(\tilde{d}|w)+\sum_{d\in\mathcal{D}\setminus\tilde{d}}f_{D|W}(d|w)$
-$\leq f_{D|W}(\tilde{d}|w)+e^{\xi}\cdot f_{D|W}(\tilde{d}|w)\cdot(|\mathcal{D}|-1)$
-$=(1+e^{\xi}\cdot(|\mathcal{D}|-1))\cdot f_{D|W}(\tilde{d}|w).$
-Therefore, we have that
-$f_{D|W}(\tilde{d}|w)\geq\frac{1}{1+e^{\xi}\cdot(|\mathcal{D}|-1)}.$
-Similarly, we have that
-11 11 $=\sum_{d\in\mathcal{D}}f_{D|W}(d|w)$
-$=f_{D|W}(\tilde{d}|w)+\sum_{d\in\mathcal{D}\setminus\tilde{d}}f_{D|W}(d|w)$
-$\geq f_{D|W}(\tilde{d}|w)+e^{-\xi}\cdot f_{D|W}(\tilde{d}|w)\cdot(|\mathcal{D}|-1)$
-$=(1+e^{-\xi}\cdot(|\mathcal{D}|-1))\cdot f_{D|W}(\tilde{d}|w).$
-Therefore, we have
-$f_{D|W}(\tilde{d}|w)$ $\leq\frac{1}{1+e^{-\xi}\cdot(|\mathcal{D}|-1)}$
-$\leq\frac{e^{\xi}}{|\mathcal{D}|}.$
-Therefore, $\forall\tilde{d}\in\mathcal{D}$, we have that
-$f_{D|W}(\tilde{d}|w)\leq e^{\xi}f_{D}(\tilde{d}).$
-We say a mechanism is $\xi$-MBP if $\forall w\in\mathcal{W}^{d}$,
-$\frac{f(d|w)}{f(d)}=\frac{f(w|d)f(d)}{f(w)f(d)}=\frac{f(w|d)}{f(w)}\leq e^{\xi}.$ (7)
-Therefore, $f_{D|W}(\cdot|\cdot)$ is $\xi$-MBP. ∎
-A-B MBP implies LDP
-The following lemma illustrates that if an algorithm is $\xi$-maximum Bayesian privacy, then it is $2\xi$-Differentially Private.
-Lemma A.2 (MBP implies LDP).
-Consider the scenario where the prior belief distribution of the adversary is a uniform distribution.
-That is, $f_{D}(d)=f_{D}(d^{\prime})$ for any $d,d^{\prime}\in\mathcal{D}$.
-Let $f_{W|D}(\cdot|\cdot)$ be a privacy-preserving mapping that guarantees $\xi$-maximum Bayesian privacy.
-That is, $\xi=\max_{w\in\mathcal{W},d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|$.
-Then, the mapping $f_{W|D}(\cdot|\cdot)$ is $2\xi$-Local Differential Private:
-$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-2\xi},e^{2\xi}].$
-Proof.
-From the definition of maximum Bayesian privacy leakage, we know that
-$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\xi},e^{\xi}],$
-for any $w\in\mathcal{W}$ and $d\in\mathcal{D}$. Therefore, for any $d,d^{\prime}\in\mathcal{D}$ we have
-$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}=\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}\in[e^{-2\xi},e^{2\xi}].$
-From the definition of Differential Privacy, we know that it is $2\xi$-Differentially Private. ∎
-Combining the results in C-A and C-B, we are now ready to analyze the relationship between MBP and LDP, which is shown in the following theorem.
-Theorem A.3 (Relationship between MBP and LDP).
-[…]
-Appendix B Relationship Between Average Bayesian Privacy and Maximum Bayesian Privacy
+Appendix C Relationship Between Average Bayesian Privacy and Maximum Bayesian Privacy
 In this section, we introduce the relationship between average Bayesian privacy and maximum Bayesian privacy.
-Theorem B.1 (Relationship Between MBP and ABP).
+Theorem C.6 (Relationship Between MBP and ABP).
 Assume that the prior belief of the attacker satisfies that
 $\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}(d)}\in[e^{-\epsilon},e^{\epsilon}].$
-Let MBP be defined using Definition IV.3. For any $\epsilon_{p,m}\geq 0$, let $f_{W|D}(\cdot|\cdot)$ be a privacy preserving mapping that guarantees $\epsilon_{p,m}$-maximum Bayesian privacy. That is,
+Let MBP be defined using Definition 4.3. For any $\epsilon_{p,m}\geq 0$, let $f_{W|D}(\cdot|\cdot)$ be a privacy preserving mapping that guarantees $\epsilon_{p,m}$-maximum Bayesian privacy. That is,
 $\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\epsilon_{p,m}},e^{\epsilon_{p,m}}],$
@@
 $\epsilon_{p,a}\leq\frac{1}{\sqrt{2}}\sqrt{(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right)},$
-where $\epsilon_{p,a}$ is introduced in Definition IV.1.
+where $\epsilon_{p,a}$ is introduced in Definition 4.1.
 Proof.
@@
 $\geq e^{-(\epsilon_{p,m}+\epsilon)}\cdot f^{\mathcal{B}}_{D}(d).$ (13)
-Combining Eq. (D) and Eq. (D), we have
+Combining Eq. (C) and Eq. (C), we have
 $\left|\log\left(\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|\leq\epsilon_{p,m}+\epsilon.$
@@
 $\left|\log\left(\frac{f_{D}^{\mathcal{M}}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|\leq\epsilon_{p,m}+\epsilon.$ (14)
-Combining Eq. (D), Eq. (D) and Eq. (14), we have that
+Combining Eq. (C), Eq. (C) and Eq. (14), we have that
 $JS(F^{\mathcal{A}}||F^{\mathcal{B}})$ $\leq\frac{1}{2}\left[\int_{\mathcal{S}}\left|f^{\mathcal{A}}_{D}(d)-f^{\mathcal{B}}_{D}(d)\right|\left|\log\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right|\textbf{d}\mu(d)\right]$
@@
 $\epsilon_{p,a}=\sqrt{JS(F^{\mathcal{A}}||F^{\mathcal{B}})}\leq\frac{1}{\sqrt{2}}\sqrt{(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right)}.$
-Appendix C Generalized Relationship between LDP and MBP
-In this section, we derive the generalized relationship between LDP and MBP.
-C-A LDP implies MBP
-The following lemma asserts that when the adversary’s prior belief distribution is close to uniform, the mapping function provides a high level of privacy protection by limiting the information leakage to a maximum of $\xi$ in a Bayesian sense when the adversary observes a transformed data point $w$.
-Lemma C.1 (LDP implies MBP).
-Consider the scenario where the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution.
-That is, $\forall d,d^{\prime},\ |\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$.
-If $f_{W|D}(\cdot|\cdot)$ is $\xi$-LDP.
-That is, the mapping $f_{W|D}(\cdot|\cdot)$ satisfies that
-$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
-then $f_{W|D}(\cdot|\cdot)$ is a privacy-preserving mapping that guarantees $\xi$-maximum Bayesian privacy.
-That is, $\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|\leq\xi$, $\forall w\in\mathcal{W},d\in\mathcal{D}$.
-Proof.
-From the definition of $\xi$-LDP, we know that
-$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
-for all $d,d^{\prime}\in\mathcal{D}$. Notice that
-$e^{-\xi}\leq\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}\leq e^{\xi}.$
-It implies that $f_{D|W}(d|w)\leq f_{D|W}(d^{\prime}|w)\cdot e^{\xi}\cdot\frac{f_{D}(d)}{f_{D}(d^{\prime})}\leq f_{D|W}(d^{\prime}|w)\cdot e^{\xi+\epsilon}$.
-We assume that $\frac{f_{D}(d)}{f_{D}(d^{\prime})}\in[e^{-\epsilon},e^{\epsilon}]$. Notice that
-$\sum_{d\in\mathcal{D}}f_{D|W}(d|w)=1.$
-For any fixed $\tilde{d}\in\mathcal{D}$, we have that
-11 11 $=\sum_{d\in\mathcal{D}}f_{D|W}(d|w)$
-$=f_{D|W}(\tilde{d}|w)+\sum_{d\in\mathcal{D}\setminus\tilde{d}}f_{D|W}(d|w)$
-$\leq f_{D|W}(\tilde{d}|w)+e^{\xi+\epsilon}\cdot f_{D|W}(\tilde{d}|w)\cdot(|\mathcal{D}|-1)$
-$=(1+e^{\xi+\epsilon}\cdot(|\mathcal{D}|-1))\cdot f_{D|W}(\tilde{d}|w).$
-Therefore, we have that
-$f_{D|W}(\tilde{d}|w)\geq\frac{1}{1+e^{\xi+\epsilon}\cdot(|\mathcal{D}|-1)}.$
-Similarly, we have that
-11 11 $=\sum_{d\in\mathcal{D}}f_{D|W}(d|w)$
-$=f_{D|W}(\tilde{d}|w)+\sum_{d\in\mathcal{D}\setminus\tilde{d}}f_{D|W}(d|w)$
-$\geq f_{D|W}(\tilde{d}|w)+e^{-\xi-\epsilon}\cdot f_{D|W}(\tilde{d}|w)\cdot(|\mathcal{D}|-1)$
-$=(1+e^{-\xi-\epsilon}\cdot(|\mathcal{D}|-1))\cdot f_{D|W}(\tilde{d}|w).$
-Therefore, we have
-$f_{D|W}(\tilde{d}|w)$ $\leq\frac{1}{1+e^{-\xi-\epsilon}\cdot(|\mathcal{D}|-1)}$
-$\leq\frac{e^{\xi+\epsilon}}{|\mathcal{D}|}.$
-Therefore, $\forall\tilde{d}\in\mathcal{D}$, we have that
-$f_{D|W}(\tilde{d}|w)\leq e^{\xi+\epsilon}f_{D}(\tilde{d}).$
-We say a mechanism is $(\xi+\epsilon)$-MBP if $\forall w\in\mathcal{W}^{d}$,
-$\frac{f(d|w)}{f(d)}=\frac{f(w|d)f(d)}{f(w)f(d)}=\frac{f(w|d)}{f(w)}\leq e^{\xi+\epsilon}.$ (11)
-Therefore, $f_{D|W}(\cdot|\cdot)$ is $(\xi+\epsilon)$-MBP. ∎
-C-B MBP implies LDP
-The following lemma establishes a connection between MBP and LDP under the assumption that the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution.
-It states that if a mapping $f_{W|D}(\cdot|\cdot)$ guarantees $\xi$-MBP, then it also satisfies $(2\xi+\epsilon)$-LDP.
-In other words, if the log-ratio of the posterior belief $f_{D|W}(d|w)$ and the prior belief $f_{D}(d)$ is bounded by $\xi$ for all $w\in\mathcal{W}$ and $d\in\mathcal{D}$, and the prior belief distribution is "$\epsilon$-close" to uniform (i.e., $|\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$ for all $d,d^{\prime}$), then the ratio of the conditional probabilities $f_{W|D}(w|d)$ and $f_{W|D}(w|d^{\prime})$ is bounded within the range $[e^{-2\xi-\epsilon},e^{2\xi+\epsilon}]$ for any $w\in\mathcal{W}$ and $d,d^{\prime}\in\mathcal{D}$.
-This lemma provides a useful connection between MBP and LDP, which can be leveraged to analyze the privacy guarantees of algorithms satisfying MBP under certain conditions on the prior belief distribution.
-Lemma C.2 (MBP implies LDP).
-Consider the scenario where the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution.
-That is, $\forall d,d^{\prime},\ |\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$.
-Let $f_{W|D}(\cdot|\cdot)$ be a privacy-preserving mapping that guarantees $\xi$-maximum Bayesian privacy.
-That is, $\xi=\max_{w\in\mathcal{W},d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|$.
-Then, the mapping $f_{W|D}(\cdot|\cdot)$ is $(2\xi+\epsilon)$-Local Differential Private:
-$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-2\xi},e^{2\xi}].$
-Proof.
-From the definition of maximum Bayesian privacy leakage, we know that
-$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\xi},e^{\xi}],$
-for any $w\in\mathcal{W}$ and $d\in\mathcal{D}$. Therefore, for any $d,d^{\prime}\in\mathcal{D}$ we have
-$e^{-2\xi-\epsilon}\leq\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}\cdot e^{-\epsilon}\leq\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\leq\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}\cdot e^{\epsilon}\leq e^{2\xi+\epsilon}.$
-From the definition of Differential Privacy, we know that it is $(2\xi+\epsilon)$-Differentially Private. ∎
-Combining the results in C-A and C-B, we are now ready to analyze the generalized relationship between MBP and LDP, which is shown in the following theorem.
-Theorem C.3 (Relationship between MBP and LDP).
-Consider the scenario where the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution.
-That is, $\forall d,d^{\prime},\ |\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$.
-If an algorithm is $\xi$-LDP, then it is $(\xi+\epsilon)$-MBP.
-If an algorithm is $\xi$-maximum Bayesian privacy, then it is $(2\xi+\epsilon)$-LDP.
-Appendix D Generalized Relationship Between Average Bayesian Privacy and Maximum Bayesian Privacy
-In this section, we introduce the generalized relationship between average Bayesian privacy and maximum Bayesian privacy.
-Theorem D.1 (Relationship Between MBP and ABP).
-Assume that the prior belief of the attacker satisfies that
-$\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}(d)}\in[e^{-\epsilon},e^{\epsilon}].$
-Let MBP be defined using Definition IV.3. For any $\epsilon_{p,m}\geq 0$, let $f_{W|D}(\cdot|\cdot)$ be a privacy preserving mapping that guarantees $\epsilon_{p,m}$-maximum Bayesian privacy. That is,
-$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\epsilon_{p,m}},e^{\epsilon_{p,m}}],$
-for any $w\in\mathcal{W}$. Then, we have that the ABP (denoted as $\epsilon_{p,a}$) is bounded by
-$\epsilon_{p,a}\leq\frac{1}{\sqrt{2}}\sqrt{(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right)},$
-where $\epsilon_{p,a}$ is introduced in Definition IV.1.
-Proof.
-Let $F^{\mathcal{M}}=\frac{1}{2}(F^{\mathcal{A}}+F^{\mathcal{B}})$. We have
-$JS(F^{\mathcal{A}}||F^{\mathcal{B}})$ $=\frac{1}{2}\left[KL\left(F^{\mathcal{A}}||F^{\mathcal{M}}\right)+KL\left(F^{\mathcal{B}}||F^{\mathcal{M}}\right)\right]$
-$=\frac{1}{2}\left[\int_{\mathcal{S}}f^{\mathcal{A}}_{D}(d)\log\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)+\int_{\mathcal{S}}f^{\mathcal{B}}_{D}(d)\log\frac{f^{\mathcal{B}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)\right]$
-$=\frac{1}{2}\left[\int_{\mathcal{S}}f^{\mathcal{A}}_{D}(d)\log\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)-\int_{\mathcal{S}}f^{\mathcal{B}}_{D}(d)\log\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\textbf{d}\mu(d)\right]$
-$\leq\frac{1}{2}\left[\int_{\mathcal{D}}\left|f^{\mathcal{A}}_{D}(d)-f^{\mathcal{B}}_{D}(d)\right|\left|\log\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right|\textbf{d}\mu(d)\right],$
-where the inequality is due to $\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\leq\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}$.
-Therefore, we have that
-$JS(F^{\mathcal{A}}||F^{\mathcal{B}})\leq\frac{1}{2}\left[\int_{\mathcal{D}}\left|f^{\mathcal{A}}_{D}(d)-f^{\mathcal{B}}_{D}(d)\right|\left|\log\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right|\textbf{d}\mu(d)\right].$
-Let $w^{\star}=\arg\max_{w\in\mathcal{W}}f_{D|W}(d|w)$. Then we have that
-[…]
-[…]
-$\leq e^{\epsilon_{p,m}}\cdot f_{D}(d)$
-$\leq e^{\epsilon_{p,m}+\epsilon}\cdot f^{\mathcal{B}}_{D}(d).$ (12)
-Similarly, let $w_{l}=\arg\min_{w\in\mathcal{W}}f_{D|W}(d|w)$. Then we have that
-[…]
-[…]
-$\geq e^{-\epsilon_{p,m}}\cdot f_{D}(d)$
-$\geq e^{-(\epsilon_{p,m}+\epsilon)}\cdot f^{\mathcal{B}}_{D}(d).$ (13)
-Combining Eq. (D) and Eq. (D), we have
-$\left|\log\left(\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|\leq\epsilon_{p,m}+\epsilon.$
-Now we analyze according to the following two cases.
-Case 1: $f^{\mathcal{B}}_{D}(d)\leq f^{\mathcal{A}}_{D}(d)$. Then we have
-$\left|\log\left(\frac{f_{D}^{\mathcal{M}}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|=\log\left(\frac{f_{D}^{\mathcal{M}}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\leq\log\left(\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\leq\left|\log\left(\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|\leq\epsilon_{p,m}+\epsilon.$
-Case 2: $f^{\mathcal{B}}_{D}(d)\geq f^{\mathcal{A}}_{D}(d)$. Then we have
-$\left|\log\left(\frac{f_{D}^{\mathcal{M}}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|=\log\left(\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}^{\mathcal{M}}(d)}\right)\leq\log\left(\frac{f^{\mathcal{B}}_{D}(d)}{f^{\mathcal{A}}_{D}(d)}\right)\leq\left|\log\left(\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|\leq\epsilon_{p,m}+\epsilon.$
-Therefore,
-$\left|\log\left(\frac{f_{D}^{\mathcal{M}}(d)}{f^{\mathcal{B}}_{D}(d)}\right)\right|\leq\epsilon_{p,m}+\epsilon.$ (14)
-Combining Eq. (D), Eq. (D) and Eq. (14), we have that
-$JS(F^{\mathcal{A}}||F^{\mathcal{B}})$ $\leq\frac{1}{2}\left[\int_{\mathcal{S}}\left|f^{\mathcal{A}}_{D}(d)-f^{\mathcal{B}}_{D}(d)\right|\left|\log\frac{f^{\mathcal{M}}_{D}(d)}{f^{\mathcal{B}}_{D}(d)}\right|\textbf{d}\mu(d)\right]$
-$\leq\frac{1}{2}(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right).$
-Therefore,
-$\epsilon_{p,a}=\sqrt{JS(F^{\mathcal{A}}||F^{\mathcal{B}})}\leq\frac{1}{\sqrt{2}}\sqrt{(\epsilon_{p,m}+\epsilon)\cdot\left(e^{\epsilon_{p,m}+\epsilon}-1\right)}.$
-Appendix E Analysis for Theorem IV.7
-Theorem E.1 (PAC Robustness of $\epsilon$-MBP Protection Mechanisms).
-Let $\mathcal{M}$ be an $(\epsilon,\beta,\alpha)$-MBP (Definition IV.5) algorithm.
+C.1 Theoretical Relationship between PAC Robustness and MBP
+Theorem C.7 (PAC Robustness of $\epsilon$-MBP Protection Mechanisms).
+Let $\mathcal{M}$ be an $(\epsilon,\beta,\alpha)$-MBP (Definition 4.5) algorithm.
 Then $\mathcal{M}$ is $\alpha$-PAC-Robust with probability $1-\gamma$, where $\gamma\in(0,1)$, $n$ represents the size of the dataset, and $\beta=\frac{\gamma}{1+4\epsilon}$ represents the error probability.
@@
 $\mathcal{M}$ is an $(\epsilon,\beta,\alpha)$-MBP algorithm for the estimation of statistic $\mu$.
-We know that the protection mechanism $\mathcal{M}$ is $\epsilon$-MBP (measured using Definition IV.3).
-From the relationship between MBP and LDP (Theorem IV.4), we know that $\mathcal{M}$ is $2\epsilon$-LDP.
+We know that the protection mechanism $\mathcal{M}$ is $\epsilon$-MBP (measured using Definition 4.3).
+From the relationship between MBP and LDP (Lemma 4.1), we know that $\mathcal{M}$ is $2\epsilon$-LDP.
 From the definition of private algorithm, we know that for the original private information $d$, we have that
@@
 (16), the third and fourth inequalities are due to $0\leq\epsilon\leq 1$, and the last inequality is due to the relationship between $\beta$ and $\epsilon$.
-From Definition IV.6, $\mathcal{M}$ is $\alpha$-PAC-Robust with probability $1-\gamma$.
-Appendix F Practical Estimation for Maximum Bayesian Privacy
+From Definition 4.6, $\mathcal{M}$ is $\alpha$-PAC-Robust with probability $1-\gamma$.
+Appendix D Practical Estimation for Maximum Bayesian Privacy
 $\xi$ represents the maximum privacy leakage over all possible released information $w$ from the set $\mathcal{W}$ and all the dataset $d$, expressed as
 $\xi=\max_{w\in\mathcal{W},d}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|.$ (23)
-When the extent of the attack is fixed, $\xi$ becomes a constant. Given Assumption G.1, we can infer that
+When the extent of the attack is fixed, $\xi$ becomes a constant. Given Assumption E.1, we can infer that
 $\xi$ $=\max_{d}\left|\log\left(\frac{f_{D|W}(d|w^{*})}{f_{D}(d)}\right)\right|.$ (24)
@@
 The task of estimating the maximum Bayesian Privacy $\xi$ reduces to determining the value of $\hat{f}_{D|W}(d|w_{m})$, which will be addressed in the subsequent section.
-F-A Estimation for $\hat{f}_{{D}|{W}}(d|w_{m})$
+D.1 Estimation for $\hat{f}_{{D}|{W}}(d|w_{m})$
 We first generate a set of models $\{w_{m}\}_{m=1}^{M}$ stochastically: to generate the $m$-th model $w_{m}$, we run an algorithm (e.g., stochastic gradient descent), using a mini-batch $\mathcal{S}$ randomly sampled from $\mathcal{D}=\{z_{1},\ldots,z_{|\mathcal{D}|}\}$.
@@
 We then estimate $\hat{f}_{{D}|{W}}(d|w_{m})$ with $\{w_{m}\}_{m=1}^{M}$.
-As mentioned in subsection III-A, the attacker can conduct Bayesian inference attack, i.e., estimate dataset $d$ with $\hat{d}$ that optimizes posterior belief, denoted by:
+Recall that the attacker can conduct Bayesian inference attack, i.e., estimate dataset $d$ with $\hat{d}$ that optimizes posterior belief, denoted by:
 $\hat{d}=\arg\max_{d}\text{log}f_{D|W}(d|w)=\arg\max_{d}\text{log}[f_{W|D}(w|d)\times f^{\mathcal{B}}_{D}(d)],$ (27)
@@
 By performing multiple iterations and accumulating counts, the algorithm provides an estimation of the conditional probability.
-F-B Estimation Error of MBP
+D.2 Estimation Error of MBP
 To streamline the error bound analysis, it is presumed that both the data and model parameters take on discrete values.
@@
 This means that the larger the number $T$ or the smaller the error parameter $\epsilon$, the more confident we are that our estimate is accurate.
-For the detailed analysis, please refer to Zhang et al.
-(2023e).
-Lemma F.1.
+For the detailed analysis, please refer to [64].
+Lemma D.8.
 With probability at least $1-2\exp\left(\frac{-\epsilon^{2}T\kappa_{1}(d)}{3}\right)$, we have that
 $|\hat{\kappa}_{1}(d)-\kappa_{1}(d)|\leq\epsilon\kappa_{1}(d).$ (28)
-Appendix G Practical Estimation for Average Bayesian Privacy
-In the scenario of federated learning, Zhang et al.
-(2023e) derived the upper bound of privacy leakage based on the specific form of $\text{TV}(P^{\mathcal{O}}||P^{\mathcal{D}})$ of $\mathcal{M}$ according to this equation: $\widetilde{\epsilon}_{p}=2C_{1}-C_{2}\cdot{\text{TV}}(P^{\mathcal{O}}||P^{\mathcal{D}})$.
+Appendix E Practical Estimation for Average Bayesian Privacy
+In the scenario of federated learning, [64] derived the upper bound of privacy leakage based on the specific form of $\text{TV}(P^{\mathcal{O}}||P^{\mathcal{D}})$ of $\mathcal{M}$ according to this equation: $\widetilde{\epsilon}_{p}=2C_{1}-C_{2}\cdot{\text{TV}}(P^{\mathcal{O}}||P^{\mathcal{D}})$.
 In this section, we elaborate on the methods we propose to estimate $C_{1}$ and $C_{2}$.
 First, we introduce the following assumption.
-Assumption G.1.
+Assumption E.1.
 We assume that the amount of Bayesian privacy leaked from the true parameter is maximum over all possible parameters. That is, for any $w\in\mathcal{W}$, we have that
@@
 where $w^{*}$ represents the true model parameter.
-G-A Estimation for $C_{1}=\sqrt{{\text{JS}}(F^{\mathcal{O}}||F^{\mathcal{B}})}$
+E.1 Estimation for $C_{1}=\sqrt{{\text{JS}}(F^{\mathcal{O}}||F^{\mathcal{B}})}$
 $C_{1}$ quantifies the average square root of the Jensen-Shannon divergence between the adversary’s belief distribution about the private information before and after observing the unprotected parameter, independent of the employed protection mechanisms.
@@
 $\hat{C}_{1}^{2}=\frac{1}{2}\int_{\mathcal{D}}\hat{\kappa}_{1}\log\frac{\hat{\kappa}_{1}}{\frac{1}{2}(\hat{\kappa}_{1}+\kappa_{2})}\,d\mu(d)+\frac{1}{2}\int_{\mathcal{D}}\kappa_{2}\log\frac{\kappa_{2}}{\frac{1}{2}(\hat{\kappa}_{1}+\kappa_{2})}\,d\mu(d).$
-G-B Estimation for $C_{2}=\frac{1}{2}(e^{2\xi}-1)$
+E.2 Estimation for $C_{2}=\frac{1}{2}(e^{2\xi}-1)$
 Recall that the Maximum Bayesian Privacy, is defined as $\xi=\max_{w\in\mathcal{W},d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|$ and represents the maximal privacy leakage across all possible information $w$ released.
 The quantity $\xi$ remains constant when the extent of the attack is fixed.
-Under Assumption G.1, we have that
+Under Assumption E.1, we have that
 $\xi=\max_{d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w^{*})}{f_{D}(d)}\right)\right|.$ (31)
@@
 $\hat{C}_{2}=\frac{1}{2}\left(e^{2\hat{\kappa}_{3}}-1\right).$ (33)
-Ultimately, the computation of $C_{1}$ and $C_{2}$ hinges on the estimation of $\hat{f}_{{D}|{W}}(d|w_{m})$, which is further elaborated upon in Section F-A.
-G-C Estimation Error of ABP
+Ultimately, the computation of $C_{1}$ and $C_{2}$ hinges on the estimation of $\hat{f}_{{D}|{W}}(d|w_{m})$, which is further elaborated upon in Section D.1.
+E.3 Estimation Error of ABP
 In this section, we analyze the estimation error of ABP.
 To facilitate the analysis of the error bound, we consider that both the data and the model parameter are in discrete form.
-For the detailed analysis, please refer to Zhang et al.
-(2023e).
+For the detailed analysis, please refer to [64].
 The following theorem establishes that a bounded estimation error for $\kappa_{1}(d)$—specifically, not exceeding $\epsilon$ times the value of $\kappa_{1}(d)$—implies a bounded estimation error for another variable, $C_{1}$.
@@
 In summary, the theorem provides a calculated limit for how large the estimation error for $C_{1}$ can be, given the known estimation error for $\kappa_{1}(d)$.
-Theorem G.1 (The Estimation Error of $C_{1}$).
+Theorem E.9 (The Estimation Error of $C_{1}$).
 Assume that $|\kappa_{1}(d)-\hat{\kappa}_{1}(d)|\leq\epsilon\kappa_{1}(d)$. We have that
@@
 This threshold is important in applications where it is crucial to know how accurately the privacy parameter can be estimated.
-Theorem G.2.
+Theorem E.10.
 We have that
 $|\hat{\epsilon}_{p}-\epsilon_{p}|\leq\frac{3}{2}\cdot\sqrt{\frac{(1+\epsilon)\log\frac{1+\epsilon}{1-\epsilon}}{2}+\frac{\epsilon+\max\left\{\log(1+\epsilon),\log\frac{1}{(1-\epsilon)}\right\}}{2}}.$ (35)
+Appendix F Analysis for Theorem 5.5
+In this section, we introduce the relationship between privacy leakage and distortion extent.
+Now we introduce the definition of privacy leakage.
+The privacy leakage is measured using the gap between the estimated dataset and the original dataset.
+The semi-honest attacker uses an optimization algorithm ([67, 29, 66, 57]) to reconstruct the original dataset of the client given the exposed model information.
+Let $d^{(i)}_{m}$ represent the reconstructed $m$-th data at iteration $i$.
+Let $D$ be a positive constant satisfying that $||d^{(i)}_{m}-d||\in[0,D]$, $\forall i$ and $m$.
+The privacy leakage is measured using protection extent and attacking extent.
+Assume that the semi-honest attacker uses an optimization algorithm to infer the original dataset of client $k$ based on the released parameter $w$.
+Let $\Delta=||W-\breve{W}||$ represent the distortion of the model parameter, where $\breve{W}$ represents the original parameter, and $W$ represents the protected parameter.
+The regret of the optimization algorithm in a total of $I$ rounds is $\Theta(I^{p})$.
+Definition F.1 (Privacy Leakage).
+Let $d$ represent the original private dataset, and $d^{(i)}$ represent the dataset inferred by the attacker at iteration $i$, $I$ represent the total number of learning rounds.
+The privacy leakage $\epsilon_{p}$ is defined as
+$\epsilon_{p}=\left\{\begin{array}[]{cl}\frac{D-\frac{1}{I}\sum_{t=1}^{T}\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d_{I}^{(m)}-\breve{d}^{(m)}||}{D},&T>0\\ 0,&T=0\\ \end{array}\right.$ (36)
+Remark:
+(1) We assume that $||d^{(i)}-d||\in[0,D]$. Therefore, $\epsilon_{p}\in[0,1]$.
+(2) If the estimated data of the adversary $d^{(i)}$ is equal to the original data $d$, i.e., $d^{(i)}=d$, then the privacy leakage is maximal. In this case, the privacy leakage is equal to 111.
+(3) When the adversary does not attack ($I=0$), the privacy leakage $\epsilon_{p}=0$ from Eq. (36).
+To derive bounds for privacy leakage, we need the following assumptions.
+Assumption F.1.
+Assume that $||d||\leq 1$.
+Assumption F.2.
+For any two datasets $d_{1}$ and $d_{2}$, assume that $c_{a}||g(d_{1})-g(d_{2})||\leq||d_{1}-d_{2}||\leq c_{b}||g(d_{1})-g(d_{2})||$.
+Assumption F.3.
+Assume that $c_{0}\cdot I^{p}\leq\sum_{i=1}^{I}||g(d^{(i)})-g(d)||=\Theta(I^{p})\leq c_{2}\cdot I^{p}$, where $d^{(i)}$ represents the dataset reconstructed by the attacker at round $I$, $d$ represents the dataset satisfying that $g(d)=w$, and $g(d^{(i)})=\frac{\partial\mathcal{L}(d^{(i)},w)}{\partial w}$ represents the gradient of the reconstructed dataset at round $I$.
+With the regret bounds of the optimization algorithms, we are now ready to provide bounds for privacy leakage.
+The following lemma presents bounds for privacy leakage, following the analyses in [60, 62, 65].
+Lemma F.11.
+Assume that the semi-honest attacker uses an optimization algorithm to infer the original dataset of client $k$ based on the released parameter $W$.
+Let $d$ represent the dataset satisfying that $g(d)=W$, and $\breve{d}$ represent the dataset satisfying that $g(\breve{d})=\breve{W}$, where $\breve{W}$ represents the original parameter, $W$ represents the protected parameter, and $g(d)=\frac{\partial\mathcal{L}(W,d)}{\partial W}$.
+Let $d^{(m)}$ represent $m$-th data of dataset $d$, $\breve{d}^{(m)}$ represent $m$-th data of dataset $\breve{d}$.
+Let $\Delta=\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||g(d^{(m)})-g(\breve{d}^{(m)})||$ represent the distortion of the parameter.
+The expected regret of the optimization algorithm in a total of $I$ ($I>0$) rounds is $\Theta(I^{p})$.
+If $\Delta\geq\frac{2c_{2}c_{b}}{c_{a}}\cdot I^{p-1}$, then we have that
+$\epsilon_{p}\leq 1-\frac{c_{a}\cdot\Delta+c_{a}\cdot c_{0}\cdot I^{p-1}}{4D}.$ (37)
+Proof.
+Recall the privacy leakage of client $k$ $\epsilon_{p}^{(k)}$ is defined as
+$\epsilon_{p}^{(k)}=\left\{\begin{array}[]{cl}\frac{D-\frac{1}{I}\sum_{i=1}^{I}\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d_{i}^{(m)}-\breve{d}^{(m)}||}{D},&I>0\\ 0,&I=0\\ \end{array}\right.$ (38)
+To protect privacy, client $k$ selects a protection mechanism $M_{k}$, which maps the original parameter $\breve{W}$ to a protected parameter $W$.
+After observing the protected parameter, a semi-honest adversary infers the private information using the optimization approaches.
+Let $d_{i}$ represent the reconstructed data at iteration $i$ using the optimization algorithm.
+Therefore the cumulative regret over $I$ rounds
+$R(I)$ $=\sum_{i=1}^{I}[||g(d_{i}^{(m)})-W^{(m)}||-||g(d^{(m)})-W^{(m)}||]$
+$=\sum_{i=1}^{I}[||g(d_{i}^{(m)})-g(d^{(m)})||]$
+$=\Theta(I^{p}).$
+Therefore, we have
+$c_{0}\cdot I^{p}\leq\sum_{i=1}^{I}||g(d_{i}^{(m)})-g(d^{(m)})||=\Theta(I^{p})\leq c_{2}\cdot I^{p},$ (39)
+where $c_{0}$ and $c_{2}$ are constants independent of $I$.
+Let $x$ and $\widetilde{x}$ represent two data. From our assumption, we have that
+$c_{a}||g(x)-g(\widetilde{x})||\leq||x-\widetilde{x}||\leq c_{b}||g(x)-g(\widetilde{x})||.$ (40)
+Let $d^{(m)}_{i}$ represent the reconstructed $m$-th data at iteration $i$ using the optimization algorithm. We have that
+$\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d_{i}^{(m)}-\breve{d}^{(m)}||$ $\geq\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d^{(m)}-\breve{d}^{(m)}||-\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d_{i}^{(m)}-d^{(m)}||$
+$\geq c_{a}\cdot\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||g(d^{(m)})-g(\breve{d}^{(m)})||-c_{b}\cdot\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||g(d_{i}^{(m)})-g(d^{(m)})||$
+$\geq c_{a}\cdot||\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}(g(d^{(m)})-g(\breve{d}^{(m)}))||-c_{b}\cdot\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||g(d_{i}^{(m)})-g(d^{(m)})||$
+where the second inequality is due to $\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d^{(m)}-\breve{d}^{(m)}||\geq c_{a}\cdot\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||g(d^{(m)})-g(\breve{d}^{(m)})||$ and $\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d_{i}^{(m)}-d^{(m)}||\leq\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}c_{b}||g(d_{i}^{(m)})-g(d^{(m)})||$.
+We denote $g^{(k)}$ as the original gradient of client $k$, and $\widetilde{g}^{(k)}$ as the distorted gradient uploaded from client $k$ to the server. Therefore, we have that
+$\|\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}(g(d^{(m)})-g(\breve{d}^{(m)}))\|$ $=\|\widetilde{g}^{(k)}-g^{(k)}\|$ (41)
+$=\|\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}(g(\breve{d}^{(m)})+\delta^{(k)}-g(\breve{d}^{(m)}))\|$ (42)
+$=\|\delta^{(k)}\|$ (43)
+$=\Delta^{(k)}.$ (44)
+Therefore, we have
+$D(1-\epsilon_{p}^{(k)})=\frac{1}{I}\sum_{i=1}^{I}\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d_{i}^{(m)}-\breve{d}^{(m)}||$ $\geq c_{a}\Delta^{(k)}-c_{b}\cdot\frac{1}{I}\cdot\frac{1}{|\mathcal{D}^{(k)}|}\sum_{i=1}^{I}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||g(d_{i}^{(m)})-g(d^{(m)})||$
+$\geq c_{a}\Delta^{(k)}-c_{2}\cdot c_{b}I^{p-1}$
+$\geq\frac{1}{2}\max\{c_{a}\Delta^{(k)},c_{2}\cdot c_{b}I^{p-1}\}$
+$\geq\frac{c_{a}\Delta^{(k)}+c_{2}\cdot c_{b}I^{p-1}}{4}.$
+Therefore, we have that
+$\epsilon_{p}^{(k)}\leq 1-\frac{c_{a}\Delta^{(k)}+c_{2}\cdot c_{b}I^{p-1}}{4D}.$
+Therefore, we have that
+$\epsilon_{p}^{(k)}$ $\leq 1-\frac{c_{a}\Delta^{(k)}+c_{2}\cdot c_{b}I^{p-1}}{4D}$ (45)
+$=1-\frac{c_{a}\|\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}(g(d^{(m)})-g(\breve{d}^{(m)}))\|+c_{2}\cdot c_{b}I^{p-1}}{4D}.$ (46)
+Appendix G Analysis for Theorem 5.5
+Theorem G.12 (Relationship between Privacy and Robustness).
+Let the privacy leakage of the protection mechanism $\mathcal{A}$ be measured using Definition 5.1, and let the robustness of $\mathcal{A}$ be measured using Definition 5.2.
+Let the privacy leakage be $\epsilon_{p}$, then $\left(r,\frac{Cr}{2}+\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{2c_{a}}\right)$, where $c_{a}$ and $D$ represent constants, $\mathcal{D}$ represents the dataset, and $\gamma$ represents the error rate.
+Proof.
+The distortion extent measures the discrepancy between the gradient of the original data and the distorted data.
+From the relationship between distortion extent and privacy leakage (Lemma 5.4), we know that the distortion extent is bounded as follows:
+$\|\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}(g(d^{(m)})-g(\breve{d}^{(m)}))\|\leq\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{c_{a}}.$ (47)
+On one hand, we have that
+$\|\nabla\mathcal{L}(\theta,(X+\delta,y))-\nabla\mathcal{L}(\theta,(X,y))\|$ $=\|\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}(g(d^{(m)})-g(\breve{d}^{(m)}))\|$ (48)
+$\leq\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{c_{a}}.$ (49)
+On the other hand, from Lipschitz’s condition, we know that
+$\|\nabla\mathcal{L}(\theta,(X+\delta,y))-\nabla\mathcal{L}(\theta,(X,y))\|\leq C\|\delta\|,$ (50)
+where $C$ represents a constant. Therefore, we have that
+$\|\nabla\mathcal{L}(\theta,(X+\delta,y))-\nabla\mathcal{L}(\theta,(X,y))\|\leq\frac{C\|\delta\|}{2}+\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{2c_{a}}.$ (51)
+Therefore, we have that
+$\mathbb{E}_{X\sim P}[\text{sup}_{\|\delta\|_{p}\leq r}|\nabla\mathcal{L}(\theta,(X+\delta,y))-\nabla\mathcal{L}(\theta,(X,y))|]$ (52)
+$\leq\frac{C\|\delta\|}{2}+\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{2c_{a}}$ (53)
+$\leq\frac{Cr}{2}+\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{2c_{a}}.$ (54)
+Recall that a model is $(r,\epsilon)$-input-robust, if it holds that
+$\mathbb{E}_{X\sim P}[\text{sup}_{\|\delta\|_{p}\leq r}|\nabla\mathcal{L}\left(w,(X+\delta,y)\right)-\nabla\mathcal{L}\left(w,(X,y)\right)|]\leq\epsilon,$ (55)
+where $X$ denotes the data (training sample or feature), and $y$ denotes the corresponding label.
+Therefore, from the definition of robustness, we know that the model is $\left(r,\frac{Cr}{2}+\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{2c_{a}}\right)$-input-robust. ∎